### v1

Abstract
Classical statistical decision theory evaluates treatment choices based solely on observed outcomes. However, by ignoring counterfactual outcomes, it cannot assess the quality of decisions relative to feasible alternatives. For example, the quality of a physician’s decision may depend not only on patient survival, but also on whether a less invasive treatment could have produced a similar result. To address this limitation, we extend standard decision theory to incorporate counterfactual losses–criteria that evaluate decisions using all potential outcomes. The central challenge in this generalization is identification: because only one potential outcome is observed for each unit, the associated risk under a counterfactual loss is generally not identifiable. We show that under the assumption of strong ignorability, a counterfactual risk is identifiable if and only if the counterfactual loss function is additive in the potential outcomes. Moreover, we demonstrate that additive counterfactual losses can yield treatment recommendations that differ from those based on standard loss functions, provided that the decision problem involves more than two treatment options.
Keywords: causal inference, expected utility, identification, policy learning, potential outcomes, risk
1 Introduction
Statistical decision theory, pioneered by Wald [1950], formulates decision-making as a game against nature. A decision-maker selects an action $D=d$ in the face of an unknown state of nature $\theta$, and incurs a prespecified loss or negative utility $\ell(d;\theta)$, which quantifies the consequence of the action when the true state of nature is $\theta$. Better decisions correspond to lower losses. Given covariates $\bm{X}$, we can define a decision rule $D=\pi(\bm{X})$ that maps observed data to actions. The performance of a decision rule is commonly evaluated by its risk, that is, the expected loss:
$R(\pi;\theta,\ell)\ =\ \mathbb{E}\left[\ell(\pi(\bm{X});\theta)\right],$
where the expectation is taken across over the distribution of $\bm{X}$. Since $\theta$ is unknown, the true risk is typically unidentifiable. This motivates the use of optimality criteria such as admissibility, minimax, minimax-regret, and Bayes risk [Wald, 1950, Savage, 1972, Berger, 1985]. Furthermore, statistical decision theory also provides a foundational framework for classical estimation theory [e.g., Lehmann and Casella, 2006] and hypothesis testing [e.g., Lehmann et al., 1986].
In a series of influential papers, Manski [2000, 2004, 2011] applies this framework to the problem of treatment choice. In its simplest form, the goal is to choose a treatment $D$ to minimize the expected loss associated with an outcome of interest $Y$. Because the treatment may influence the outcome, the loss $\ell(d;Y(d))$ depends on the potential outcome under the chosen treatment $d$, denoted by $Y(d)$ [Neyman, 1923, Rubin, 1974], and possibly the treatment itself. As before, we can define a treatment rule $D=\pi(\bm{X})$ based on observed covariates, and evaluate its performance via risk:
$R(\pi;\ell)\ =\ \mathbb{E}{\left[\ell(\pi(\bm{X});Y(\pi(\bm{X})))\right]},$
where the expectation is taken over the joint distribution of $\bm{X}$ and $Y(\pi(\bm{X}))$. Since only one potential outcome is observed per unit, the risk is generally unidentifiable unless the treatment rule coincides with the one used to generate the data. Consequently, additional assumptions — most notably, strong ignorability of treatment assignment [Rosenbaum and Rubin, 1983] — are required to identify the risk and determine optimal treatment rules [e.g., Manski, 2000, Kitagawa and Tetenov, 2018, Athey and Wager, 2021].
In this paper, we generalize the standard framework of statistical decision theory for treatment choice by introducing a broader class of loss functions. Specifically, we consider counterfactual losses, which depend on the full set of potential outcomes rather than the outcome under the selected treatment alone. That is, the loss takes the form $\ell(d;\{Y(d^{\prime})\}_{d^{\prime}\in\mathcal{D}}\})$, where $\mathcal{D}$ denotes the support of the treatment variable $D$. We refer to this as a counterfactual loss because it evaluates each decision in light of the alternative outcomes that could have occurred had a different treatment been chosen. Given this loss, we define the counterfactual risk for measuring the performance of treatment rule $\pi$ as:
$R(\pi;\ell)\ =\ \mathbb{E}[\ell(\pi(\bm{X});\{Y(d)\}_{d\in\mathcal{D}})],$
where the expectation is taken over the distribution of covariates $\bm{X}$ and all potential outcomes $\{Y(d)\}_{d\in\mathcal{D}}$.
The idea of counterfactual loss includes familiar examples such as regret [Savage, 1972] as special cases. A regret-based loss can be defined as $\ell(d;\{Y(d^{\prime})\}_{d^{\prime}\in\mathcal{D}\}})=\max_{d^{\prime}\in\mathcal{D}}Y(d^{\prime})-Y(d)$. However, our definition of counterfactual risk differs from another notion of regret commonly used in the reinforcement learning literature, where maximization occurs outside the expectation [e.g., Lattimore and Szepesvári, 2020]. Specifically, the expression of this version of regret, which is given by $\max_{\pi^{\prime}}\mathbb{E}[Y(\pi^{\prime}(\bm{X}))]-\mathbb{E}[Y(\pi(\bm{X}))]$, is not equivalent to $R(\pi;\ell)=\mathbb{E}[\max_{\pi^{\prime}}Y(\pi^{\prime}(\bm{X}))]-\mathbb{E}[Y(\pi(\bm{X}))]$.
A key challenge introduced by counterfactual losses is that their associated risks are generally not identifiable, even under the standard strong ignorability assumption. Although this assumption enables identification of the standard risk in treatment choice settings, it is insufficient for identifying counterfactual risk when the loss depends on unobserved potential outcomes.
The main contribution of this paper is to propose a structural condition on the counterfactual loss that enables identification of counterfactual risk. Specifically, we show that under strong ignorability, the additivity of the counterfactual loss in potential outcomes is a necessary and sufficient condition for identifying the difference in counterfactual risk between any two decision rules. This result, presented in Section 3, implies that when the counterfactual loss is additive, one can not only identify the optimal treatment rule but also evaluate risk differences between competing rules. Moreover, we also show that counterfactual risks give treatment recommendations different from those based on standard risks, provided that the decision problem involves more than two treatment options.
2 Illustrative Examples
To clarify the ideas introduced above, we now present a series of illustrative examples in the context of medical decision-making.
2.1 Binary decision
Consider a stylized setting in which a physician must choose whether to provide a patient with a standard care treatment ($D=1$) or do nothing ($D=0$). The outcome of interest is patient survival, denoted by $Y$, where $Y=1$ indicates survival and $Y=0$ otherwise. For notational simplicity, we assume no covariates are observed.
Suppose we evaluate the physician’s decision using the standard loss $\ell(D;Y(D))$, which assigns a value to each pair of decision and realized outcome $(D,Y(D))$. Let $c_{1}>0=c_{0}$ denote the cost of providing treatment relative to doing nothing. We also use $\ell_{y}$ to denote the loss associated with the observed outcome $Y(D)=y$. For example, if the physician administers treatment and the patient does not survive, the loss is $\ell(1;0)=\ell_{0}+c_{1}$. If the physician does nothing and the patient survives, the loss is $\ell(0;1)=\ell_{1}$. Since patient survival is a desired outcome, we assume $\ell_{1}<\ell_{0}$. This is an example of a standard loss without counterfactuals which takes the form,
$\ell(D;Y(D))=\ell_{Y(D)}+c_{D}.$ (1)
Decision
[…]
Negative ($Y(0)=0$) False Negative (FN) True Positive
$\ell_{0}$$\tilde{\ell}_{0}+c_{1}$
Positive ($Y(0)=1$) True Negative False Positive (FP)
$\ell_{1}$$\tilde{\ell}_{1}+c_{1}$
Potential Outcomes Negative ($Y(1)=0$) True Negative False Positive (FP)
~$\tilde{\ell}_{0}$ $\ell_{0}+c_{1}$
Positive ($Y(1)=1$) False Negative (FN) True Positive
~$\tilde{\ell}_{1}$ $\ell_{1}+c_{1}$
Table 1: Confusion matrix for each combination of potential outcome $Y(d)=y_{d}$ and decision $D=d$. Each cell is assigned a loss $\ell(d;y_{0},y_{1})$ for $d,y_{0},y_{1}\in\{0,1\}$. Darker colors indicate realized outcomes, while lighter colors indicate counterfactual outcomes.
Coston et al. [2020] and Ben-Michael et al. [2024a] consider a loss function that depends on the baseline potential outcome rather than the realized outcome. This is an example of the counterfactual loss $\ell(D;Y(0))$. Under this formulation, the decision is incorrect (i.e., a false negative) if the physician does nothing and the patient dies, $(D,Y(0))=(0,0)$. Since $Y(0)$ is realized in this case, the loss is $\ell_{0}$, as defined earlier. Conversely, a false positive occurs when the physician provides treatment even though the patient would have survived without it, $(D,Y(0))=(1,1)$. In this case, $Y(0)$ is counterfactual, and the loss is ~$\tilde{\ell}_{1}$. Under this setup, Ben-Michael et al. [2024a] propose the following counterfactual classification loss:
$\ell(D;Y(0))=(1-D)(1-Y(0))\ell_{0}+D\cdot Y(0)\tilde{\ell}_{1}+c_{D}.$ (2)
As illustrated in Table 1, this counterfactual loss can be further generalized. First, we may also incorporate loss terms for incorrect decisions based on the potential outcome under treatment — specifically, cases where $(D,Y(1))=(0,1)$ (false negative) or $(D,Y(1))=(1,0)$ (false positive). Second, we can include losses associated with correct decisions: $(D,Y(0))=(0,1)$ and $(D,Y(1))=(1,1)$ are true positives, while $(D,Y(0))=(0,1)$ and $(D,Y(1))=(0,0)$ are true negatives. Letting ~$\tilde{\ell}_{0}$ denote the loss incurred when a patient would have died under an alternative (unchosen) decision, we can write a general counterfactual loss as:
$\ell(D;Y(0),Y(1))=\ell_{Y(D)}+\tilde{\ell}_{Y(1-D)}+c_{D}.$ (3)
These examples illustrate that the counterfactual loss framework allows us to formalize aspects of decision-making that are difficult to capture using standard loss functions.
2.2 Asymmetric counterfactual loss
The counterfactual classification loss introduced above considers only one potential outcome at a time, $Y(d)$ for $d=0,1$. In contrast, a fully general counterfactual loss may depend on both potential outcomes at the same time. This leads naturally to the concept of principal strata, which classifies individuals according to their joint potential outcomes [Frangakis and Rubin, 2002]:
  • Never survivors $(Y(0),Y(1))=(0,0)$: individuals who would not survive regardless of treatment.
  • Responders $(Y(0),Y(1))=(0,1)$: individuals who would survive only if given the treatment.
  • Harmed $(Y(0),Y(1))=(1,0)$: individuals who would survive only if not treated.
  • Always survivors $(Y(0),Y(1))=(1,1)$: individuals who would survive regardless of treatment.
For any treatment decision $D$ and observed outcome $Y(D)$, the counterfactual loss can vary depending on the principal stratum to which the individual belongs.
Building on this framework, Ben-Michael et al. [2024b] propose a counterfactual loss function designed to encode the principle of the Hippocratic Oath — “Do no harm” — a longstanding ethical guideline in medicine [Wiens et al., 2019]. According to this principle, causing harm with a treatment is considered more serious than failing to provide a potentially life-saving intervention. This ethical asymmetry motivates the following loss function:
$\ell(D;Y(0),Y(1))$ $=(1-Y(0))Y(1)\ell_{D}^{\text{R}}+Y(0)(1-Y(1))\ell_{1-D}^{\text{H}}$ (4)
$\qquad+Y(0)Y(1)\ell_{1}+(1-Y(0))(1-Y(1))\ell_{0}+c_{D},$
where $\ell_{0}^{\text{R}}>\ell_{1}^{\text{R}}\geq 0$ and $\ell_{0}^{\text{H}}>\ell_{1}^{\text{H}}\geq 0$ represent the respective losses for incorrect and correct decisions within the responder and harmed strata, respectively.
The following inequality encodes the Hippocratic Oath:
$\underbrace{\Delta^{\text{R}}=\ell_{0}^{\text{R}}-\ell_{1}^{\text{R}}}_{\begin{subarray}{c}\text{Failure to treat}\\ \text{a responder}\end{subarray}}<\underbrace{\Delta^{\text{H}}=\ell_{0}^{\text{H}}-\ell_{1}^{\text{H}}}_{\begin{subarray}{c}\text{Harming}\\ \text{a patient}\end{subarray}}.$
That is, failing to provide life-saving treatment (a false negative among responders) is considered less costly than harming a patient by administering an unnecessary or detrimental treatment (a false positive among the harmed).
This counterfactual loss framework enables ethical principles to be explicitly incorporated into the treatment decision-making process. Moreover, the structure of principal strata and asymmetric loss can be extended to address other concerns, such as fairness in decisions [Imai and Jiang, 2023] or prioritizing treatment equity across subpopulations.
2.3 Trichotomous decision
Finally, we consider a slightly more general treatment choice setting by introducing an experimental treatment as a third option, denoted $D=2$. A physician may now choose among no treatment ($D=0$), standard treatment ($D=1$), and experimental treatment ($D=2$). Let $c_{2}$ represent the cost of the experimental treatment (relative to the cost of no treatment $c_{0}=0$), which we assume to be greater than the cost of standard treatment $c_{1}$. Let $\ell_{y}$ denote the loss associated with the observed outcome $Y(D)=y$ under decision $D$. Then, the standard loss function from Equation (1) takes again the form:
$\ell^{\textsc{Std}}(D,Y(D))=\ell_{Y(D)}+c_{D}.$
Suppose further that the experimental treatment is more invasive than the standard one, which itself is more invasive than giving no treatment. From an ethical or clinical perspective, a physician may prefer to avoid overtreatment by prescribing the least invasive option that ensures patient survival. A counterfactual loss function that encodes this preference is:
$\ell^{\textsc{Cof}}(D;Y(0),Y(1),Y(2))=\ell_{Y(D)}+c_{D}+\sum_{k<D}r_{k}Y(k),$ (5)
where $r_{k}$ represents the “regret” or penalty for administering a more invasive treatment when a less invasive one would have sufficed (i.e., when $Y(k)=1$ for some $k<D$).
When does the counterfactual loss recommend a different treatment than the standard loss? Let $\mu_{d}=\mathbb{E}[Y(d)]$ denote the average potential outcome under treatment $d$. Then, the difference between the counterfactual and standard risks for a given treatment $D=d$ can be written as:
$R(d,\ell^{\textsc{Cof}})-R(d,\ell^{\textsc{Std}})=\sum_{k<d}r_{k}\Pr(Y(k)=1\mid D=d)\Pr(D=d).$
To gain further intuition, we normalize the loss values such that $\ell_{0}=1$ and $\ell_{1}=0$, and again let $\mu_{d}=\mathbb{E}[Y(d)]$. Under this normalization, the counterfactual risk may favor a less invasive treatment $k<d$, even when the standard risk prefers $d$, if the regret from overtreatment is sufficiently large. Specifically, counterfactual risk favors $k$ over $d$ whenever:
$0<\frac{(\mu_{d}-c_{d})-(\mu_{k}-c_{k})}{\mu_{k}}<r_{k}.$
This expression highlights the trade-off between improved survival probabilities, treatment costs, and the ethical penalty for overtreatment. If the increase in cost and survival probability from $k$ to $d$ is small relative to the regret for overtreatment, the counterfactual risk will recommend a less invasive option. Thus, the counterfactual loss formally incorporates treatment conservatism and ethical preferences into decision-making.
3 Identification Analysis under Strong Ignorability
Despite its potential, the counterfactual loss framework faces a fundamental challenge: it is difficult to identify the counterfactual risk, which depends on the joint distribution of multiple potential outcomes — quantities that are never jointly observed. This raises the central question we address in the remainder of the paper: Under what conditions is the counterfactual risk identifiable? We will derive a necessary and sufficient condition that a counterfactual loss function must satisfy for its associated counterfactual risk to be identifiable under the standard assumption of strong ignorability, which posits that treatment decisions are unconfounded given observed variables.
3.1 Setup
Consider an independently and identically distributed (IID) sample of size $n$ drawn from a population. For each unit $i=1,\ldots,n$, we observe a vector of pre-treatment covariates $\bm{X}_{i}\in\mathcal{X}$, a discrete decision $D_{i}\in\mathcal{D}:=\{0,1,\ldots,K-1\}$ with $K\geq 2$, and an outcome $Y_{i}\in\mathcal{Y}:=\{0,1,\ldots,M-1\}$ with $M\geq 2$. Let $Y_{i}(d)\in\mathcal{Y}$ denote the potential outcome of unit $i$ under decision $d\in\mathcal{D}$.
Our goal is to evaluate the quality of a generic decision rule $D_{i}^{\ast}\in\mathcal{D}$, which may differ from the observed decision $D_{i}$. We assume that this target decision $D_{i}^{\ast}$ is observed for each unit and, when it shares the same potential outcomes as the observed decision, yields the same outcome. Formally, we adopt the following assumptions:
Assumption 1 (IID Sampling)
The tuple $\{Y_{i},D_{i},D_{i}^{\ast},\bm{X}_{i}\}$ is independently and identically distributed across $i=1,\ldots,n$.
Assumption 2 (Consistency)
For all $i$, the observed outcome satisfies $Y_{i}=Y_{i}(D_{i})$. Moreover, if $D_{i}^{\ast}=D_{i}$, then $Y_{i}(D_{i}^{\ast})=Y_{i}(D_{i})$.
In this setting, we do not require knowledge of how $D_{i}^{\ast}$ is generated, only that it is observed for each unit. For example, $D_{i}^{\ast}$ may correspond to an individualized treatment rule that depends on both observed and unobserved covariates: $D_{i}^{\ast}=\pi(\bm{X}_{i},\bm{U}_{i})$ for some possibly unknown function $\pi:\mathcal{X}\times\mathcal{U}\to\mathcal{D}$, where $\bm{U}_{i}\in\mathcal{U}$ denotes unobserved covariates. Alternatively, $D_{i}^{\ast}$ may arise from a stochastic policy, where $\Pr(D_{i}^{\ast}=d)=\pi_{d}(\bm{X}_{i},\bm{U}_{i})$ for $d\in\mathcal{D}$ and $\sum_{d=0}^{K-1}\pi_{d}(\bm{X}_{i},\bm{U}_{i})=1$. Importantly, we do not observe the corresponding potential outcome $Y_{i}(D_{i}^{\ast})$.
We study identifiability of counterfactual risk under standard assumptions from the causal inference literature:
Assumption 3 (Strong Ignorability)
For $i=1,\ldots,n$, the following assumptions hold,
  1. (a)
    Unconfoundedness: $D_{i}\mbox{$\perp\!\!\!\perp$}(D_{i}^{\ast},\{Y_{i}(d)\}_{d\in\mathcal{D}})\mid\bm{X}_{i}$.
  2. (b)
    Overlap: there exists $\eta>0$ such that, for all $d\in\mathcal{D}$, $\eta<\Pr(D_{i}=d\mid\bm{X}_{i})<1-\eta$.
Assumption 3(a) implies that, conditional on covariates, the observed decision $D_{i}$ is independent of both the potential outcomes and the decision rule $D_{i}^{\ast}$. Notably, $D_{i}^{\ast}$ may still depend on unobserved covariates, and thus on potential outcomes, in ways that do not violate this assumption. Assumption 3(b) requires that every treatment level in $\mathcal{D}$ occurs with nonzero probability given covariates, but only for the observed decision $D_{i}$, not for the hypothetical decision of interest $D_{i}^{\ast}$.
3.2 Counterfactual Loss and Risk
We define the loss (or negative utility) associated with a decision as a function of the covariates and all potential outcomes. A lower value of the loss indicates a better decision. We refer to this as the counterfactual loss because it depends on the full vector of potential outcomes rather than just the realized one.
Definition 1 (Counterfactual Loss)
Given potential outcomes $(Y_{i}(0),\ldots,Y_{i}(K-1))=(y_{0},\ldots,y_{K-1})$ and covariates $\bm{X}_{i}=\bm{x}$, the counterfactual loss of choosing decision $D^{\ast}=d$ is defined as a function $\ell:\mathcal{D}\times\mathcal{Y}^{K}\times\mathcal{X}\rightarrow\mathbb{R}$.
We now define the counterfactual risk, which is the expected counterfactual loss and serves as the primary estimand of interest. We also define the conditional counterfactual risk, which conditions on the observed covariates.
Definition 2 (Counterfactual Risk and Conditional Counterfactual Risk)
Given the counterfactual loss $\ell$, the counterfactual risk $R(D^{\ast};\ell)$ of decision $D^{\ast}$ is given by,
$R(D^{\ast};\ell):=\mathbb{E}{\left[\ell(D^{\ast};Y(0),\ldots,Y(K-1),\bm{X})\right]}=\mathbb{E}{\left[R_{\bm{X}}(D^{\ast};\ell)\right]},$
where $R_{\bm{x}}(D^{\ast};\ell)$ is the conditional counterfactual risk given $\bm{X}=\bm{x}$,
$R_{\bm{x}}(D^{\ast};\ell):=\sum_{d\in\mathcal{D}}\sum_{\{y_{k}\}_{k=0}^{K-1}\in\mathcal{Y}^{K}}\ell(d;y_{0},\ldots,y_{k-1},\bm{x})\cdot\Pr(D^{\ast}=d,Y(0)=y_{0},\ldots,Y(K-1)=y_{K-1}\mid\bm{X}=\bm{x}).$
When clear from context, we abbreviate the counterfactual risk as $R(D^{\ast})=R(D^{\ast};\ell)$ and the conditional counterfactual risk as $R_{\bm{x}}(D^{\ast})=R_{\bm{x}}(D^{\ast};\ell)$.
The central goal of this paper is to study the identifiability of counterfactual risk. Intuitively, the challenge arises from the fact that, for each unit, we observe only a single realized outcome out of $K$ possible potential outcomes. Although the counterfactual loss depends on the entire set of potential outcomes, we will show that the expected loss, i.e., the counterfactual risk, can be identified under a set of relatively mild structural conditions on the loss function.
3.3 Additivity
To overcome the identification problem, we consider the following restricted class of counterfactual loss that is additive in the potential outcomes.
Definition 3 (Additive Counterfactual Loss)
Let $\bm{y}=(y_{0},\ldots,y_{K-1})$. Then, additive counterfactual loss is defined as,
$\ell^{\textsc{Add}}(d;\bm{y},\bm{x}):=\sum_{k\in\mathcal{D}}\omega_{k}(d,y_{k},\bm{x})+\varpi(\bm{y},\bm{x}),$
where $\omega_{k}:\mathcal{D}\times\mathcal{Y}\times\mathcal{X}\rightarrow\mathbb{R}$ is the weight function and $\varpi:\mathcal{Y}^{K}\times\mathcal{X}\rightarrow\mathbb{R}$ is the intercept function.
The additive counterfactual loss equals the sum of two terms; the first term depends on the decision $D^{\ast}=d$, while the second term does not. Moreover, the first term is the sum of weight functions $\omega_{k}$, each depending only on the corresponding potential outcomes $Y_{i}(k)$ given the covariates. The second term does not depend on the decision and includes all possible interactions of potential outcomes.
We illustrate Definition 3 by revisiting the examples shown in Section 2.
Example 1 (Counterfactual Classification Loss)
The counterfactual loss function given in Equation (2) is additive in the potential outcomes with $\omega_{0}(0,0)=\ell_{0}$, $\omega_{0}(1,1)=\tilde{\ell}_{1}+c$, $\omega_{1}(d,y_{1})=0$ for $d,y_{1}\in\{0,1\}$, and $\varpi(\bm{y})=0$. The extension of this counterfactual loss given in Equation (3) is also additive with $\omega_{d}(d,y_{d})=\ell_{y_{d}}+c_{d}$, $\omega_{1-d}(d,y_{1-d})=\tilde{\ell}_{y_{1-d}}$ for $d,y\in\{0,1\}$, and $\varpi(\bm{y})=0$.
Example 2 (Asymmetric Counterfactual Loss)
The asymmetric counterfactual loss given in Equation (4) is an example of a non-additive loss whenever $\Delta^{\text{R}}\neq\Delta^{\text{H}}$ (see Section 3.6 for details). When $\Delta^{\text{R}}=\Delta^{\text{H}}=\Delta$ the loss is additive. The weights for this loss are given by $\omega_{d}(d,y_{d})=\ell_{y_{d}}+c_{d}$, $\omega_{1-d}(d,y_{1-d})=\{\Delta-(\ell_{0}-\ell_{1})\}y_{1-d}$, $\varpi(\bm{y})=(\ell_{1}^{\text{H}}-\ell_{1})y_{0}(1-y_{1})+(\ell_{1}^{\text{R}}-\ell_{1})(1-y_{0})y_{1}-\{\Delta-(\ell_{0}-\ell_{1})\}y_{0}y_{1}$. Note that $\Delta-(\ell_{0}-\ell_{1})$ is the difference (of the loss differences) between units for whom the treatment matters and those for whom it does not.
Example 3 (Trichotomous decision)
The counterfactual loss given in Equation (5) is additive in the potential outcomes with $\omega_{d}(d,y_{d})=\ell_{y_{d}}+c_{d}$, $\omega_{k}(d,y_{k})=r_{k}y_{k}\mathds{1}{\left\{k<d\right\}}$ for $d,k\in\{0,1,2\},y_{d},y_{k}\in\{0,1\}$, and $\varpi(\bm{y})=0$.
3.4 Nonparametric Identification
We now investigate conditions under which the counterfactual risk is identifiable, focusing on the class of additive counterfactual loss functions. Identification has been extensively studied in the causal inference literature [e.g., Matzkin, 2007, Pearl, 2009]. We adopt the following general definition of identification from Basse and Bojinov [2020]:
Definition 4 (Identifiability)
Let $\mathcal{P}$ denote the set of all joint probability distributions under consideration, and let $\mathcal{Q}$ be the set of all observable distributions. Let $Q:\mathcal{P}\to\mathcal{Q}$ denote a surjective mapping that associates each joint distribution with an observable distribution. Let $\theta(P)$ denote the (causal) estimand of interest under distribution $P\in\mathcal{P}$. For a fixed observable distribution $Q_{0}\in\mathcal{Q}$ define $\mathcal{S}_{0}=\{P\in\mathcal{P}:Q(P)=Q_{0}\}.$ We say that $\theta$ is identifiable at $Q_{0}$ with respect to $Q$ if there exists a constant $\theta_{0}$ such that $\theta(P)=\theta_{0}$ for all $P\in\mathcal{S}_{0}$. We say that $\theta$ is identifiable with respect to $Q$ if it is identifiable at every $Q_{0}\in\mathcal{Q}$.
According to the definition, a causal estimand is identifiable if it takes the same value across all joint probability measures that map to any given observable distribution. Surjectivity of $Q$ ensures that this property holds for every possible observable distribution.
In our setting, $\mathcal{P}$ denotes the set of all joint probability distributions $P(D^{\ast},D,Y(0),\ldots,Y(K-1),\bm{X})$ that satisfy Assumptions 13. The causal estimand of interest is the counterfactual risk $R(D^{\ast};\ell)$, which, according to Definition 2, is fully determined by $P\in\mathcal{P}$. Let $Q(P)=P(D^{\ast},D,Y,\bm{X})$ denote the surjective mapping that yields this observable distribution from each $P\in\mathcal{P}$. Then, according to Definition 4, the counterfactual risk $R(D^{\ast};\ell)$ is identifiable with respect to $Q$ if it can be expressed as a function of the observable distribution $P(D^{\ast},D,Y,\bm{X})$.
As formally shown in Lemma 2 (see Appendix G.2), since $P(\bm{X})$ is observable, the counterfactual risk $R(D^{\ast};\ell)$ is identifiable with respect to $P(D^{\ast},D,Y,\bm{X})$ if and only if the conditional counterfactual risk $R_{\bm{X}}(D^{\ast};\ell)$ is identifiable with respect to $P(D^{\ast},D,Y\mid\bm{X})$. Since $R_{\bm{X}}(D^{\ast};\ell)$ depends on the conditional distribution $P(D^{\ast},Y(0),\ldots,Y(K-1)\mid\bm{X})$, we need to express this distribution of potential outcomes in terms of observable distributions. Under Assumption 3, we can only identify the joint distribution of the decision $D^{\ast}$ and a single potential outcome given covariates,
$\Pr(D^{\ast}=d,Y(k)=y_{k}\mid\bm{X})\ =$ $\ \Pr(D^{\ast}=d,Y(k)=y_{k}\mid D=k,\bm{X})$
$=$ $\ \Pr(D^{\ast}=d,Y=y_{k}\mid D=k,\bm{X}),$
for any $k\in\mathcal{D}$. In particular, without additional assumptions, we cannot identify a joint distribution that involves multiple potential outcomes. Thus, it follows that $R_{\bm{x}}(D^{\ast};\ell)$ is identifiable if and only if it can be expressed solely in terms of the distribution $\Pr(D^{\ast}=d,Y(k)=y\mid\bm{X}=\bm{x})$ for any $d,k\in\mathcal{D},y\in\mathcal{Y}$. This observation suggests that the additivity assumption given in Definition 3 can facilitate identification.
We first show that under additive loss, we can identify the difference in conditional counterfactual risk between any two decision-making systems even though each conditional counterfactual risk is only identifiable up to an unknown additive constant. That is, the conditional counterfactual risk of decision-making system $D^{\ast}$ can be decomposed into an identifiable part and an unidentifiable part, the latter is a function of covariates and does not depend on $D^{\ast}$.
Theorem 1 (The Conditional Counterfactual Risk under the Additive Loss)
Suppose that a counterfactual loss function takes the additive form given in Definition 3. Then, under Assumptions 13, the conditional counterfactual risk can be decomposed as,
$R_{\bm{x}}(D^{\ast};\ell^{\textsc{Add}})=\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D}}\sum_{y\in\mathcal{Y}}\omega_{k}(d,y,\bm{x})\Pr(D^{*}=d,Y(k)=y\mid\bm{X}=\bm{x})+C(\bm{x}),$
where
$C(\bm{x})=\sum_{\bm{y}\in\mathcal{Y}^{K}}\varpi(\bm{y},\bm{x})\Pr(\bm{Y}(\mathcal{D})=\bm{y}\mid\bm{X}=\bm{x}),$
and $\bm{Y}(\mathcal{D})=(Y(0),\ldots,Y(K-1))$. Thus, the difference in conditional counterfactual risk between any pair of decision-making systems can be identified.
The proof is given in Appendix A.
Taking the expectation of the expression in Theorem 1 over the distribution of covariates, we find that the counterfactual risk is identified up to an unknown additive constant $\mathbb{E}[C(\bm{X})]$. We state this result as the following corollary without proof.
Corollary 1 (The Counterfactual Risk under the Additive Loss)
Suppose that a counterfactual loss function takes the additive form given in Definition 3. Then, under Assumptions 13, the counterfactual risk can be decomposed as,
$R(D^{\ast};\ell^{\textsc{Add}})=\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D}}\sum_{y\in\mathcal{Y}}\mathbb{E}\left[\omega_{k}(d,y,\bm{X})\Pr(D^{*}=d,Y=y\mid D=k,\bm{X})+C(\bm{X})\right].$
Thus, the difference in counterfactual risk between any pair of decision-making systems is identified.
Theorem 1 immediately implies that one can find an optimal decision, e.g., $\pi:\mathcal{X}\rightarrow\mathcal{D}$, by minimizing the identifiable term alone, i.e.,
$\pi^{\text{opt}}\ \in$ $\ \operatornamewithlimits{argmin}_{\pi\in\Pi}R(\pi(\bm{X});\ell^{\textsc{Add}})$
$=$ $\ \operatornamewithlimits{argmin}_{\pi\in\Pi}\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D}}\sum_{y\in\mathcal{Y}}\mathbb{E}\left[\omega_{k}(d,y,\bm{X})\Pr(\pi(\bm{X})=d,Y=y\mid D=k,\bm{X})\right],$
where $\Pi$ is a class of decision rules to be considered. Moreover, if the constant term is zero, then we can exactly identify the counterfactual risk. This happens when the second term in the additive counterfactual loss (Definition 3 is zero, i.e., $\varpi(\bm{y},\bm{x})=0$ for all $\bm{y}\in\mathcal{Y}^{K}$ and $\bm{x}\in\mathcal{X}$). Lastly, by placing assumptions on $\varpi$, we can control the magnitude of $\mathbb{E}{\left[C(X)\right]}$, which leads to partial identification results.
A natural question arises as to whether or not the above identification results can be obtained under a different form of counterfactual loss function. The next theorem shows that the additivity of the loss function given in Definition 3 is a necessary and sufficient condition for the identification of the difference in counterfactual risk. In other words, all counterfactual risks that are identifiable up to an additive constant must have an additive loss.
Theorem 2 (Additivity as Necessary and Sufficient Condition for Identification)
The difference in counterfactual risk between any pair of decision-making systems can be identified under Assumptions 13 if and only if the corresponding loss takes the additive form given in Definition 3.
The proof is given in Appendix B.
Finally, as a corollary of this identification result, we show that the exact identification of the counterfactual risk is possible if and only if the interaction term $\varpi$ is zero in the additive counterfactual loss (Definition 3).
Corollary 2 (Necessary and Sufficient Condition for Exact Identification)
The counterfactual risk given in Definition 2 can be exactly identified under Assumptions 13 if and only if the corresponding loss takes the following form,
~$\tilde{\ell}^{\textsc{Add}}(d;\bm{y},\bm{x})=\sum_{k\in\mathcal{D}}\omega_{k}(d,y_{k},\bm{x}),$
for a weight function $\omega_{k}:\mathcal{D}\times\mathcal{Y}\times\mathcal{X}\rightarrow\mathbb{R}$. The identification formula is given by,
$R(D^{*},\tilde{\ell}^{\textsc{Add}})$ $=\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D}}\sum_{y\in\mathcal{Y}}\mathbb{E}\left[\omega_{k}(d,y,\bm{X})\Pr(D^{\ast}=d,Y=y\mid D=k,\bm{X})\right].$
The proof is given in Appendix C.
3.5 Binary Outcome Case
To further build intuition for the above results, we study the special case of binary outcome, i.e., $\mathcal{Y}=\{0,1\}$. For notational simplicity, we ignore covariates. Without loss of generality, assume that $Y=1$ is a desirable outcome while $Y=0$ is undesirable. Under this setting, it is natural to consider the following ordering of weight functions in the additive counterfactual loss for any given decision $d$:
$\omega_{d}(d,1)\leq\{\omega_{d^{\prime}}(d,0)\}_{d^{\prime}\neq d}\leq 0\leq\{\omega_{d^{\prime}}(d,1)\}_{d^{\prime}\neq d}\leq\omega_{d}(d,0),$ (6)
where the terms $\omega_{d}(d,y_{d})$ for $y_{d}=0,1$ account for the accuracy of decision, and the other weights $\{\omega_{d^{\prime}}(d,y_{d^{\prime}})\}_{d^{\prime}\neq d}$ for $y_{d^{\prime}}=0,1$ capture its difficulty. The counterfactual loss is smaller when fewer alternative decisions would yield a desirable outcome (i.e., making a correct decision is difficult). The loss is greater when many alternatives would result in a desirable outcome (making a correct decision is easier). While accuracy is an essential part of standard decision theory, difficulty is unique to the counterfactual loss. In this way, the counterfactual loss rewards a consequential decision that changes the outcome.
We first discuss the accuracy component. The correct “true positive” decision, i.e., $(D^{\ast},Y(d))=(d,1)$ would decrease the counterfactual loss by the largest amount, which is represented by the largest negative loss of $\omega_{d}(d,1)$. In contrast, the incorrect “false negative” decision, i.e., $(D^{\ast},Y(d))=(d,0)$ would increase the loss by the greatest amount, which corresponds to a positive loss of $\omega_{d}(d,0)$.
While accuracy is an essential part of standard decision theory, difficulty is unique to the counterfactual loss. Avoidance of an incorrect decision (“true negative”), i.e., $(D^{\ast},Y(d))=(d^{\prime},0)$ in the top right cell results in a decrease of the loss, while failure to make a correct “true positive” decision, i.e., $(D^{\ast},Y(d))=(d^{\prime},1)$ increases the loss. Because these counterfactual outcomes do not occur, their weights are smaller in magnitude than those of the realized outcome. In this way, the counterfactual loss rewards a consequential decision that changes the outcome.
Lastly, since each weight $\omega_{k}(d,y)$ depends on both the decision and the outcome, we can decompose it into a baseline decision term and an interaction component:
$\omega_{k}(d,y)=\omega_{k}(d,0)+\mathds{1}{\left\{y=1\right\}}\left(\omega_{k}(d,1)-\omega_{k}(d,0)\right).$
The first term corresponds to the baseline cost of making decision $D^{\ast}=d$ on the potential outcome $Y(k)$, while the second term accounts for the interaction of decision and potential outcome.
We formalize the above discussion as the following corollary of Theorem 1. Specifically, we show that the conditional counterfactual risk can be decomposed into four components — accuracy based on the realized outcome $\Pr(D^{\ast}=d,Y(d)=1\mid\bm{X}=\bm{x})$, difficulty based on counterfactual outcomes $\Pr(D^{\ast}=d,Y(k)=1\mid\bm{X}=\bm{x})$ for $k\neq d$, and baseline decision probability $\Pr(D^{\ast}=d\mid\bm{X}=\bm{x})$, and the constant $C(\bm{x})$— with each term having weights based on the specified additive loss function.
Corollary 3 (Accuracy and Difficulty for the Binary Outcome Case)
Assume $Y\in\{0,1\}$. Suppose that the counterfactual loss function takes the additive form given in Definition 3. Then, under Assumptions 13, the conditional counterfactual risk can be written as,
$R_{\bm{x}}(D^{\ast};\ell^{\textsc{Add}})$
$=$ $\sum_{d\in\mathcal{D}}\zeta_{d}(d,\bm{x})\Pr(D^{\ast}=d,Y(d)=1\mid\bm{X}=\bm{x})+\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D},k\neq d}\zeta_{k}(d,\bm{x})\Pr(D^{\ast}=d,Y(k)=1\mid\bm{X}=\bm{x})$
$+\sum_{d\in\mathcal{D}}\xi(d,\bm{x})\Pr(D^{\ast}=d\mid\bm{X}=\bm{x})+C(\bm{x}).$
where $\zeta_{k}(d,\bm{x})=\omega_{k}(d,1,\bm{x})-\omega_{k}(d,0,\bm{x})$, $\xi(d,\bm{x})=\sum_{k\in\mathcal{D}}\omega_{k}(d,0,\bm{x})$, and
$C(\bm{x})=\sum_{\bm{y}\in\{0,1\}^{K}}\varpi(\bm{y},\bm{x})\Pr(\bm{Y}(\mathcal{D})=\bm{y}\mid\bm{X}=\bm{x}).$
The proof is given in Appendix F. If we choose the weights such that Equation (6) holds, then we have $\zeta_{d}(d,\bm{x})\leq 0\leq\zeta_{k}(d,\bm{x})$ for all $k\neq d$. This implies that the risk decreases with greater accuracy, i.e., when the realized outcome is desirable, and increases with counterfactual regret, i.e., when other decisions also would yield a desirable outcome, reflecting the difficulty of the decision. The baseline decision term is a regularizer that penalizes decisions with higher baseline losses.
3.6 Connection to Principal Strata
Principal Strata Decision
$D^{\ast}=0$ $D^{\ast}=1$
Never Survivor
$(Y(0),Y(1))=(0,0)$
$\ell(0;0,0)$ $\ell(1;0,0)$
Harmed
$(Y(0),Y(1))=(1,0)$
$\ell(0;1,0)$ $\ell(1;1,0)$
Responsive
$(Y(0),Y(1))=(0,1)$
$\ell(0;0,1)$ $\ell(1;0,1)$
Always Survivor
$(Y(0),Y(1))=(1,1)$
$\ell(0;1,1)$ $\ell(1;1,1)$
Table 2: Counterfactual Loss for Principal Strata in the Binary Case $\ell(D^{\ast};Y(0),Y(1))$.
Next, we consider the relationship between the principal strata and the weights of an additive counterfactual loss. For simplicity, we further restrict our attention to the case of binary decision and binary outcome, though the results can be generalized.
In this setting, as shown in Table 2, there are a total of eight different values the loss function can take, i.e., $\ell(d;y_{0},y_{1})$ for $d,y_{0},y_{1}\in\{0,1\}$, corresponding to two decisions for each of four principal strata. We can place different restrictions on the loss function to satisfy additivity.
Standard statistical decision theory assumes that the loss depends only on choice and its realized outcome. This implies that when $D^{\ast}=0$, never survivors and responsive units, for whom we have $Y(0)=0$, incur the same loss. Similarly, always survivors and the harmed units, for whom we have $Y(0)=1$, yield the same loss. When $D^{\ast}=1$, never survivors and the harmed are assumed to have the same loss, as do always survivors and responsive units. Together, standard statistical decision theory imposes a total of four equality constraints,
$\ell(0;0,0)=\ell(0;0,1),\quad\ell(0;1,0)=\ell(0;1,1),\quad\ell(1;0,0)=\ell(1;1,0),\quad\ell(1;0,1)=\ell(1;1,1).$ (7)
In Table 2, these constraints imply that within each column, the light orange and purple cells equal one another, while the dark orange and purple cells are equal to each other. In this case, four out of the twelve possible weights are non-zero and take the form,
$\omega_{0}(0,0)=\ell(0;0,y_{1}),\quad\omega_{0}(0,1)=\ell(0;1,y_{1}),\quad\omega_{1}(1,0)=\ell(1;y_{0},0),\quad\omega_{1}(1,1)=\ell(1;y_{0},1),$
for $y_{0},y_{1}\in\{0,1\}$.
While the above set of constraints leads to the exact identification of risk, we can further reduce the number of constraints without losing the exact identification. For example, we can assume that for each decision $d\in\{0,1\}$, the sum of losses for never survivors and always survivors equals the sum of losses for the harmed and responsive,
$\ell(d;0,0)+\ell(d;1,1)=\ell(d;0,1)+\ell(d;1,0).$ (8)
This represents a necessary and sufficient condition for exact identification of the counterfactual risk under strong ignorability. In Table 2, this implies that the sum of orange cells equals the sum of purple cells within each column. Given these two equality restrictions, if we let $\omega_{0}(1,1)=a$ and $\omega_{1}(0,1)=b$, we can express the remaining six weights as follows,
$\omega_{0}(0,0)=\ell(0;0,1)-b,\quad\omega_{0}(0,1)=\ell(0;1,1)-b,$ $\quad\omega_{1}(1,0)=\ell(1;1,0)-a,\quad\omega_{1}(1,1)=\ell(1;1,1)-a,$
$\omega_{0}(1,0)=\ell(1;0,1)-\ell(1;1,1)+a,$ $\quad\omega_{1}(0,0)=\ell(0;1,0)-\ell(0;1,1)+b.$
We can further relax the equality constraints and achieve identification up to an unknown constant. We may assume that the sum of loss differences between the two decisions for never-survivors and always-survivors is equal to the corresponding sum for the responsive and harmed units. This restriction represents a necessary and sufficient condition for identification up to an unknown constant and can be written as,
$[\ell(1;0,0)-\ell(0;0,0)]+[\ell(1;1,1)-\ell(0;1,1)]=[\ell(1;1,0)-\ell(0;1,0)]+[\ell(1;0,1)-\ell(0;0,1)].$ (9)
In Table 2, this equality constraint implies that the sum of orange cells is equal to the sum of purple cells across the two columns.
Restrictions Degrees Strong
on loss functions of freedom ignorability
light (dark) orange = light (dark) purple
within each column (Equation (7))
4 Exact
sum of orange $=$ sum of purple
within each column (Equation (8))
6 Exact*
sum of orange = sum of purple
across two columns (Equation (9))
7 Constant*
No restriction 8 Unidentifiable
Table 3: The table summarizes the identifiability restrictions under which the risk is either exactly identifiable (“Exact” in green), identifiable up to a constant (“Constant” in yellow), or unidentifiable (in red), under strong ignorability. An asterisk * signifies a necessary and sufficient condition. The second column shows the number of free parameters in the loss specification.
To examine identifiability and obtain weights under this and more general settings with multi-valued decisions and outcomes, we consider the following linear system,
$\bm{A}\bm{w}=\bm{\ell},$ (10)
where =$\bm{\ell}=\{\ell(d,y_{0},\ldots,y_{K-1})\}_{d\in\mathcal{D},\bm{y}\in\mathcal{Y}^{K}}$ is the vector of all possible loss values, $\bm{w}=\{\omega_{k}(d,y),\varpi(\bm{y})\}_{d\in\mathcal{D},k\in\mathcal{D},y\in\mathcal{Y},\bm{y}\in\mathcal{Y}^{K}}$ is the vector of all weights and intercept terms, and $\bm{A}$ is a ||×$\left|\bm{\ell}\right|\times\left|\bm{w}\right|$ matrix of zeros and ones, defined by the additive structure given in Definition 3. Then, the image of $\bm{A}$, denoted by $\text{im}(\bm{A})=\text{ker}(\bm{a}^{\top})^{\perp}$ [Lax, 2007], characterizes the space of counterfactual loss functions that is additive. If $\bm{\ell}\in\text{im}(\bm{A})$, then the loss is identifiable and there exists a solution $\bm{w}$ satisfying Equation (10). Otherwise, $\bm{\ell}\notin\text{im}(\bm{A})$ and the loss is not identifiable.
For exact identification, we consider a restricted version of Equation (10), $\widetilde{\bm{A}}\tilde{\bm{w}}=\tilde{\bm{\ell}}$ where $\tilde{\bm{w}}=\{\omega_{k}(d,y)\}_{d\in\mathcal{D},k\in\mathcal{D},y\in\mathcal{Y}}$ contains only the weights, and the matrix $\widetilde{\bm{A}}$ is restricted accordingly. An explicit example of this construction is provided in Appendix H. Solving this linear equation with the restriction given in Equation (9) gives the following weights,
$\omega_{0}(0,0)=\ell(0;0,1)-b-c,$ $\quad\omega_{0}(0,1)=\ell(0;1,1)-b-e,$
$\omega_{1}(1,0)=\ell(1;1,0)-a-d,$ $\quad\omega_{1}(1,1)=\ell(1;1,1)-a-e,$
$\omega_{0}(1,0)=\ell(1;0,1)-\ell(1;1,1)+a-c+e,$ $\quad\omega_{1}(0,0)=\ell(0;1,0)-\ell(0;1,1)+b-d+e,$
$\varpi_{0,1}(0,0)=\ell(1;0,0)-\ell(1;1,0)-$ $\ell(1;0,1)+\ell(1;1,1)+c+d-e.$
where $\omega_{0}(1,1)=a,\omega_{1}(0,1)=b,\varpi_{0,1}(0,1)=c,\varpi_{0,1}(1,0)=d,\varpi_{0,1}(1,1)=e$ are free parameters. The third column of Table 3 summarizes the identification results discussed above.
3.7 When is the Additive Counterfactual Loss Different from the Standard Loss?
Although the additive counterfactual loss given in Definition 1 generalizes the additive standard loss, one may wonder when they yield different treatment recommendations. We first show that, when the decision is binary, i.e., $K=2$, for any given additive counterfactual loss, we can always find a standard loss that gives the same treatment recommendations. However, when the decision is non-binary, the additive counterfactual loss no longer generates the same treatment recommendations as those based on the standard loss.
Proposition 1 (Additive Counterfactual Risk with Binary Decision)
Suppose that the decision is binary, i.e., $\mathcal{D}=\{0,1\}$.
  1. 1.
    For any additive counterfactual loss $\ell^{\textsc{Add}}(d;\bm{y})$ defined in Definition 3, we can construct a standard loss $\ell^{\textsc{Std}}(d,y_{d})$ such that the risk difference $R(D^{\ast};\ell^{\textsc{Add}})-R(D^{\ast};\ell^{\textsc{Std}})$ does not depend on $D^{\ast}$. Moreover, this standard loss is unique up to a constant only depending on covariates and takes the following form,
    $\ell^{\textsc{Std}}(d,y_{d},\bm{x})=\omega_{d}(d,y_{d},\bm{x})-\omega_{d}(1-d,y_{d},\bm{x})+c(\bm{x}).$
  2. 2.
    While the mapping from the space of counterfactual losses to standard losses that yield the same treatment recommendations is surjective, it is not injective. That is, for any standard loss, there exists an infinite number of counterfactual losses that give the same treatment recommendations.
The proof is given in Appendix D. This proposition implies that in the binary decision setting, any additive counterfactual risk admits a standard risk that yields identical treatment recommendations. As a consequence, both of these risks lead to the same optimal policy. Perhaps surprisingly, this result is not limited to binary outcomes. In addition, the corresponding standard loss $\ell^{\textsc{Std}}(d,y_{d})$ is given by (up to an additive constant only depending on covariates) the difference between the weight for the realized outcome $Y(D)=y_{d}$ under the chosen decision $D^{\ast}=d$ and the weight for the same outcome $Y(D)=y_{d}$ under the alternative decision $D^{\ast}=1-d$.
Although every additive counterfactual loss has a unique corresponding standard loss (up to an additive constant) with the same treatment recommendations, the reverse does not hold. That is, any given standard loss has infinitely many additive counterfactual losses with the same treatment recommendations. Each of such counterfactual losses assigns different values to principal strata. This asymmetry explains why, even in the binary case, additive counterfactual loss has a straightforward interpretation based on principal strata whereas standard loss does not.
We now revisit our binary decision examples given in Section 3.3.
Example 1 (Counterfactual Classification Loss)
The counterfactual loss given in Equation (3) can be written as the following standard loss $\ell^{\textsc{Std}}(d,y_{d})=\ell_{y_{d}}+c_{d}-\tilde{\ell}_{y_{d}}=\ell_{y_{d}}-\tilde{\ell}_{y_{d}}+c_{d}$, which corresponds to the difference between the realized and counterfactual loss associated with outcome $y_{d}$ with the additional cost term.
Example 2 (Asymmetric Counterfactual Loss)
When $\Delta^{\text{R}}=\Delta^{\text{H}}=\Delta$, the loss in Equation (4) can be rewritten as a standard loss with $\ell^{\textsc{Std}}(d,y_{d})=\ell_{y_{d}}+c_{d}-\{\Delta-(\ell_{0}-\ell_{1})\}y_{d}=(\ell_{1}^{\text{H}}-\ell_{0}^{\text{H}})y_{d}+c_{d}+\ell_{0}=(\ell_{1}^{\text{R}}-\ell_{0}^{\text{R}})y_{d}+c_{d}+\ell_{0}$. This is the loss difference of a positive and negative outcome for units for which the treatment matters when the treatment is applied. Note that the loss $\ell_{y}$ associated with never-and always-survivors does not impact the treatment choice.
Finally, we show that in general when the decision is non-binary, no standard loss corresponds directly to the additive counterfactual loss.
Proposition 2 (Additive Counterfactual Risk with Non-binary Decision)
Assume that the decision is non-binary, i.e., $K=|\mathcal{D}|\geq 3$. Then, for any additive counterfactual loss defined in Definition 3 with at least one counterfactual weight $\omega_{k}(d,y_{k},\bm{x})$ depending on the decision $d\in\mathcal{D}$ and potential outcome $y_{k}\in\mathcal{Y}$ for $d\neq k$, there exists no standard loss $\ell^{\textsc{Std}}(d;y_{d})$ such that the risk difference $R(D^{\ast};\ell^{\textsc{Add}})-R(D^{\ast};\ell^{\textsc{Std}})$ does not depend on $D^{\ast}$.
The proof is given in Appendix E. Proposition 2 shows that in the non-binary setting, no standard loss can replicate the decision behavior induced by a genuinely counterfactual additive loss. This demonstrates that standard losses are inadequate when decision-counterfactual interactions matter.
4 Concluding Remarks
In this paper, we generalize the classical statistical decision theory for treatment choice by introducing counterfactual losses that depend on all potential outcomes. While the standard loss evaluates decisions based solely on observed outcomes, the counterfactual loss also considers what would have happened under alternative actions. This enables us to assess not only the accuracy of a decision but also its difficulty, that is, whether the chosen action truly mattered or whether other choices would have led to similar results.
A central challenge in this framework is identification: since only one potential outcome is observed per unit, counterfactual risks are generally unidentifiable. We show that under the standard assumption of strong ignorability, a counterfactual loss leads to an identifiable risk if and only if it is additive in the potential outcomes. Moreover, we establish that such additive counterfactual losses do not have an equivalent standard loss representation so long as decisions are non-binary.
Counterfactual loss expands the scope of statistical decision theory by allowing analysts to incorporate ethical and practical considerations that are especially relevant in domains such as medicine and public policy. Promising directions for future work include extending the framework to continuous outcomes and relaxing the strong ignorability assumption in the identification analysis [see Ben-Michael et al., 2024a].
Supplementary Appendix
Appendix A Proof of Theorem 1
By the definition of the counterfactual conditional risk (Definition 2), we have,
$R_{\bm{x}}(D^{\ast};\ell^{\textsc{Add}})$
=\ = $\sum_{d\in\mathcal{D}}\sum_{\bm{y}\in\mathcal{Y}^{K}}\left(\sum_{k\in\mathcal{D}}\omega_{k}(d,y_{k},\bm{x})+\varpi(\bm{y},\bm{x})\right)$
×$\hskip 72.26999pt\times\Pr(D^{\ast}=d,Y(0)=y_{1},\ldots,Y(K-1)=y_{K-1}\mid\bm{X}=\bm{x})$
=\ = $\sum_{d\in\mathcal{D}}\sum_{\bm{y}\in\mathcal{Y}^{K}}\sum_{k\in\mathcal{D}}\omega_{k}(d,y_{k},\bm{x})\Pr(D^{\ast}=d,Y(0)=y_{0},\ldots,Y(K-1)=y_{K-1}\mid\bm{X}=\bm{x})$
$\quad+\sum_{d\in\mathcal{D}}\sum_{\bm{y}\in\mathcal{Y}^{K}}\varpi(\bm{y},\bm{x})\Pr(D^{\ast}=d,Y(0)=y_{0},\ldots,Y(K-1)=y_{K-1}\mid\bm{X}=\bm{x}),$ (A.1)
where the first equality follows from Definition 3. We now consider each of the two summation terms in Equation (A.1). For notational convenience, we write $\bm{y}_{-k}=(y_{0},\ldots,y_{k-1},y_{k+1},\ldots,y_{K-1})$ for any $k\in\mathcal{D}$. We can rewrite the first term as,
$\sum_{d\in\mathcal{D}}\sum_{\bm{y}\in\mathcal{Y}^{K}}\sum_{k\in\mathcal{D}}\omega_{k}(d,y_{k},\bm{x})\Pr(D^{\ast}=d,Y(0)=y_{0},\ldots,Y(K-1)=y_{K-1}\mid\bm{X}=\bm{x})$
=\ = $\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D}}\sum_{y_{k}\in\mathcal{Y}}\sum_{\bm{y}_{-k}\in\mathcal{Y}^{K-1}}\omega_{k}(d,y_{k},\bm{x})\Pr(D^{\ast}=d,Y(0)=y_{0},\ldots,Y(K-1)=y_{K-1}\mid\bm{X}=\bm{x})$
=\ = $\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D}}\sum_{y_{k}\in\mathcal{Y}}\omega_{k}(d,y_{k},\bm{x})\sum_{\bm{y}_{-k}\in\mathcal{Y}^{K-1}}\Pr(D^{\ast}=d,Y(0)=y_{0},\ldots,Y(K-1)=y_{K-1}\mid\bm{X}=\bm{x})$
=\ = $\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D}}\sum_{y_{k}\in\mathcal{Y}}\omega_{k}(d,y_{k},\bm{x})\Pr(D^{\ast}=d,Y(k)=y_{k}\mid\bm{X}=\bm{x})$
=\ = $\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D}}\sum_{y\in\mathcal{Y}}\omega_{k}(d,y,\bm{x})\Pr(D^{\ast}=d,Y(k)=y\mid\bm{X}=\bm{x})$
=\ = $\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D}}\sum_{y\in\mathcal{Y}}\omega_{k}(d,y,\bm{x})\Pr(D^{\ast}=d,Y=y\mid D=k,\bm{X}=\bm{x}),$
where the last equality follows from Assumptions 2and 3.
We next consider the second summation in Equation (A.1).
$\sum_{d\in\mathcal{D}}\sum_{\bm{y}\in\mathcal{Y}^{K}}\varpi(\bm{y},\bm{x})\Pr(D^{\ast}=d,Y(0)=y_{0},\ldots,Y(K-1)=y_{K-1}\mid\bm{X}=\bm{x})$
=\ = $\sum_{\bm{y}\in\mathcal{Y}^{K}}\varpi(\bm{y},\bm{x})\sum_{d\in\mathcal{D}}\Pr(D^{\ast}=d,Y(0)=y_{0},\ldots,Y(K-1)=y_{K-1}\mid\bm{X}=\bm{x})$
=\ = $\sum_{\bm{y}\in\mathcal{Y}^{K}}\varpi(\bm{y},\bm{x})\Pr(Y(0)=y_{0},\ldots,Y(K-1)=y_{K-1}\mid\bm{X}=\bm{x}).$
This completes the proof. ∎
Appendix B Proof of Theorem 2
Theorem 1 immediately establishes the sufficiency by Definition 4, as the counterfactual risk is a function only of the observable marginals (up to an unknown additive constant). We now prove the necessity. That is, we will show that if the counterfactual risk is identifiable up to an unknown additive constant, then the corresponding loss function must take the additive form given in Definition 3.
By Lemma 2, the counterfactual risk is identifiable if and only if the conditional counterfactual risk is identifiable. It therefore suffices to show that if the conditional counterfactual risk is identifiable up to an unknown additive constant depending only on covariates, then the corresponding loss must be additive.
We begin by fixing $\bm{X}=\bm{x}$. Let $\bm{p}(\bm{x}):=(p_{1}(\bm{x}),\ldots,p_{N}(\bm{x}))^{\top}\in\Delta_{N-1}$ denote the (conditional) probability measure over $\{D^{\ast},Y(0),\ldots,Y(K-1)\}$ given $\bm{X}=\bm{x}$ where $N=K\times M^{K}$ and $\Delta_{N-1}$ is the probability simplex defined as,
$\Delta_{N-1}=\left\{\bm{p}\in\mathbb{R}^{N}\ \mathrel{\Big{|}}\ \sum_{i=1}^{N}p_{i}=1\text{ and }p_{i}\geq 0\text{ for }i=1,\dots,N\right\}.$
Thus, each element of $\bm{p}(\bm{x})$ is indexed by $i=i(d,\bm{y})$ and is equal to $p_{i}=\Pr(D^{\ast}=d,Y(0)=y_{0},\ldots,Y(K-1)=y_{K-1}\mid\bm{X}=\bm{x})$. Then, by Definition 2, we can write the conditional counterfactual risk as the following linear function of $\bm{p}(\bm{x})$,
$R_{\bm{x}}(D^{\ast};\ell)$ $=\bm{\ell}^{\top}\bm{p}(\bm{x}),$
where $\bm{\ell}=(\ell_{1},\ldots,\ell_{N})^{\top}$ with $\ell_{i}=\ell(d,\bm{y},\bm{x})$ for $i=i(d,\bm{y})$.
By Definition 4 and the subsequent discussion in Section 3.4 an identifiable risk can be written entirely as a function of the observable marginals $\Pr(D^{\ast}=d,Y(k)=y_{k}\mid\bm{X}=\bm{x})$. Since we allow the risk to be identifiable up to a constant, it can also depend on the joint distribution of potential outcomes that do not involve the decision $D^{\ast}$, i.e., $\Pr(\bm{Y}(\mathcal{D})=\bm{y}\mid\bm{X}=\bm{x})$.
Let $Q:\Delta_{N-1}\to[0,1]^{L}$ be the observational mapping from Definition 4 with output $Q(\bm{p}(\bm{x}))=\bm{q}(\bm{x})$ the $L$-dimensional stacked vector of observed distribution $\{\Pr(D^{\ast}=d,Y(k)=y_{k}\mid\bm{X}=\bm{x})\}_{\{k\in\mathcal{D},d\in\mathcal{D},y_{k}\in\mathcal{Y}\}}$ and unobserved one $\{\Pr(\bm{Y}(\mathcal{D})=\bm{y}\mid\bm{X}=\bm{x})\}_{\bm{y}\in\mathcal{Y}^{K}}$ with $L=K^{2}\times M+M^{K}$. By Lemma 3(b) we know that $Q$ is surjective to the space of marginals and
$Q(\bm{p}(\bm{x}))=\bm{C}\bm{p}(\bm{x})=\bm{q}(\bm{p}(\bm{x})),$
for every $\bm{p}(\bm{x})\in\Delta_{N-1}$ and corresponding marginal $\bm{q}(\bm{p}(\bm{x}))$, where $\bm{C}$ is a $L\times N$ matrix of zeros and ones. Its kernel satisfies $\text{ker}(\bm{C})=\{\bm{v}\in\mathbb{R}^{N}:\bm{C}\bm{v}=\mathbf{0}_{L},\mathbf{1}_{N}^{\top}\bm{v}=0\}$ where $\mathbf{0}_{L}$ and $\mathbf{1}_{N}$ are an $L$-dimensional vector of zeros and an $N$-dimensional vector of ones, respectively.
Because the conditional risk $R_{\bm{x}}(D^{\ast};\ell)$ is identifiable up to an additive constant, Definition 4 combined with Lemma 1 guarantees the existence of a map $f$ such that
$R_{\bm{x}}(D^{\ast};\ell)=f(\bm{q}(\bm{p}(\bm{x}))),$ (B.1)
for all $\bm{p}(\bm{x})\in\Delta_{N-1}$.
Equation (B.1) differs from our discussion in Section 3.4, in which identifiability is defined with respect to the conditional probability measures over $\{D,D^{\ast},Y(0),\ldots,Y(K-1)\}$ given $\bm{X}=\bm{x}$ that satisfy Assumptions 13. In contrast, Equation (B.1) considers all conditional probability measures over $\{D^{\ast},Y(0),\ldots,Y(K-1)\}$ given $\bm{X}=\bm{x}$. This difference does not affect our conclusion for two reasons. First, the latter distribution fully characterizes the conditional counterfactual risk $R_{\bm{x}}(D^{\ast};\ell)$. Second, Assumptions 13 do not place any restriction on the distribution of $\{D^{\ast},Y(0),\ldots,Y(K-1)\}$ given $\bm{X}=\bm{x}$. Thus, identifiability must hold uniformly over every possible observable collection of the distribution $\{D^{\ast},Y(0),\ldots,Y(K-1)\}$ given $\bm{X}=\bm{x}$, without any restriction. Thus, Equation (B.1) holds for every $\bm{p}(\bm{x})\in\Delta_{N-1}$.
Finally, Lemma 4 implies that there exists a vector $\bm{w}\in\mathbb{R}^{L}$ such that
$R_{\bm{x}}(D;\ell)$ $=\bm{w}^{\top}\bm{q}(\bm{p}(\bm{x}))$
$=\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D}}\sum_{y\in\mathcal{Y}}w_{k}(d,y,\bm{x})\Pr(D^{\ast}(z)=d,Y(k)=y\mid\bm{X}=\bm{x})$
$\quad+\sum_{\bm{y}\in\mathcal{Y}^{K}}\varpi(\bm{y},\bm{x})\Pr(\bm{Y}(\mathcal{D})=\bm{y}_{\mathcal{S}}\mid\bm{X}=\bm{x})$
where $w_{k}(d,y,\bm{x})$ and $\varpi(\bm{y},\bm{x})$ correspond to different entries of $\bm{w}$. We must show that the additive counterfactual loss given in Definition 3 is the only representation that leads to this conditional counterfactual risk. Assume that there exists another loss function representation $\tilde{\bm{\ell}}$ such that
$R_{\bm{x}}(D;\ell)=\bm{\ell}^{\top}\bm{p}(\bm{x})=\tilde{\bm{\ell}}^{\top}\bm{p}(\bm{x}).$
Since this relationship must hold for any $\bm{p}(\bm{x})\in\Delta_{N-1}$, we can choose $\bm{p}(\bm{x})=\bm{e}(i)\in\mathbb{R}^{N}$, which denotes the standard basis vector with zero entries except the $i$th entry equaling one. This yields $\bm{\ell}_{i}=\tilde{\bm{\ell}}_{i}$, for all $i\in\{1,\ldots,N\}$ and thus $\bm{\ell}=\tilde{\bm{\ell}}$. This completes the proof. ∎
Appendix C Proof of Corollary 2
The result follows immediately from the proof of Theorem 2 by restricting $\bm{q}(\bm{p}(\bm{x}))$ to the probabilities $\{\Pr(D^{\ast}=d,Y(k)=y_{k}\mid\bm{X}=\bm{x})\}_{d\in\mathcal{D},k\in\mathcal{D},y_{k}\in\mathcal{Y}}$, while omitting the other terms involving multiple potential outcomes. Then, instead of Lemma 3(b), we apply Lemma 3(a).∎
Appendix D Proof of Proposition 1
We first show the existence of such a standard loss. When $\mathcal{D}=\{0,1\}$, Theorem 1 implies,
$R_{\bm{x}}(D^{\ast};\ell^{\textsc{Add}})$
=\ = $\sum_{d\in\{0,1\}}\sum_{k\in\{0,1\}}\sum_{y\in\mathcal{Y}}\omega_{k}(d,y,\bm{x})\Pr(D^{*}=d,Y(k)=y\mid\bm{X}=\bm{x})+C(\bm{x})$
=\ = $\sum_{y\in\mathcal{Y}}\left\{\omega_{0}(0,y,\bm{x})\Pr(D^{*}=0,Y(0)=y\mid\bm{X}=\bm{x})+\omega_{1}(1,y,\bm{x})\Pr(D^{*}=1,Y(1)=y\mid\bm{X}=\bm{x})\right.$
$\left.\qquad+\omega_{0}(1,y,\bm{x})\Pr(D^{*}=1,Y(0)=y\mid\bm{X}=\bm{x})+\omega_{1}(0,y,\bm{x})\Pr(D^{*}=0,Y(1)=y\mid\bm{X}=\bm{x})\right\}+C(\bm{x})$
where $C(\bm{x})=\sum_{\bm{y}\in\mathcal{Y}^{K}}\varpi(\bm{y},\bm{x})\Pr(\bm{Y}(\mathcal{D})=\bm{y}\mid\bm{X}=\bm{x})$. Observe,
$\Pr(D^{*}=d,Y(1-d)=y\mid\bm{X}=\bm{x})$
$=$ $\Pr(Y(1-d)=y\mid\bm{X}=\bm{x})-\Pr(D^{*}=1-d,Y(1-d)=y\mid\bm{X}=\bm{x}),$
for $d=0,1$. Substituting this into the above equation yields,
$R_{\bm{x}}(D^{\ast};\ell^{\textsc{Add}})$
$=$ $\sum_{y\in\mathcal{Y}}\omega_{0}(0,y,\bm{x})\left\{\Pr(D^{*}=0,Y(0)=y\mid\bm{X}=\bm{x})+\omega_{1}(1,y,\bm{x})\Pr(D^{*}=1,Y(1)=y\mid\bm{X}=\bm{x})\right.$
$\quad+\omega_{0}(1,y,\bm{x})\left(\Pr(Y(0)=y\mid\bm{X}=\bm{x})-\Pr(D^{*}=0,Y(0)=y\mid\bm{X}=\bm{x})\right)$
$\left.\quad+\omega_{1}(0,y,\bm{x})\left(\Pr(Y(1)=y\mid\bm{X}=\bm{x})-\Pr(D^{*}=1,Y(1)=y\mid\bm{X}=\bm{x})\right)\right\}+C(\bm{x})$
$=$ $\sum_{y\in\mathcal{Y}}\left(\omega_{0}(0,y,\bm{x})-\omega_{0}(1,y,\bm{x})\right)\Pr(D^{*}=0,Y(0)=y\mid\bm{X}=\bm{x})$
$\quad+\left(\omega_{1}(1,y,\bm{x})-\omega_{1}(0,y,\bm{x})\right)\Pr(D^{*}=1,Y(1)=y\mid\bm{X}=\bm{x})+C^{\prime}(\bm{x})$
$=$ $\sum_{d\in\mathcal{D}}\sum_{y\in\mathcal{Y}}\ell^{\textsc{Std}}(d,y,\bm{x})\Pr(D^{*}=d,Y(d)=y\mid\bm{X}=\bm{x})+C^{\prime}(\bm{x})$
$=$ $R_{\bm{x}}(D^{\ast};\ell^{\textsc{Std}})+C^{\prime}(\bm{x}),$
where $\ell^{\textsc{Std}}(d,y,\bm{x})=\omega_{d}(d,y,\bm{x})-\omega_{d}(1-d,y,\bm{x})$ and
$C^{\prime}(\bm{x})=$ $\sum_{y\in\mathcal{Y}}\omega_{0}(1,y,\bm{x})\Pr(Y(0)=y\mid\bm{X}=\bm{x})+\omega_{1}(0,y,\bm{x})\Pr(Y(1)=y\mid\bm{X}=\bm{x})$
$\quad+\sum_{\bm{y}\in\mathcal{Y}^{K}}\varpi(\bm{y},\bm{x})\Pr(\bm{Y}(\mathcal{D})=\bm{y}\mid\bm{X}=\bm{x}).$
Thus,
$R_{\bm{x}}(D^{\ast};\ell^{\textsc{Add}})-R_{\bm{x}}(D^{\ast};\ell^{\textsc{Std}})$ $=C^{\prime}(\bm{x}),$
which does not depend on $D^{\ast}$. Taking the expectation on both sides yields the existence of the desired standard loss. This completes the proof of the first part of the proposition.
Next, we prove the uniqueness. Assume there exists another standard loss ~$\tilde{\ell}^{\textsc{Std}}$ such that $R(D^{\ast},\ell^{\textsc{Add}})-R(D^{\ast},\tilde{\ell}^{\textsc{Std}})=\tilde{c}$ where $\tilde{c}$ is a constant. Then, this assumption implies that the conditional risk difference is given by
$R_{\bm{x}}(D^{\ast};\ell^{\textsc{Std}})-R_{\bm{x}}(D^{\ast};\tilde{\ell}^{\textsc{Std}})$ $=\widetilde{C}(\bm{x}),$ (D.1)
where $\widetilde{C}(\bm{x})$ is a function that satisfies $\tilde{c}=\mathbb{E}[\widetilde{C}(\bm{X})]$ (otherwise, for any $D^{\ast}$ independent of $\bm{X}$, the resulting risk difference would depend on $D^{\ast}$, contradicting our assumption).
By choosing $D^{\ast}\equiv d$, we obtain
$R_{\bm{x}}(d;\ell^{\textsc{Std}})-R_{\bm{x}}(d;\tilde{\ell}^{\textsc{Std}})$ $=\sum_{y\in\mathcal{Y}}\left(\ell^{\textsc{Std}}(d,y,\bm{x})-\tilde{\ell}^{\textsc{Std}}(d,y,\bm{x})\right)\Pr(Y(d)=y\mid\bm{X}=\bm{x})$
$=\mathbb{E}\left[\ell^{\textsc{Std}}(d,Y(d),\bm{X})-\tilde{\ell}^{\textsc{Std}}(d,Y(d),\bm{X})\ \bigr{|}\ \bm{X}=\bm{x}\right].$
Since this holds for $d=0,1$ and as the risk difference is invariant to the decision, we obtain the following equality,
$\mathbb{E}\left[\ell^{\textsc{Std}}(0,Y(0),\bm{X})-\tilde{\ell}^{\textsc{Std}}(0,Y(0),\bm{X})\ \bigr{|}\ \bm{X}=\bm{x}\right]=\mathbb{E}\left[\ell^{\textsc{Std}}(1,Y(1),\bm{X})-\tilde{\ell}^{\textsc{Std}}(1,Y(1),\bm{X})\ \bigr{|}\ \bm{X}=\bm{x}\right].$
for all $\bm{x}$. The left hand side depends on the distribution of $Y(0)$ given $\bm{X}=\bm{x}$, while the right hand side depends on the distribution of $Y(1)$ given $\bm{X}=\bm{x}$. Since there is no restriction on these distributions, the above equality must hold for every choice of these two conditional distributions. This is only possible if the following equality holds,
$\ell^{\textsc{Std}}(d,Y(d),\bm{x})-\tilde{\ell}^{\textsc{Std}}(d,Y(d),\bm{x})=\overline{C}(d,\bm{x}).$
By Equation (D.1), we must have $\overline{C}(d,\bm{x})\equiv\widetilde{C}(\bm{x})$ for $d=0,1$. That is, the difference in loss does not depend on the decision. Hence, we have shown $\ell^{\textsc{Std}}(d,y,\bm{x})=\tilde{\ell}^{\textsc{Std}}(d,y,\bm{x})+\widetilde{C}(\bm{x})$, implying the uniqueness of $\ell^{\textsc{Std}}(d,Y(d),\bm{x})$ up to an additive constant that only depends on covariates.
Finally, the second part of the proposition concerning surjectivity of the mapping follows immediately because any standard loss is a special case of an additive counterfactual loss. For any standard loss $\ell^{\textsc{Std}}(d,y_{d},\bm{x})$ and $\lambda\in\mathbb{R}$, define the weights
$\omega_{d}(d,y,\bm{x})=(1+\lambda)\ell^{\textsc{Std}}(d,y,\bm{x}),\ \quad\ \omega_{d}(1-d,y,\bm{x})=\lambda\ell^{\textsc{Std}}(d,y,\bm{x}),$
for $d=0,1$ and $y\in\mathcal{Y}$. Then, by the first part of this proposition, this corresponds to an additive counterfactual loss with the same treatment recommendations. Since $\lambda$ is arbitrary, there exist infinitely many additive counterfactual losses that map to the same standard loss. ∎
Appendix E Proof of Proposition 2
Let $\ell^{\textsc{Add}}(d;\bm{y})$ denote an additive counterfactual loss defined in Definition 3. Without loss of generality let $\varpi\equiv 0$, as this term does not depend on the decision. Suppose that there exists a standard loss $\ell^{\textsc{Std}}(d;y_{d})$ such that the risk difference $R(D^{\ast};\ell^{\textsc{Add}})-R(D^{\ast};\ell^{\textsc{Std}})$ does not depend on $D^{\ast}$. This implies that for fixed $\bm{X}=\bm{x}$, there exists $C(\bm{x})$ such that the following conditional risk difference does not depend on $D^{\ast}$,
$\Delta_{\bm{x}}(D^{\ast})=R_{\bm{x}}(D^{\ast};\ell^{\textsc{Add}})-R_{\bm{x}}(D^{\ast};\ell^{\textsc{Std}})=C(\bm{x}).$ (E.1)
Otherwise, we could choose $D^{\ast}$ to be independent of $\bm{X}$, making the expectation of Equation (E.1) dependent on $D^{\ast}$ and leading to a contradiction.
Because this conditional risk difference must remain fixed no matter how we choose $D^{\ast}$, it must stay identical across all deterministic decision rules. Consider the following constant decision rule $\pi_{j}=\pi_{j}(\bm{X})=j$ for $j\in\mathcal{D}$. Let
$\Delta_{\bm{x}}(\pi_{j})$ $=\sum_{k\in\mathcal{D}}\sum_{y_{k}\in\mathcal{Y}}\omega_{k}(j,y_{k},\bm{x})\Pr(Y(k)=y_{k}\mid\bm{X}=\bm{x})-\sum_{y_{j}\in\mathcal{Y}}\tilde{\omega}(j,y_{j},\bm{x})\Pr(Y(j)=y_{j}\mid\bm{X}=\bm{x})$
$=\sum_{k\in\mathcal{D}}\mathbb{E}[\omega_{k}(j,Y(k),\bm{X})-\tilde{\omega}(j,Y(j),\bm{X})\mid\bm{X}=\bm{x}],$
where $\tilde{\omega}(j,y_{j},\bm{x})$ is the weight function for the standard loss. Note that $\Delta(\pi_{j})=C(\bm{x})$ by Equation (E.1) for every $j\in\mathcal{D}$. Let $j,j^{\prime}\in\mathcal{D}$ with $j\neq j^{\prime}$. Then, $\Delta_{\bm{x}}(\pi_{j})-\Delta_{\bm{x}}(\pi_{j^{\prime}})=C(\bm{x})-C(\bm{x})=0$ or equivalently,
$\sum_{k\in\mathcal{D}}\mathbb{E}[\omega_{k}(j,Y(k),\bm{X})-\omega_{k}(j^{\prime},Y(k),\bm{X})\mid\bm{X}=\bm{x}]=\mathbb{E}[\tilde{\omega}(j,Y(j),\bm{X})-\tilde{\omega}(j^{\prime},Y(j^{\prime}),\bm{X})\mid\bm{X}=\bm{x}].$
Since the potential outcomes $\{Y(0),\ldots,Y(K-1)\}$ can have an arbitrary conditional distribution given $\bm{X}=\bm{x}$, the above equality must hold for every choice of this distribution. Since the right-hand side of the equation depends at most on the conditional distribution of $\{Y(j),Y(j^{\prime})\}$, so must its left-hand side. In particular, for every $k\neq\{j,j^{\prime}\}$ we must have $\omega_{k}(j,y_{k},\bm{x})=\omega_{k}(j^{\prime},y_{k},\bm{x})$, for all $y_{k}\in\mathcal{Y}$ and $\bm{x}\in\mathcal{X}$. Note that such $k$ exists so long as $K\geq 3$. Thus, $\omega_{k}(d,y_{k},\bm{x})$ cannot depend on the decision $d$ whenever $d\neq k$, implying that there exists functions $c_{k}$ such that $\omega_{k}(d,y,\bm{x})=c_{k}(y,\bm{x})$, for $d\neq k$. However, this contradicts the assumption that $\ell^{\textsc{Add}}$ is an additive counterfactual loss with at least one counterfactual weight depending on the decision and potential outcome.
Appendix F Proof of Corollary 3
By Theorem 2, the conditional counterfactual risk can be written as,
$R_{\bm{x}}(D)$ $=\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D}}\sum_{y=0}^{1}\omega_{k}(d,y,\bm{x})\Pr(D=d,Y(k)=y\mid\bm{X}=\bm{x})+C(\bm{x}),$ (F.1)
where $\sum_{\bm{y}\in\{0,1\}^{K}}\varpi(\bm{y},\bm{x})\Pr(\bm{Y}(\mathcal{S})=\bm{y}_{\mathcal{S}}\mid\bm{X}=\bm{x})$. First note,
$\Pr(D^{*}=d,Y(k)=0\mid\bm{X}=\bm{x})=\Pr(D^{*}=d\mid\bm{X}=\bm{x})-\Pr(D^{*}=d,Y(k)=1\mid\bm{X}=\bm{x}).$
Substituting this expression into the first term of Equation (F.1) yields,
$\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D}}\sum_{y=0}^{1}\omega_{k}(d,y,\bm{x})\Pr(D^{*}=d,Y(k)=y\mid\bm{X}=\bm{x})$
=\ = $\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D}}\omega_{k}(d,0,\bm{x})\Pr(D^{*}=d,Y(k)=0\mid\bm{X}=\bm{x})+\omega_{k}(d,1,\bm{x})\Pr(D^{*}=d,Y(k)=1\mid\bm{X}=\bm{x})$
=\ = $\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D}}\omega_{k}(d,0,\bm{x})\left\{\Pr(D^{*}=d\mid\bm{X}=\bm{x})-\Pr(D^{*}=d,Y(k)=1\mid\bm{X}=\bm{x})\right\}$
$\quad+\omega_{k}(d,1,\bm{x})\Pr(D^{*}=d,Y(k)=1\mid\bm{X}=\bm{x})$
=\ = $\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D}}\omega_{k}(d,0,\bm{x})\Pr(D^{*}=d\mid\bm{X}=\bm{x})+\left\{\omega_{k}(d,1,\bm{x})-\omega_{k}(d,0,\bm{x})\right\}\Pr(D^{*}=d,Y(k)=1\mid\bm{X}=\bm{x})$
=\ = $\sum_{d\in\mathcal{D}}\xi(d,\bm{x})\Pr(D^{*}=d\mid\bm{X}=\bm{x})+\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D}}\zeta_{k}(d,\bm{x})\Pr(D^{*}=d,Y(k)=1\mid\bm{X}=\bm{x})$
=\ = $\sum_{d\in\mathcal{D}}\xi(d,\bm{x})\Pr(D^{*}=d\mid\bm{X}=\bm{x})+\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D},k\neq d}\zeta_{k}(d,\bm{x})\Pr(D^{*}=d,Y(k)=1\mid\bm{X}=\bm{x})$
$+\sum_{d=1}^{K}\zeta_{d}(d,\bm{x})\Pr(D^{*}=d,Y(d)=1\mid\bm{X}=\bm{x}),$
where $\zeta_{k}(d,\bm{x})=\omega_{k}(d,1,\bm{x})-\omega_{k}(d,0,\bm{x})$, $\xi(d,\bm{x})=\sum_{k\in\mathcal{D}}\omega_{k}(d,0,\bm{x})$. ∎
Appendix G Lemmas and their Proofs
G.1 Lemma 1
Lemma 1
Let $\mathcal{P}$ denote the set of all joint probability distributions under consideration, and let $\mathcal{Q}$ be the set of all observable distributions. Let $Q:\mathcal{P}\to\mathcal{Q}$ denote a surjective mapping that associates each joint distribution with an observable distribution. Let $\theta(P)$ denote the (causal) estimand of interest under distribution $P\in\mathcal{P}$. For a fixed observable distribution $Q_{0}\in\mathcal{Q}$ define $\mathcal{S}_{0}=\{P\in\mathcal{P}:Q(P)=Q_{0}\}.$ Then, the following statements are equivalent:
  1. (a)
    For every $Q_{0}\in\mathcal{Q}$ there exists a constant $\theta_{0}$ such that $\theta(P)=\theta_{0}$ for all $P\in\mathcal{S}_{0}$.
  2. (b)
    There exists a function $f$ such that $\theta(P)=f(Q(P))$ for every $P\in\mathcal{P}$.
Proof.
We first prove that (b) implies (a). Fix any $Q_{0}\in\mathcal{Q}$ and let $P\in\mathcal{S}_{0}$ such that $Q(P)=Q_{0}$. Note that $\mathcal{S}_{0}$ is non-empty by $Q$ surjective. Then, by (b), there exists $f$ such that
$\theta(P)=f(Q(P))=f(Q_{0})=\theta_{0},$
for every $P\in\mathcal{S}_{0}$. Because $Q_{0}$ was arbitrary, the claim follows.
We next prove that (a) implies (b). Let $Q_{0}\in\mathcal{Q}$. Surjectivity of $Q$ ensures $\mathcal{S}_{0}$ is non-empty. By (a), there is a constant $\theta_{0}$ such that $\theta(P)=\theta_{0}$ for all $P\in\mathcal{S}_{0}$. Define $f:\mathcal{Q}\to\mathbb{R}$ by setting $f(Q_{0}):=\theta_{0}$ for every such $Q_{0}\in\mathcal{Q}$. Now, let $P\in\mathcal{P}$ be an arbitrary probability measure and let $Q_{0}=Q(P)$. Then, there exists $\theta_{0}=f(Q_{0})$ such that $\theta(P)=\theta_{0}=f(Q_{0})=f(Q(P))$. This proves the claim. ∎
G.2 Lemma 2
Lemma 2
Given a counterfactual loss $\ell$ defined in Definition 1, the counterfactual risk $R(D^{\ast},\ell)$ defined in Definition 2 is identifiable with respect to $\mathbb{P}(D^{\ast},D,Y,\bm{X})$ if and only if the conditional counterfactual risk $R_{\bm{X}}(D^{\ast},\ell)$ (also defined in Definition 2) is identifiable with respect to $\mathbb{P}(D^{\ast},D,Y\mid\bm{X})$.
Proof.
Let $Q:\mathcal{P}\to\mathcal{Q}$ be the mapping from $P\in\mathcal{P}$ to the distribution $Q(P)=P(D^{\ast},D,Y,\bm{X})$. Further, let $Q_{\bm{X}}:\mathcal{P}\to\mathcal{Q}_{\bm{X}}$ be the mapping from $P\in\mathcal{P}$ to the conditional distribution $Q_{\bm{X}}(P)=P(D^{\ast},D,Y\mid\bm{X})$. We use $P_{\bm{X}}$ to denote the marginal probability distribution of $\bm{X}$. Note that for any $P\in\mathcal{P}$ we can write
$P(D^{\ast},D,Y(\mathcal{D}),\bm{X})$ $=P(D^{\ast},D,Y,Y(\mathcal{D}),\bm{X})$
$=P(Y(\mathcal{D})\mid D^{\ast},D,Y,\bm{X})Q(P)$
$=P(Y(\mathcal{D})\mid D^{\ast},D,Y,\bm{X})Q_{\bm{X}}(P)P_{\bm{X}}(\bm{X}),$
where the first equality follows from Assumption 2. To ease notation, we write $R=R(D^{\ast},\ell)$ and $R_{\bm{x}}=R_{\bm{x}}(D^{\ast},\ell)$.
We first assume that $R$ is identifiable. We use proof by contradiction. Suppose that $R_{\bm{x}}$ is not identifiable. Then, by Definition 4, there exist a conditional observable distribution $Q_{0}(\bm{X})\subseteq\mathcal{Q}_{\bm{X}}$ and distributions $P_{1},P_{2}\in\mathcal{P}$ such that $Q_{\bm{X}}(P_{1})=Q_{\bm{X}}(P_{2})=Q_{0}(\bm{X})$ but $\Delta(\bm{x})=R_{\bm{x}}(P_{1})-R_{\bm{x}}(P_{2})\neq 0$ on some measurable set $\widetilde{\mathcal{X}}\subseteq\mathcal{X}$. Let $\widetilde{P}_{\bm{X}}$ be a probability measure such that
[…]
We can always construct such a probability measure $P_{\bm{X}}(\bm{x})$, for example, by concentrating the probability mass on a set $\widetilde{\mathcal{X}}$ where $\Delta(\bm{x})>0$ for all $\bm{x}\in\widetilde{X}$. Next, define another probability measure $\widetilde{P}_{i}$ for $i=1,2$:
$\widetilde{P}_{i}(D^{\ast},D,Y(\mathcal{D}),\bm{X})=P_{i}(Y(\mathcal{D})\mid D^{\ast},D,Y,\bm{X})Q_{0}(\bm{X})\widetilde{P}_{\bm{X}}(\bm{X}).$
where each $\widetilde{P}_{i}\in\mathcal{P}$ satisfies Assumptions 13, as these assumptions do not restrict the marginal of $\bm{X}$. Since the marginal distribution of $\bm{X}$ does not affect the conditional risk, we have $R_{\bm{x}}(\widetilde{P}_{i})=R_{\bm{x}}(P_{i})$ for $i=1,2$ and
$R_{\bm{x}}(\widetilde{P}_{1})-R_{\bm{x}}(\widetilde{P}_{2})=R_{\bm{x}}(P_{1})-R_{\bm{x}}(P_{2})=\Delta(\bm{x}).$
Further, by construction, we have $Q(\widetilde{P}_{1})=Q_{0}(\bm{X})\widetilde{P}_{\bm{X}}(\bm{X})=Q(\widetilde{P}_{2})\in\mathcal{Q}$. Since $R$ is identifiable, Definition 4 implies that $R(\widetilde{P}_{1})=R(\widetilde{P}_{2})$. However,
$R(\widetilde{P}_{1})-R(\widetilde{P}_{2})=$ $\int_{\bm{x}\in\mathcal{X}}R_{\bm{x}}(\widetilde{P}_{1})-R_{\bm{x}}(\widetilde{P}_{2})d\widetilde{P}_{\bm{X}}(\bm{x})$
$=$ $\int_{\bm{x}\in\mathcal{X}}R_{\bm{x}}(P_{1})-R_{\bm{x}}(P_{2})d\widetilde{P}_{\bm{X}}(\bm{x})$
[…]
$\neq$ $0,$
which leads to a contradiction. Thus, $R_{\bm{x}}$ must be identifiable.
Assume next that $R_{\bm{X}}$ is identifiable. Then, by Lemma 1, there exists a function $g$ such that $R_{\bm{X}}(P)=g(Q_{\bm{X}}(P))$. Define a function $f:\mathcal{Q}\to\mathbb{R}$ such that
[…]
Note that we can define $f$ in such a way as both $Q_{\bm{X}}(P)$ and $P_{\bm{X}}$ are obtainable from $Q$. Hence $R(P)=f(Q(P))$ and so $R$ is identifiable by Lemma 1. ∎
G.3 Lemma 3
Lemma 3
Let $D,Y_{0},\ldots,Y_{K-1}$ be discrete random variables such that $D\in\mathcal{D}=\{0,1,\ldots,K-1\}$ and $Y_{0},\ldots,Y_{K-1}\in\mathcal{Y}=\{0,1,\ldots,M-1\}$. Let $\Delta_{N-1}=\{\bm{p}\in\mathbb{R}^{N}\mid\sum_{i=1}^{N}p_{i}=1\text{ and }p_{i}\geq 0\text{ for }i=1,\ldots,N\}$ denote the probability simplex in $\mathbb{R}^{N}$ where $N=K\times M^{K}$. Then, for every probability distribution $P(D,Y_{0},\ldots,Y_{K-1})$, there exists $\bm{p}\in\Delta_{N-1}$ where $p_{i}=\Pr(D=d,Y_{0}=y_{0},\ldots,Y_{K-1}=y_{K-1})$ for each $i=i(d,\bm{y})\in\{1,\ldots,N\}$. Consider the following two cases that yield the same conclusion.
  1. (a)
    Let $\bm{q}=\bm{q}(\bm{p})\in[0,1]^{L}$ be a vector of probabilities $\{\Pr(D=d,Y_{k}=y_{k})\}_{d\in\mathcal{D},k\in\mathcal{D},y_{k}\in\mathcal{Y}}$ where $q_{j}=\Pr(D=d,Y_{k}=y_{k})$ for $j=j(d,k,y_{k})\in\{1,\ldots,L\}$ and $L=K^{2}\times M$.
  2. (b)
    Let $\bm{q}=\bm{q}(\bm{p})\in[0,1]^{L}$ be the extended vector of probabilities that also includes the joint distribution of $Y_{0},\ldots,Y_{K-1}$, i.e., probabilities $\{\Pr(D=d,Y_{k}=y_{k})\}_{d\in\mathcal{D},\,k\in\mathcal{D},\,y_{k}\in\mathcal{Y}}$ and $\{\Pr(Y_{0}=y_{0},\ldots,Y_{K-1}=y_{K-1})\}_{\bm{y}\in\mathcal{Y}^{K}}$, with $L=K^{2}\times M+M^{K}$.
Then, in each case, there exists an $L\times N$ matrix $\bm{C}$ of zeros and ones such that $\bm{q}=\bm{C}\bm{p}$ for all $\bm{p}\in\Delta_{N-1}$ and
$\text{ker}(\bm{C})=\left\{\bm{v}\in\mathbb{R}^{N}:\sum_{\bm{y}_{-k}\in\mathcal{Y}^{K-1}}v_{d,\bm{y}}=0\ \text{for all }d\in\mathcal{D},k\in\mathcal{D},y_{k}\in\mathcal{Y},\ \mathbf{1}_{N}^{\top}\bm{v}=0\ \text{and}\ \bm{C}\bm{v}=\mathbf{0}_{L}\right\},$
where $\bm{y}_{-k}=(y_{0},\ldots,y_{k-1},y_{k+1},\ldots,y_{K-1})$ and $\mathbf{1}_{N}$ represents an $N$-dimensional vector of ones. Furthermore, $\bm{C}$ is a surjective map from the space of joint distribution $P(D,Y_{0},\ldots,Y_{K-1})$ onto the space of the respective marginal distribution.
We now prove each case in turn.
Case (a).
For every joint probability distribution $P(D,Y_{0},\ldots,Y_{K-1})$, we can express the marginal distribution of $P(D,Y_{k})$ as,
$\Pr(D=d,Y_{k}=y_{k})=\sum_{d^{\prime}\in\mathcal{D}}\sum_{\bm{y}^{\prime}\in\mathcal{Y}^{K}}\mathds{1}{\left\{d^{\prime}=d,y^{\prime}_{k}=y_{k}\right\}}\Pr(D=d^{\prime},Y_{0}=y^{\prime}_{0},\ldots,Y_{K-1}=y^{\prime}_{K-1}).$ (G.1)
Define a matrix $\bm{C}\in\{0,1\}^{L\times N}$ whose $(j,i)$ entry is given by,
$C_{\,j(d,k,y_{k}),\,i(d^{\prime},\bm{y}^{\prime})}=\mathds{1}{\left\{d^{\prime}=d,y^{\prime}_{k}=y_{k}\right\}}.$
Then, Equation (G.1) can be written as,
$q_{j(d,k,y_{k})}=\sum_{d^{\prime}\in\mathcal{D}}\sum_{\bm{y}^{\prime}\in\mathcal{Y}^{K}}C_{\,j(d,k,y_{k}),\,i(d^{\prime},\bm{y}^{\prime})}p_{i(d^{\prime},\bm{y}^{\prime})},$
implying $\bm{q}=\bm{C}\bm{p}$. Since every marginal distribution has a corresponding joint probability distribution, $\bm{C}$ is a surjective map from the space of joint probability distributions onto the space of marginal distributions.
We proceed to examine the kernel of the matrix $\bm{C}$. Let $\bm{v}=(v_{d,\bm{y}})\in\mathbb{R}^{N}$. The definition of $\bm{C}$ implies that for all $j=j(d,k,y_{k})\in\{1,\ldots,L\}$, we have,
$(\bm{C}\bm{v})_{j(d,k,y_{k})}=\sum_{d^{\prime}\in\mathcal{D}}\sum_{\bm{y}^{\prime}\in\mathcal{Y}^{K}}\mathds{1}{\left\{d^{\prime}=d,y^{\prime}_{k}=y_{k}\right\}}\bm{v}_{(d^{\prime},\bm{y})}=\sum_{\bm{y}_{-k}\in\mathcal{Y}^{K-1}}v_{d,\bm{y}_{-k}}.$
Therefore, $\bm{C}\bm{v}=\mathbf{0}_{L}$ if and only if $\sum_{\bm{y}_{-k}\in\mathcal{Y}^{K-1}}v_{d,\bm{y}}=0$ for all $d\in\mathcal{D},\;k\in\mathcal{D},\;y_{k}\in\mathcal{Y}$. This proves the first part of the kernel property. The second part follows immediately by observing,
$\mathbf{1}_{N}^{\top}\bm{v}=\sum_{d\in\mathcal{D}}\sum_{\bm{y}\in\mathcal{Y}^{K}}v_{d,\bm{y}}=\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D}}\sum_{y_{k}\in\mathcal{Y}}\sum_{\bm{y}_{-k}\in\mathcal{Y}^{K-1}}v_{d,\bm{y}}=0.$
Case (b).
The existence and surjectivity of $\bm{C}$ follow from the same arguments as in the above proof of Case (a). Thus, we only need to verify the kernel property of $\bm{C}$. We start by partitioning
$\bm{C}=\begin{bmatrix}\bm{C}_{1}\\ \bm{C}_{2}\\ \end{bmatrix},$
where $\bm{C}_{1}$ is an $K^{2}M\times N$ matrix of zeros and ones, containing all the rows corresponding to the marginal distributions $\{\Pr(D=d,Y_{k}=y_{k})\}_{d\in\mathcal{D},k\in\mathcal{D},y_{k}\in\mathcal{Y}}$, which are included in $\bm{C}$ by assumption. Let $\bm{v}\in\text{Ker}(\bm{C})$, then,
$\bm{C}\bm{v}=\begin{bmatrix}\bm{C}_{1}\bm{v}\\ \bm{C}_{2}\bm{v}\\ \end{bmatrix}=\begin{bmatrix}\bm{0}\\ \bm{0}\\ \end{bmatrix},$
implying $\bm{C}_{1}\bm{v}=0$. Thus, $\bm{v}\in\text{ker}(\bm{C}_{1})$ and consequently, $\text{ker}(\bm{C})\subseteq\text{ker}(\bm{C}_{1})$. Since $\bm{C}_{1}$ corresponds to the matrix described in Case (a), this completes the proof. ∎
G.4 Lemma 4
Lemma 4
Let $\Delta_{N-1}=\{\bm{p}\in\mathbb{R}^{N}\mid\sum_{i=1}^{N}p_{i}=1\text{ and }p_{i}\geq 0\text{ for }i=1,\ldots,N\}$ denote the probability simplex in $\mathbb{R}^{N}$. Let $\bm{C}\in\mathbb{R}^{L\times N}$ be a matrix such that $\text{ker}(\bm{C})=\left\{\bm{v}\in\mathbb{R}^{N}:\mathbf{1}_{N}^{\top}\bm{v}=0\ \text{and}\ \bm{C}\bm{v}=\mathbf{0}_{L}\right\}$. Set $\bm{C}(\Delta_{n-1})=\{\bm{q}\in\mathbb{R}^{L}\mid\exists\bm{p}\in\Delta_{n-1}:\bm{q}=\bm{C}\bm{p}\}$ to the image of $\bm{C}$ on $\Delta_{n-1}$. Let $f_{\bm{\ell}}:\bm{C}(\Delta_{n-1})\to\mathbb{R}$ be a function for which there exists $\bm{\ell}\in\mathbb{R}^{N}$ such that for all $\bm{q}(\bm{p})\in\bm{C}(\Delta_{n-1})$, we have $f_{\bm{\ell}}(\bm{q})=f_{\bm{\ell}}(\bm{C}\bm{p})=\bm{\ell}^{\top}\bm{p}$.
Then, $f_{\bm{\ell}}$ is well-defined and linear. That is, $\bm{\ell}\in\text{im}(\bm{C}^{\top})$, and there exists $\bm{w}\in\mathbb{R}^{L}$ only depending on $\bm{C}$ and $\bm{\ell}$ such that $f_{\bm{\ell}}(\bm{q})=\bm{w}^{\top}\bm{q}$ holds.
Proof.
Observe that if $\bm{\ell}\in\text{im}(\bm{C}^{\top})$, then there exists $\bm{w}\in\mathbb{R}^{L}$ such that $\bm{\ell}=\bm{C}^{\top}\bm{w}$. Then, for any $\bm{q}=\bm{C}\bm{p}$ we have
$f_{\bm{\ell}}(\bm{q})=\bm{\ell}^{\top}\bm{p}=\bigl{(}\bm{C}^{\top}\bm{w}\bigr{)}^{\!\top}\bm{p}=\bm{w}^{\top}\bm{C}\bm{p}=\bm{w}^{\top}\bm{q}.$
Hence, $f_{\bm{\ell}}(\bm{q})=\bm{w}^{\top}\bm{q}$.
Assume now that $\bm{\ell}\notin\text{im}(\bm{C}^{\top})$. We will prove that this yields a contradiction. By standard linear algebra [Lax, 2007, Theorem 5, Chapter 3], $\text{im}(\bm{C}^{\top})=\text{ker}(\bm{C})^{\perp}$. This implies that $\exists\bm{v}\in\text{ker}(\bm{C})\setminus\{\bm{0}\}:\bm{\ell}^{\top}\bm{v}\neq 0$. Let $\bm{p}_{0}\in\mathring{\Delta}_{n-1}$ be in the interior of the simplex with strictly positive entries and set $\bm{q}_{0}=\bm{C}\bm{p}_{0}$. Define
$\bm{p}=\bm{p}_{0}+\alpha\bm{v}\quad\text{with}\ 0<\alpha<\frac{\min_{i\in[N]}{p_{0}}_{i}}{\max_{i\in[N]}\left|v_{i}\right|}.$
Note that such an $\alpha$ exists. Indeed, by assumption, $0<{p_{0}}_{i}$ for all entries and $\bm{v}$ has at least one non-zero entry as $\bm{v}\neq\bm{0}$. Observe,
$\min_{i\in[n]}p_{i}\geq\min_{i\in[n]}{p_{0}}_{i}-\alpha\max_{i\in[n]}\left|v_{i}\right|>0,$
and,
$\sum_{i=1}^{n}p_{i}=\sum_{i=1}^{n}{p_{0}}_{i}+\alpha\sum_{i=1}^{n}v_{i}=\sum_{i=1}^{n}{p_{0}}_{i}=1,$
where the second equality follows from $\sum_{i=1}^{n}v_{i}=0$ by $\bm{v}\in\text{ker}(\bm{C})$. Thus, $\bm{p}\in\Delta_{N-1}$.
Because $\bm{v}\in\text{ker}(\bm{C})$,
$\bm{C}\bm{p}=\bm{C}\bigl{(}\bm{p}_{0}+\alpha\bm{v}\bigr{)}=\bm{C}\bm{p}_{0}=\bm{q}_{0}.$
This implies that $f_{\bm{\ell}}(\bm{q}_{0})=f_{\bm{\ell}}(\bm{C}\bm{p})=f_{\bm{\ell}}(\bm{C}\bm{p}_{0})$ and,
$0=f_{\bm{\ell}}(\bm{C}\bm{p})-f_{\bm{\ell}}(\bm{C}\bm{p}_{0})=\bm{\ell}^{\top}\bm{p}-\bm{\ell}^{\top}\bm{p}_{0}=\bm{\ell}^{\top}\left(\bm{p}-\bm{p}_{0}\right)=\alpha\bm{\ell}^{\top}\bm{v}\neq 0,$
as $\bm{\ell}^{\top}\bm{v}\neq 0$ by assumption. This yields a contradiction. ∎
Appendix H Derivation of Weights
We illustrate the linear system described in Equation 10 by considering the binary decision and outcome case $Y\in\{0,1\}$. This example corresponds to the identifiability conditions discussed in Section 3.6. Recall that we are trying to solve a set of linear equations
$\bm{A}^{\textsc{df}}\bm{w}^{\textsc{df}}=\bm{\ell},$
where $\textsc{df}\in\{4,6,7\}$ and
$A=\left[\begin{array}[]{cccccccccccc}1&0&0&0&0&0&1&0&1&0&0&0\\ 0&0&1&0&1&0&0&0&1&0&0&0\\ 0&1&0&0&0&0&1&0&0&0&1&0\\ 0&0&1&0&0&1&0&0&0&0&1&0\\ 1&0&0&0&0&0&0&1&0&1&0&0\\ 0&0&0&1&1&0&0&0&0&1&0&0\\ 0&1&0&0&0&0&0&1&0&0&0&1\\ 0&0&0&1&0&1&0&0&0&0&0&1\end{array}\right],\qquad\bm{w}=\begin{bmatrix}\omega_{0}(0,0)\\ \omega_{0}(0,1)\\ \omega_{1}(1,0)\\ \omega_{1}(1,1)\\ \omega_{0}(1,0)\\ \omega_{0}(1,1)\\ \omega_{1}(0,0)\\ \omega_{1}(0,1)\\ \varpi_{0,1}(0,0)\\ \varpi_{0,1}(0,1)\\ \varpi_{0,1}(1,0)\\ \varpi_{0,1}(1,1)\end{bmatrix},\qquad\bm{\ell}=\begin{bmatrix}\ell(0;0,0)\\[5.0pt] \ell(1;0,0)\\[5.0pt] \ell(0;1,0)\\[5.0pt] \ell(1;1,0)\\[5.0pt] \ell(0;0,1)\\[5.0pt] \ell(1;0,1)\\[5.0pt] \ell(0;1,1)\\[5.0pt] \ell(1;1,1)\end{bmatrix}$
with
$\bm{A}^{4}=\bm{A}_{1:4},\bm{w}^{4}=\bm{w}_{1:4},\qquad\bm{A}^{6}=A_{1:8},\bm{w}^{6}=\bm{w}_{1:8},\qquad\bm{A}^{7}=\bm{A},\bm{w}^{7}=\bm{w}.$
Here, $A_{1:k}$ denotes the first $k$ rows of matrix $A$, and $\bm{w}_{1:k}$ refers to the first $k$ entries of the weight vector $\bm{w}$. Note that the rank of $\bm{A}$ is seven. These three systems correspond to the identifiability conditions summarized in Table 3. The most restrictive case, with four degrees of freedom, corresponds to $\textsc{df}=4$; the intermediate case, with six degrees of freedom, yielding exact identifiability under strong ignorability,corresponds to $\textsc{df}=6$; and the most relaxed case, with seven degrees of freedom, corresponds to $\textsc{df}=7$. By solving each linear system and characterizing its image, $\text{im}(\bm{A}^{\textsc{df}})$, we recover the precise constraints on the loss function that determine whether the counterfactual risk is exactly identifiable, identifiable up to a constant, or unidentifiable, as discussed in Section 3.5.