### v2

Abstract
Many researchers have applied classical statistical decision theory to evaluate treatment choices and learn optimal policies. However, because this framework is based solely on realized outcomes under chosen decisions and ignores counterfactual outcomes, it cannot assess the quality of a decision relative to feasible alternatives. For example, in bail decisions, a judge must consider not only crime prevention but also the avoidance of unnecessary burdens on arrestees. To address this limitation, we generalize standard decision theory by incorporating counterfactual losses, allowing decisions to be evaluated using all potential outcomes. The central challenge in this counterfactual statistical decision framework is identification: since only one potential outcome is observed for each unit, the associated counterfactual risk is generally not identifiable. We prove that, under the assumption of strong ignorability, the counterfactual risk is identifiable if and only if the counterfactual loss function is additive in the potential outcomes. Moreover, we demonstrate that additive counterfactual losses can yield treatment recommendations, which differ from those based on standard loss functions when the decision problem involves more than two treatment options. One interpretation of this result is that additive counterfactual losses can capture the accuracy and difficulty of a decision, whereas standard losses account for accuracy alone. Finally, we formulate a symbolic linear inverse program that, given a counterfactual loss, determines whether its risk is identifiable, without requiring data.
Keywords: causal inference, expected utility, identification, policy learning, potential outcomes, risk
1 Introduction
Statistical decision theory, pioneered by Wald [1950], formulates decision-making as a game against nature. A decision-maker selects an action $D=d$ in the face of an unknown state of nature $\theta$, and incurs a prespecified loss (or negative utility) $\ell(d;\theta)$, which quantifies the consequence of the chosen action when the true state of nature is $\theta$. A better decision corresponds to a lower loss. Given covariates $\bm{X}$, we can define a decision rule $D=\pi(\bm{X})$ that maps observed data to actions. The performance of a decision rule is commonly evaluated by its risk, that is, the expected loss:
$R(\pi;\theta,\ell)\ :=\ \mathbb{E}\left[\ell(\pi(\bm{X});\theta)\right],$
where the expectation is taken over the distribution of $\bm{X}$. Since $\theta$ is unknown, the true risk is typically unidentifiable. This motivates the use of optimality criteria such as admissibility, minimax, minimax-regret, and Bayes risk [Wald, 1950, Savage, 1972, Berger, 1985]. Statistical decision theory also provides a foundational framework for classical estimation theory [e.g., Lehmann and Casella, 2006] and hypothesis testing [e.g., Lehmann et al., 1986].
In a series of influential papers, Manski [2000, 2004, 2011] applies this framework to the problem of treatment choice. In its simplest form, the goal is to choose a treatment $D$ to minimize the expected loss associated with an outcome of interest $Y$. Because the treatment may influence the outcome, the loss $\ell(d;Y(d))$ depends on the potential outcome $Y(d)$ under the chosen treatment $d\in\mathcal{D}$ and possibly the treatment itself, where $\mathcal{D}$ denotes the support of decision $D$ [Neyman, 1923, Rubin, 1974]. As before, we can define a treatment rule $D=\pi(\bm{X})$ based on observed covariates, and evaluate its performance via the following risk:
$R(\pi;\ell)\ =\ \mathbb{E}{\left[\ell(\pi(\bm{X});Y(\pi(\bm{X})))\right]},$ (1)
where the expectation is taken over the joint distribution of $\bm{X}$ and $Y(\pi(\bm{X}))$. This contrasts with the classical statistical decision theory, which would condition on the potential outcome vector. Since only one potential outcome is observed per unit, i.e., $Y=Y(D)$, the risk is generally unidentifiable unless the treatment rule coincides with the one used to generate the data. Consequently, additional assumptions, such as the strong ignorability of treatment assignment [Rosenbaum and Rubin, 1983], are required to identify the risk and learn optimal treatment rules [e.g., Manski, 2000, Kitagawa and Tetenov, 2018, Athey and Wager, 2021].
In this paper, we generalize the standard framework of statistical decision theory for treatment choice by introducing a broader class of loss functions. Specifically, we consider counterfactual losses, which depend on the full set of potential outcomes rather than the outcome under the selected treatment alone. That is, the loss takes the form of $\ell(d;\{Y(d^{\prime})\}_{d^{\prime}\in\mathcal{D}}\})$. We refer to this as a counterfactual loss because it evaluates each decision in light of the alternative outcomes that could have occurred had a different treatment been chosen. Given this generalized loss, we define the counterfactual risk for measuring the performance of treatment rule $\pi$ as:
$R(\pi;\ell)\ =\ \mathbb{E}[\ell(\pi(\bm{X});\{Y(d)\}_{d\in\mathcal{D}})],$ (2)
where the expectation is taken over the distribution of covariates $\bm{X}$ and all potential outcomes $\{Y(d)\}_{d\in\mathcal{D}}$. A central challenge introduced by this generalized framework is that the counterfactual risk given in Equation (2) is not identifiable even under the standard strong ignorability assumption, which is sufficient to identify the standard risk given in Equation (1).
The main contribution of this paper is to propose a structural condition on the counterfactual loss that enables the identification of counterfactual risk. Specifically, we consider a class of counterfactual losses that are additive in the potential outcomes. We show that under strong ignorability of treatment assignment, additivity is both necessary and sufficient for the identification of difference in the counterfactual risk between any two decision rules, permitting the identification of optimal treatment rules. Therefore, additivity represents the weakest possible condition on counterfactual losses that guarantees the identification of counterfactual risk under strong ignorability. Importantly, our results extend beyond the binary treatment and outcome cases that have been a focus of the existing literature.
Furthermore, we show that an identifiable counterfactual risk, based on an additive counterfactual loss, can yield an optimal treatment rule that differs from the one based on a standard risk, provided that the decision problem involves more than two treatment options, i.e., $|\mathcal{D}|>2$. This implies that in the binary treatment case, any identifiable counterfactual risk can be expressed equivalently as a standard risk. The difference between additive counterfactual and standard losses arises because the former accounts for both accuracy and difficulty of a decision whereas the latter captures the notion of accuracy alone. Finally, we formulate a linear inverse problem that offers a practical test for whether a given loss is additive and hence its risk identifiable. This test does not require any data, and the solution to this linear inverse problem provides a closed-form decomposition of counterfactual risk into identifiable marginal distribution of potential outcomes.
Related Literature.
The application of counterfactual logic to decision-theoretic problems has appeared in several strands of the literature. One of the most prominent examples of counterfactual loss is the notion of regret, $\ell(d;\{Y(d^{\prime})\}_{d^{\prime}\in\mathcal{D}})=\max_{d^{\prime}\in\mathcal{D}}Y(d^{\prime})-Y(d)$ [Savage, 1972]. Yet, because of the aforementioned identification challenge, prior work has mostly focused on the settings with binary treatment and binary outcome. For example, Coston et al. [2020] and Ben-Michael et al. [2025] consider the counterfactual loss is a function of the baseline potential outcome $Y(0)$ alone. A more general binary case has been studied by Li and Pearl [2019, 2022], Mueller and Pearl [2023], Ben-Michael et al. [2024], who address non-identifiability through partial-identification bounds and by leveraging observational data in some cases. In addition, Christy and Kowalski [2024], Christy and Kowalski [2025] circumvent non-identifiability by evaluating decisions with a finite-sample maximum-likelihood rule.
The most relevant work in the binary case is Li and Pearl [2019] who provides a sufficient condition for identification. We show that this condition follows directly from our general identification result and it is indeed necessary for the identification of counterfactual risk. However, our result implies that their condition can be relaxed if the goal is to identify a counterfactual risk difference and can be expressed as a mathematically equivalent standard loss.
Extensions to non-binary treatments have been studied by Li and Pearl [2024a, b]. Specifically, Li and Pearl [2024a] derive partial-identification bounds, while Li and Pearl [2024b] propose a non-constructive algorithm that numerically tests identifiability using the experimental data. In contrast, we provide a closed-form, analytical characterization of counterfactual losses that is both necessary and sufficient for the identification of counterfactual risk. We also extend these results to the identification of counterfactual risk difference, which suffices for learning optimal treatment rules. Our identification results lead to a symbolic linear inverse problem formulation that can be used, without data, to verify whether a counterfactual risk (or the difference between any pair of counterfactual risks) is identifiable. We further show that in the non-binary treatment case, an identifiable counterfactual loss need not reduce to a standard loss.
2 Illustrative Examples
Before presenting our general identification analysis in Section 3, we use illustrative examples to highlight the differences between standard and counterfactual losses. We begin by an example of binary decision and then consider a trichotomous decision example.
2.1 Binary Decisions
Consider a stylized setting, in which physician must decide whether to approve ($D=1$) a new medical treatment for a population of at-risk individuals, or not ($D=0$). The outcome of interest is survival, denoted by $Y$, where $Y=1$ indicates survival and $Y=0$ otherwise. For simplicity, we assume no covariates are observed. We now introduce standard, counterfactual, and additive counterfactual losses in the context of this example.
2.1.1 Standard Loss
Suppose that we evaluate the physician’s decision using the standard loss $\ell^{\textsc{Std}}(D;Y(D))$, which assigns a value to each realized decision–outcome pair $(D,Y(D))$. Let $c_{1}>0$ denote the cost of providing treatment relative to the cost of no intervention $c_{0}=0$. We use $\ell_{y}$ to denote the loss associated with the realized outcome $Y(D)=y$. For example, if the treatment is administered and the patient dies, the loss is given by $\ell^{\textsc{Std}}(1;0)=\ell_{0}+c_{1}$. If no treatment is given and the patient survives, the loss is $\ell^{\textsc{Std}}(0;1)=\ell_{1}$. Since survival is preferred, we assume $\ell_{1}<\ell_{0}$.
This yields the standard loss function,
$\ell^{\textsc{Std}}(D;Y(D))=\ell_{Y(D)}+c_{D}.$ (3)
An optimal decision $d^{\ast}$ minimizes the standard risk,
$d^{\ast}=\operatornamewithlimits{argmin}_{d\in\{0,1\}}R(d;\ell^{\textsc{Std}}),$
where $R(d;\ell^{\textsc{Std}})=\mathbb{E}{\left[\ell^{\textsc{Std}}(d;Y(d))\right]}$. The classical A/B test is a special case of this standard loss with $\ell^{\textsc{Std}}(D;Y(D))=-Y(D)$, where risk minimization reduces to the comparison of average survival rates.
2.1.2 Counterfactual Loss
The standard loss $\ell^{\textsc{Std}}$ only considers one potential outcome at a time, $Y(d)$ for $d\in\{0,1\}$. In contrast, the physician may decide to use a counterfactual loss that depends on both potential outcomes simultaneously. This naturally leads to the concept of principal strata, which classifies individuals according to their joint potential outcomes [Frangakis and Rubin, 2002]:
  • Never survivors $(Y(0),Y(1))=(0,0)$: patients who would not survive regardless of treatment
  • Responders $(Y(0),Y(1))=(0,1)$: patients who would survive only if given the treatment
  • Harmed $(Y(0),Y(1))=(1,0)$: patients who would survive only if not treated
  • Always survivors $(Y(0),Y(1))=(1,1)$: patients who would survive regardless of treatment
Building on this idea, Li and Pearl [2019], Ben-Michael et al. [2024], and Christy and Kowalski [2024] propose the following counterfactual loss:
$\ell^{\textsc{Cf}}(D;Y(0),Y(1))$ $=(1-Y(0))(1-Y(1))\ell_{0}+(1-Y(0))Y(1)\ell_{D}^{\text{R}}$ (4)
$\qquad+Y(0)(1-Y(1))\ell_{1-D}^{\text{H}}+Y(0)Y(1)\ell_{1}+c_{D}.$
This generalizes the standard loss given in Equation (3), which is obtained by setting $\ell_{D}^{R}=\ell_{D}$ and $\ell_{1-D}^{H}=\ell_{1-D}$.
Unlike the standard loss, therefore, the counterfactual loss incorporate the information about whether the treatment would alter the outcome. For example, responders and always survivors both incur the same standard loss $\ell^{\textsc{Std}}(1;1)=\ell_{1}+c_{1}$ under the treatment, but their counterfactual losses may differ (e.g., $\ell_{1}^{\text{R}}+c_{1}<\ell_{1}+c_{1}$ if we reward the correct decision). Ben-Michael et al. [2024] design a counterfactual loss that reflects a version of the Hippocratic Oath principle of “Do no harm” [Wiens et al., 2019], where killing a patient is worse than failing to provide life-saving treatment, i.e., $\ell_{0}^{\text{R}}-\ell_{1}^{\text{R}}<\ell_{0}^{\text{H}}-\ell_{1}^{\text{H}}$.
Counterfactual loss can also reveal heterogeneity that standard loss may conceal. For instance, Mueller and Pearl [2023] show a case where men and women have the same standard risk difference, $R(1;\ell^{\textsc{Std}})-R(0;\ell^{\textsc{Std}})$, and yet treatment is safe for women but is harmful for men. Deploying such a treatment would be ethically problematic, but this concern does not emerge under standard loss (and hence classical A/B testing).
Despite its ability to express one’s rich preferences, the main limitation of counterfactual loss is identifiability since principal strata are never observed [Dawid, 2000]. As a result, the literature primarily focused upon partial identification [e.g., Li and Pearl, 2019, 2024a, Ben-Michael et al., 2024].
2.1.3 Additive Counterfactual Loss
Decision
Negative $(D=0$) Positive $(D=1$)
Negative ($Y(0)=0$) False Negative (FN) True Positive
$\ell_{0}$ $\tilde{\ell}_{0}+c_{1}$
Positive ($Y(0)=1$) True Negative False Positive (FP)
$\ell_{1}$ $\tilde{\ell}_{1}+c_{1}$
Potential Outcomes Negative ($Y(1)=0$) True Negative False Positive (FP)
$\tilde{\ell}_{0}$ $\ell_{0}+c_{1}$
Positive ($Y(1)=1$) False Negative (FN) True Positive
$\tilde{\ell}_{1}$ $\ell_{1}+c_{1}$
Table 1: Confusion matrix for each combination of potential outcome $Y(d)=y_{d}$ and decision $D=d$. Each cell is assigned a loss $\ell(d;y_{0},y_{1})$ for $d,y_{0},y_{1}\in\{0,1\}$. Darker colors indicate realized outcomes, while lighter colors indicate counterfactual outcomes.
We propose an alternative strategy to non-identifiability by restricting the class of counterfactual losses to those that are additive in the potential outcomes. We now provide an example of additive counterfactual loss by extending a counterfactual loss considered in the literature [Coston et al., 2020, Ben-Michael et al., 2025] that depends on the baseline potential outcome $Y(0)$ alone. Under this formulation, the decision is a false negative if the physician does not provide the treatment and the patient dies, $(D,Y(0))=(0,0)$, in which case the loss is $\ell_{0}$ since $Y(0)$ is the realized outcome in this case. Conversely, a false positive occurs when the physician provides the treatment even though the patient would have survived without it, $(D,Y(0))=(1,1)$. In this case, $Y(0)$ is counterfactual, and we denote the associated loss as $\tilde{\ell}_{1}$. Under this setup, Ben-Michael et al. [2025] propose the following counterfactual classification loss:
$\ell(D;Y(0))=(1-D)(1-Y(0))\ell_{0}+D\cdot Y(0)\tilde{\ell}_{1}+c_{D}.$ (5)
As illustrated in Table 1, this counterfactual loss can be further generalized. First, we may also incorporate losses based on the potential outcome under treatment, covering cases where $(D,Y(1))=(0,1)$ (false negative) or $(D,Y(1))=(1,0)$ (false positive). Second, we may consider rewards (negative losses) for correct decisions: $(D,Y(0))=(1,0)$ and $(D,Y(1))=(1,1)$ are true positives, while $(D,Y(0))=(0,1)$ and $(D,Y(1))=(0,0)$ are true negatives. Letting $\tilde{\ell}_{0}$ denote the loss incurred when a patient would have died under the alternative (unchosen) decision, a more general counterfactual loss can be written as:
$\ell^{\textsc{Add}}(D;Y(0),Y(1))=\ell_{Y(D)}+\tilde{\ell}_{Y(1-D)}+c_{D}.$ (6)
This formulation is additive in the potential outcomes. While it does not fully capture stratum-specific interactions, additive counterfactual loss does account for the notion of counterfactual regret. For example, if $\tilde{\ell}_{0}<\tilde{\ell}_{1}$, this reflects a regret for not choosing the decision that would save the patient.
Despite these differences, as shown later, in the binary decision case, every additive counterfactual loss has an equivalent standard loss. For example, the additive counterfactual loss in Equation (6) is equivalent to $\ell^{\textsc{Equ}}(D;Y(D))=\ell_{Y(D)}-\tilde{\ell}_{Y(D)}+c_{D}$, because we have
$\ell^{\textsc{Add}}(D;Y(0),Y(1))-\ell^{\textsc{Equ}}(D;Y(D))=\tilde{\ell}_{Y(1-D)}+\tilde{\ell}_{Y(D)}=\tilde{\ell}_{Y(0)}+\tilde{\ell}_{Y(1)},$
which does not depend on the decision. Hence, $R(d;\ell^{\textsc{Equ}})$ and $R(d;\ell^{\textsc{Add}})$ induce the same ordering of treatment rules. Therefore, to highlight the material differences between standard and additive counterfactual losses, we next turn to a trichotomous decision example.
2.2 Trichotomous Decision
To illustrate counterfactual additive loss in the case of trichotomous decision, we introduce a different stylized example taken from criminal justice. Consider a first appearance hearing, where a judge must decide whether to release an arrestee on their own recognizance or ROR ($D=0$), impose supervision (e.g., monitoring via an electronic device) as a condition of release ($D=1$), or detain the arrestee until trail ($D=2$). As before, let $Y$ be the binary outcome of interest, where $Y=0$ indicates a negative outcome (e.g., the arrestee commits a crime before trial) and $Y=1$ represents a positive outcome (e.g., no new crime). For simplicity, we assume that detention prevents crime, i.e., $Y(2)=1$, though this assumption can be relaxed. We also assume no covariates are observed.
2.2.1 Standard Loss
Suppose that a judge wishes to balance the cost of crime incurred by the public against the costs of supervision and incarceration incurred, at least in part, by arrestees. Let $\ell_{y}$ denote the loss associated with outcome $Y=y$. Since we would like to avoid crime, we assume $\ell_{1}<\ell_{0}$. Let $c_{d}$ be the cost of decision $d$. The cost of detention is assumed to be greater than that of supervision, which in turn is larger than the cost of ROR, i.e., $c_{0}<c_{1}<c_{2}$. The corresponding standard loss is
$\ell^{\textsc{Std}}(D;Y(D))=\ell_{Y(D)}+c_{D}.$
2.2.2 Counterfactual Loss
This standard loss does not account for two important considerations a judge may wish to consider; harm to the public and harm to the arrestee. Harm to the public occurs when a judge’s decision caused a crime, whereas harm to the arrestee arises when a judge’s decision imposed an unnecessary cost on the arrestee through supervision or detention. It is reasonable to assume that a judge wishes to minimize both types of harms.
The following counterfactual loss incorporates the notions of these harms,
$\ell^{\textsc{Cf}}(D;Y(0),Y(1))$ (7)
$=$ $\ell_{Y(D)}+c_{D}+\mathds{1}{\left\{D=0\right\}}\,(1-Y(0))Y(1)p_{0}+\mathds{1}{\left\{D=1\right\}}\,Y(0)(1-Y(1))p_{1}+\sum_{k<d}r_{dk}Y(k),$
where $p_{d}$ is the additional penalty for causing harm to the public with decision $D=d$ for $d=0,1$ (recall that $D=2$ is assumed to lead to no crime), and $r_{dk}$ represents the regret for imposing an unnecessarily harsh decision $d$ when a more lenient decision $k<d$ would lead to no crime.
2.2.3 Additive Counterfactual Loss
Since the additional penalty term for the harm to the public depends on joint potential outcomes, its expected values are not identifiable. As a result, the counterfactual risk based on Equation (7) is not identifiable. However, the regret term for the harm to the arrestee does not involve joint potential outcomes and hence can be incorporated into an additive counterfactual loss in the following manner,
$\ell^{\textsc{Add}}(D;Y(0),Y(1))$ $=\ell_{Y(D)}+c_{D}+\sum_{k<d}r_{dk}Y(k).$ (8)
2.2.4 A Numerical Example
To further build intuition, we discuss a numerical illustration with the following setup. The crime losses are normalized as $\ell_{0}=1$ and $\ell_{1}=0$. Suppose that the cost of release on their own recognizance is negligible ($c_{0}=0$), whereas release with supervision incurs a small loss ($c_{1}=0.1$). Detention is assumed to be the most costly ($c_{2}=0.3$). We set the regret of unnecessary detention to $r_{20}=r_{21}=r$, and that of unnecessary supervision to $r_{10}=r/10$. Lastly, we assume a symmetric penalty $p_{0}=p_{1}=p$ so that harm of release to the public equals harm of supervised release to the public.
Since detention leads to no crime ($Y(2)=1$), the population consists of four principal strata: those who never commit crime $(Y(0),Y(1),Y(2))=(1,1,1)$), those who commit crime only if released on their own recognizance ($(Y(0),Y(1),Y(2))=(0,1,1)$), those who commit crime only if released under supervision ($(Y(0),Y(1),Y(2))=(1,0,1)$), who commit crime only if supervised, and those who commit crime unless detained ($(Y(0),Y(1),Y(2))=(0,0,1)$). For this example, we set the population shares of the above principal strata to 50, 20, 10, and 20%, respectively.
Figure 1: Numerical example illustrating the decision boundaries under the standard, additive counterfactual, and fully counterfactual loss. The standard loss gives the same decision recommendation as it ignores harm to the public $p$ and harm to the arrestee $r$. In contrast, under the counterfactual loss functions, the decision recommendation varies.
Figure 1 shows the decision recommendations under the three loss functions considered in this subsection. Under the standard loss, detention is always optimal as both harms are ignored. The additive loss incorporates harm to the arrestee, leading to release on their own recognizance when $r$ is sufficiently large. The fully counterfactual loss accounts for both types of harm, and recommends alternative decisions depending on their relative magnitude. The examle illustrates the three loss functions can yield different decision recommendations. In comparison to the standard loss, the counterfactual loss functions incorporate the notions of harm and yield a richer analysis.
3 Identification Analysis under Strong Ignorability
We now present our general theoretical results. Under the strong ignorability assumption, we show that additivity of counterfactual loss is a necessary and sufficient condition for the nonparametric identification of counterfactual risk. We further show that when the decision is binary, a counterfactual loss can be equivalently written as a standard loss although this result does not hold when the decision is non-binary. Finally, we develop a symbolic algorithm that can be used to verify the identification of counterfactual risk.
3.1 Setup
Consider an independently and identically distributed (IID) sample of size $n$ drawn from a super-population. For each unit $i=1,\ldots,n$, we observe a vector of pre-treatment covariates $\bm{X}_{i}\in\mathcal{X}$, a discrete decision $D_{i}\in\mathcal{D}:=\{0,1,\ldots,K-1\}$ with $K\geq 2$, and an outcome $Y_{i}\in\mathcal{Y}:=\{0,1,\ldots,M-1\}$ with $M\geq 2$. Let $Y_{i}(d)\in\mathcal{Y}$ denote the potential outcome of unit $i$ under decision $d\in\mathcal{D}$.
While the observed data are generated under the decision $D_{i}$, our goal is to evaluate the quality of a generic decision $D_{i}^{\ast}\in\mathcal{D}$, which may differ from $D_{i}$. For example, $D_{i}^{\ast}$ may correspond to an individualized treatment rule $D_{i}^{\ast}=\pi(\bm{X}_{i})$ for some policy function $\pi:\mathcal{X}\to\mathcal{D}$. More generally, we allow rules to depend on both observed and unobserved covariates, $D_{i}^{\ast}=\pi(\bm{X}_{i},\bm{U}_{i})$ for some possibly unknown $\pi:\mathcal{X}\times\mathcal{U}\to\mathcal{D}$, where $\bm{U}_{i}\in\mathcal{U}$ denotes unobserved covariates. Alternatively, $D_{i}^{\ast}$ may arise from a stochastic policy, where $\Pr(D_{i}^{\ast}=d)=\pi_{d}(\bm{X}_{i},\bm{U}_{i})$ for $d\in\mathcal{D}$ and $\sum_{d=0}^{K-1}\pi_{d}(\bm{X}_{i},\bm{U}_{i})=1$. Crucially, we do not require the knowledge of mechanism generating $D_{i}^{\ast}$ although we assume that $D_{i}^{\ast}$ is observed for each unit. In contrast, the associated potential outcome $Y_{i}(D_{i}^{\ast})$ remains unobserved unless $D_{i}^{\ast}=D_{i}$. Throughout this paper, we make the following assumptions.
Assumption 1 (IID Sampling)
The tuple $\{Y_{i}(0),\ldots,Y(K-1),D_{i},D_{i}^{\ast},\bm{X}_{i}\}_{i=1}^{n}$ is an independently and identically distributed sample from an unknown superpopulation.
We treat potential outcomes as iid draws from an unknown superpopulation but assume that they are fixed given any unit [Neyman, 1923, Rubin, 1974]. This contrasts with the stochastic potential outcomes [Greenland, 1987, VanderWeele and Robins, 2012], which remain random even at the unit level. Nevertheless, our results hold under either assumption.
Assumption 2 (Consistency)
For all $i$, the observed outcome satisfies $Y_{i}=Y_{i}(D_{i})$. Moreover, if $D_{i}^{\ast}=D_{i}$, then $Y_{i}(D_{i}^{\ast})=Y_{i}(D_{i})=Y_{i}$.
We study the identifiability of counterfactual risk under the standard strong ignorability assumptions [Rosenbaum and Rubin, 1983].
Assumption 3 (Strong Ignorability)
For $i=1,\ldots,n$, the following assumptions hold,
  1. (a)
    Unconfoundedness: $D_{i}\mbox{$\perp\!\!\!\perp$}(D_{i}^{\ast},\{Y_{i}(d)\}_{d\in\mathcal{D}})\mid\bm{X}_{i}$.
  2. (b)
    Overlap: there exists $\eta>0$ such that, for all $d\in\mathcal{D}$, $\eta<\Pr(D_{i}=d\mid\bm{X}_{i})<1-\eta$.
Assumption 3(a) implies that, conditional on covariates, the observed decision $D_{i}$ is independent of both the potential outcomes and the decision rule $D_{i}^{\ast}$. Therefore, $D_{i}^{\ast}$ may depend on unobserved covariates, and hence be correlated with potential outcomes. Assumption 3(b) states that every decision category in $\mathcal{D}$ occurs with nonzero probability given covariates. While this assumption needs to be satisfied by the observed decision $D_{i}$, it is not required for the decision rule $D_{i}^{\ast}$.
3.2 Counterfactual Loss and Risk
We define a loss (or negative utility) associated with a decision as a function of the covariates and all potential outcomes, where a lower loss value indicates a better decision. We refer to this as the counterfactual loss, as it depends on the full vector of potential outcomes. This contrasts with the standard loss, which is defined as a function of the realized outcome alone.
Definition 1 (Counterfactual Loss)
The counterfactual loss is defined as a function $\ell:\mathcal{D}\times\mathcal{Y}^{K}\times\mathcal{X}\to\mathbb{R}$, where $\ell(d;y_{0},\ldots,y_{K-1},\bm{x})$ denotes the loss of choosing decision $D_{i}^{\ast}=d$ for a unit with potential outcomes $(Y_{i}(0),\ldots,Y_{i}(K-1))=(y_{0},\ldots,y_{K-1})$ and covariates $\bm{X}_{i}=\bm{x}$. We further impose that the loss is integrable with respect to the covariates, i.e., $\mathbb{E}\left[\sup_{d\in\mathcal{D}}\left|\ell(d;Y_{i}(0),\ldots,Y_{i}(K-1),\bm{X})\right|\right]<\infty$.
This integrability condition is trivially satisfied when the loss does not depend on covariates, since the potential outcomes are discrete and finite.
Next, we define the counterfactual risk, or the expected counterfactual loss, which serves as the primary estimand of interest. We also define the conditional counterfactual risk, which conditions on the observed covariates.
Definition 2 (Counterfactual Risk and Conditional Counterfactual Risk)
Given the counterfactual loss $\ell$, the counterfactual risk $R(D^{\ast};\ell)$ of decision $D^{\ast}$ is given by,
$R(D^{\ast};\ell):=\mathbb{E}{\left[\ell(D^{\ast};Y(0),\ldots,Y(K-1),\bm{X})\right]}=\mathbb{E}{\left[R_{\bm{X}}(D^{\ast};\ell)\right]},$
where $R_{\bm{x}}(D^{\ast};\ell)$ is the conditional counterfactual risk given $\bm{X}=\bm{x}$,
$R_{\bm{x}}(D^{\ast};\ell):=\sum_{d\in\mathcal{D}}\sum_{\{y_{k}\}_{k=0}^{K-1}\in\mathcal{Y}^{K}}\ell(d;y_{0},\ldots,y_{k-1},\bm{x})\cdot\Pr(D^{\ast}=d,Y(0)=y_{0},\ldots,Y(K-1)=y_{K-1}\mid\bm{X}=\bm{x}).$
Definition 2 follows the literature and take the expectation with respect to the joint distribution of $\{D^{\ast},Y(0),\ldots,Y(K-1),\bm{X}\}$ [e.g., Manski, 2004, Li and Pearl, 2019, Coston et al., 2020, Mueller and Pearl, 2023, Ben-Michael et al., 2024]. That is, we integrate over the unknown joint distribution of potential outcomes rather than conditioning on them. Although it is possible to incorporate expert knowledge by placing a prior distribution on the potential outcomes, in this paper we remain agnostic about their distribution (as well as the decision to be evaluated, $D^{\ast}$). Like Bayesians, we integrate over uncertainty; however, we follow a frequentist tradition by relying entirely on the data-generating process and avoiding the use of a prior distribution.
3.3 Additivity
Our goal is to study the identifiability of the counterfactual risk. The challenge arises from the fact that, for each unit, we observe only a single realized outcome among the $K$ possible potential outcomes, even though the counterfactual loss depends on the entire set of potential outcomes. To overcome this identification problem, we consider the following class of counterfactual losses that are additive in the potential outcomes.
Definition 3 (Additive Counterfactual Loss)
A counterfactual loss is additive if it allows the decomposition,
$\ell^{\textsc{Add}}(d;\bm{y},\bm{x}):=\varpi(\bm{y},\bm{x})+\sum_{k\in\mathcal{D}}\omega_{k}(d,y_{k},\bm{x}),$
where $\bm{y}=(y_{0},\ldots,y_{K-1})$, and we call $\omega_{k}:\mathcal{D}\times\mathcal{Y}\times\mathcal{X}\rightarrow\mathbb{R}$ the weight function and $\varpi:\mathcal{Y}^{K}\times\mathcal{X}\rightarrow\mathbb{R}$ the intercept function.
The additive counterfactual loss equals the sum of two terms; the intercept function does not depend on the decision $D^{\ast}=d$, while the weight function does. Moreover, each weight function $\omega_{k}$ depends only on the corresponding potential outcomes $Y_{i}(k)$ given the covariate, whereas the intercept function may include all possible interactions of potential outcomes. Although we cannot identify the intercept function, we can ignore it when comparing alternative decision rules because it does not depend on the decision.
We now illustrate Definition 3 by revisiting the examples shown in Section 2.
Example 1 (Counterfactual Classification Loss)
The counterfactual loss function given in Equation (5) is additive in the potential outcomes with $\omega_{0}(0,0)=\ell_{0}$, $\omega_{0}(1,1)=\tilde{\ell}_{1}+c_{1}$, $\omega_{1}(d,y_{1})=0$ for $d,y_{1}\in\{0,1\}$, and $\varpi(\bm{y})=0$. An extension of this counterfactual loss given in Equation (6) is also additive with $\omega_{d}(d,y_{d})=\ell_{y_{d}}+c_{d}$, $\omega_{1-d}(d,y_{1-d})=\tilde{\ell}_{y_{1-d}}$ for $d,y\in\{0,1\}$, and $\varpi(\bm{y})=0$.
Example 2 (Asymmetric Counterfactual Loss)
The asymmetric counterfactual loss given in Equation (4) is an example of non-additive loss whenever $\ell_{0}^{\text{R}}-\ell_{1}^{\text{R}}\neq\ell_{0}^{\text{H}}-\ell_{1}^{\text{H}}$ (see Equation (15) in Section 3.8 for details). When $\ell_{0}^{\text{R}}-\ell_{1}^{\text{R}}=\ell_{0}^{\text{H}}-\ell_{1}^{\text{H}}=\Delta$, the loss is additive with $\omega_{d}(d,y_{d})=\ell_{y_{d}}+c_{d}$, $\omega_{1-d}(d,y_{1-d})=\{\Delta-(\ell_{0}-\ell_{1})\}y_{1-d}$, and $\varpi(\bm{y})=(\ell_{1}^{\text{H}}-\ell_{1})y_{0}(1-y_{1})+(\ell_{1}^{\text{R}}-\ell_{1})(1-y_{0})y_{1}-\{\Delta-(\ell_{0}-\ell_{1})\}y_{0}y_{1}$. Note that $\Delta-(\ell_{0}-\ell_{1})$ is the difference in loss between units for whom the treatment matters and those for whom it does not.
Example 3 (Trichotomous decision)
The counterfactual loss for individual harm, given in Equation (8), is additive in the potential outcomes with $\omega_{d}(d,y_{d})=\ell_{y_{d}}+c_{d}$, $\omega_{k}(d,y_{k})=r_{dk}y_{k}\mathds{1}{\left\{k<d\right\}}$ for $d,k\in\{0,1,2\},y_{d},y_{k}\in\{0,1\}$, and $\varpi(\bm{y})=0$. By contrast, the extended counterfactual loss given in Equation (7) that incorporates both harm to the public and harm to the arrestee is additive if and only if $p_{0}=p_{1}=0$, in which case it reduces to Equation (8).
3.4 Nonparametric Identification
We now investigate the conditions under which the counterfactual risk is identifiable, focusing on the class of additive counterfactual loss functions introduced above. We adopt the following general definition of identification from Basse and Bojinov [2020]:
Definition 4 (Identifiability)
Let $\mathcal{P}$ denote the set of all joint probability distributions under consideration, and let $\mathcal{Q}$ be the set of all observable distributions. Let $Q:\mathcal{P}\to\mathcal{Q}$ denote a surjective mapping that associates each joint distribution with an observable distribution. Let $\theta(P)$ denote the (causal) estimand of interest under distribution $P\in\mathcal{P}$. For a fixed observable distribution $Q_{0}\in\mathcal{Q}$ define $\mathcal{S}_{0}=\{P\in\mathcal{P}:Q(P)=Q_{0}\}.$ We say that $\theta$ is identifiable at $Q_{0}$ with respect to $Q$ if there exists a constant $\theta_{0}$ such that $\theta(P)=\theta_{0}$ for all $P\in\mathcal{S}_{0}$. We say that $\theta$ is identifiable with respect to $Q$ if it is identifiable at every $Q_{0}\in\mathcal{Q}$.
According to the definition, a causal estimand is identifiable if it takes the same value across all joint probability measures that map to any given observable distribution. Surjectivity of $Q$ ensures that this property holds for every possible observable distribution.
In our setting, $\mathcal{P}$ denotes the set of all joint probability distributions $P(D^{\ast},D,Y(0),\ldots,Y(K-1),\bm{X})$ that satisfy Assumptions 13. The causal estimand of interest is the counterfactual risk $R(D^{\ast};\ell)$, which, according to Definition 2, is fully determined by $P\in\mathcal{P}$. Let $Q(P)=P(D^{\ast},D,Y,\bm{X})$ denote the surjective mapping that yields the observable distribution from each $P\in\mathcal{P}$. Then, according to Definition 4, the counterfactual risk $R(D^{\ast};\ell)$ is identifiable with respect to $Q$ if it can be expressed as a function of the observable distribution $P(D^{\ast},D,Y,\bm{X})$.
As formally shown in Lemma 2 (see Appendix G.2), since $P(\bm{X})$ is observable, the counterfactual risk $R(D^{\ast};\ell)$ is identifiable with respect to $P(D^{\ast},D,Y,\bm{X})$ if and only if the conditional counterfactual risk $R_{\bm{X}}(D^{\ast};\ell)$ is identifiable with respect to $P(D^{\ast},D,Y\mid\bm{X})$. Since $R_{\bm{X}}(D^{\ast};\ell)$ depends on the conditional distribution $P(D^{\ast},Y(0),\ldots,Y(K-1)\mid\bm{X})$, we need to express this distribution of potential outcomes and decision in terms of observable distributions. Under Assumption 3, we can only identify the joint distribution of the decision $D^{\ast}$ and a single potential outcome given covariates,
$\Pr(D^{\ast}=d,Y(k)=y_{k}\mid\bm{X})\ =$ $\ \Pr(D^{\ast}=d,Y(k)=y_{k}\mid D=k,\bm{X})$
$=$ $\ \Pr(D^{\ast}=d,Y=y_{k}\mid D=k,\bm{X}),$
for any $k\in\mathcal{D}$. In particular, without additional assumptions, we cannot identify a joint distribution that involves multiple potential outcomes. Thus, it follows that $R_{\bm{x}}(D^{\ast};\ell)$ is identifiable if and only if it can be expressed solely in terms of the distribution $\Pr(D^{\ast}=d,Y(k)=y\mid\bm{X}=\bm{x})$ for any $d,k\in\mathcal{D},y\in\mathcal{Y}$. This observation suggests that the additivity assumption given in Definition 3 can facilitate identification.
We first show that under an additive loss, we can identify the difference in conditional counterfactual risk between any two decision rules even though each conditional counterfactual risk is only identifiable up to an unknown additive constant. That is, the conditional counterfactual risk of decision rule $D^{\ast}$ can be decomposed into an identifiable part and an unidentifiable part, the latter does not depend on $D^{\ast}$.
Theorem 1 (The Conditional Counterfactual Risk under the Additive Loss)
Suppose that a counterfactual loss function takes the additive form given in Definition 3. Then, under Assumptions 13, the conditional counterfactual risk can be decomposed as,
$R_{\bm{x}}(D^{\ast};\ell^{\textsc{Add}})=\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D}}\sum_{y\in\mathcal{Y}}\omega_{k}(d,y,\bm{x})\Pr(D^{*}=d,Y(k)=y\mid\bm{X}=\bm{x})+C(\bm{x}),$
where
$C(\bm{x})=\sum_{\bm{y}\in\mathcal{Y}^{K}}\varpi(\bm{y},\bm{x})\Pr(\bm{Y}(\mathcal{D})=\bm{y}\mid\bm{X}=\bm{x}),$
and $\bm{Y}(\mathcal{D})=(Y(0),\ldots,Y(K-1))$. Thus, the difference in the conditional counterfactual risk between any pair of decision rules can be identified.
The proof is given in Appendix A.
Taking the expectation of the expression in Theorem 1 over the distribution of covariates, we find that the counterfactual risk is identified up to an unknown additive constant $\mathbb{E}[C(\bm{X})]$. We state this result as the following corollary without proof.
Corollary 1 (The Counterfactual Risk under the Additive Loss)
Suppose that a counterfactual loss function takes the additive form given in Definition 3. Then, under Assumptions 13, the counterfactual risk can be decomposed as,
$R(D^{\ast};\ell^{\textsc{Add}})=\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D}}\sum_{y\in\mathcal{Y}}\mathbb{E}\left[\omega_{k}(d,y,\bm{X})\Pr(D^{*}=d,Y=y\mid D=k,\bm{X})\right]+\mathbb{E}\left[C(\bm{X})\right].$
Thus, the difference in counterfactual risk between any pair of decision rules is identified.
Theorem 1 immediately implies that one can find an optimal decision, e.g., $\pi:\mathcal{X}\rightarrow\mathcal{D}$, by minimizing the identifiable term alone, i.e.,
$\pi^{\text{opt}}\ \in$ $\ \operatornamewithlimits{argmin}_{\pi\in\Pi}R(\pi(\bm{X});\ell^{\textsc{Add}})$
$=$ $\ \operatornamewithlimits{argmin}_{\pi\in\Pi}\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D}}\sum_{y\in\mathcal{Y}}\mathbb{E}\left[\omega_{k}(d,y,\bm{X})\Pr(\pi(\bm{X})=d,Y=y\mid D=k,\bm{X})\right],$
where $\Pi$ is a class of decision rules to be considered. Moreover, if the constant term is zero, then we can exactly identify the counterfactual risk. This happens, for example, when the intercept function in the additive counterfactual loss (Definition 3) is zero, i.e., $\varpi(\bm{y},\bm{x})=0$ for all $\bm{y}\in\mathcal{Y}^{K}$ and $\bm{x}\in\mathcal{X}$. Lastly, by placing assumptions on $\varpi$, we can control the magnitude of $\mathbb{E}{\left[C(\bm{X})\right]}$, which leads to partial identification results.
A natural question arises as to whether or not the above identification results can be obtained under a different form of counterfactual loss function. The next theorem shows that the additivity of loss function given in Definition 3 is a necessary and sufficient condition for the identification of the difference in counterfactual risk. In other words, every counterfactual risk that is identifiable up to an additive constant must have an additive loss.
Theorem 2 (Additivity as Necessary and Sufficient Condition for Identification)
The difference in counterfactual risk between any pair of decision-making systems can be identified under Assumptions 13 if and only if the corresponding loss takes the additive form given in Definition 3.
The proof is given in Appendix B.
Finally, as a corollary of this identification result, we show that the exact identification of the counterfactual risk is possible if and only if there exists an equivalent counterfactual loss that sets the intercept function $\varpi$ to zero (Definition 3).
Corollary 2 (Necessary and Sufficient Condition for Exact Identification)
The counterfactual risk given in Definition 2 can be exactly identified under Assumptions 13 if and only if the corresponding loss takes the following form,
$\ell^{\textsc{Add}}(d;\bm{y},\bm{x})=\sum_{k\in\mathcal{D}}\omega_{k}(d,y_{k},\bm{x}),$
for a weight function $\omega_{k}:\mathcal{D}\times\mathcal{Y}\times\mathcal{X}\rightarrow\mathbb{R}$. The identification formula is given by,
$R(D^{*},\ell^{\textsc{Add}})$ $=\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D}}\sum_{y\in\mathcal{Y}}\mathbb{E}\left[\omega_{k}(d,y,\bm{X})\Pr(D^{\ast}=d,Y=y\mid D=k,\bm{X})\right].$
The proof is given in Appendix C.
Theorem 2 shows that additive counterfactual losses represent the largest class of counterfactual losses that are identifiable under Assumptions 13. In particular, separability alone is not sufficient for identification; what is required is linearity in the potential outcomes. The counterfactual risks associated with this class never depend on the joint distribution of potential outcomes in decision-relevant terms. Moreover, this class of counterfactual losses is consistent with the decision-analytic framework of Dawid [2000], since the resulting quantity of interest depends only on marginal distributions of potential outcomes.
3.5 When Does the Additive Counterfactual Loss Differ from the Standard Loss?
While the additivity of counterfactual loss is a necessary and sufficient condition for identification of counterfactual risk, an important question is whether or not an additive counterfactual risk yields an optimal decision rule that is different from the one based on the standard risk, which is also identifiable from the observed data. If they always yield the same optimal decision rule, an additive counterfactual risk is not different from the standard risk in meaningful ways.
We first show that when the decision is binary ($K=2$), any additive counterfactual loss admits an equivalent standard loss. That is, for any collection of decision rules, an additive counterfactual risk and its associated standard risk always give the same ordering among the decision rule. However, this no longer holds when the decision is non-binary, as long as the counterfactual weights have sufficient interactions between decision and counterfactual potential outcomes.
Proposition 1 (Additive Counterfactual Risk with Binary Decision)
Suppose that the decision is binary, i.e., $\mathcal{D}=\{0,1\}$ and Assumptions 13 hold.
  1. 1.
    For any additive counterfactual loss $\ell^{\textsc{Add}}(d;\bm{y},\bm{x})$ defined in Definition 3, we can construct a standard loss $\ell^{\textsc{Std}}(d;y_{d},\bm{x})$ such that the risk difference $R(D^{\ast};\ell^{\textsc{Add}})-R(D^{\ast};\ell^{\textsc{Std}})$ does not depend on $D^{\ast}$. Moreover, this standard loss is unique up to a constant, which only depends on covariates, and takes the following form,
    $\ell^{\textsc{Std}}(d;y_{d},\bm{x})=\omega_{d}(d,y_{d},\bm{x})-\omega_{d}(1-d,y_{d},\bm{x})+c(\bm{x}).$
  2. 2.
    While the mapping from the space of counterfactual losses to standard losses that yield the same treatment recommendations is surjective, it is not injective. That is, for any standard loss, there exists an infinite number of counterfactual losses that give the same treatment recommendations.
The proof is given in Appendix D. This proposition implies that in the binary decision setting, any additive counterfactual risk admits a standard risk that yields identical treatment recommendations. As a consequence, both of these risks lead to the same optimal policy. Intuitively, when the decision is binary, the unchosen decision 1d1-d is a deterministic function of the chosen one $d$, making it impossible to generate an identifiable counterfactual loss under additivity. Perhaps surprisingly, this result is not limited to binary outcomes. The corresponding standard loss $\ell^{\textsc{Std}}(d;y_{d})$ is given by the difference between the weight for the realized outcome $Y(D)=y_{d}$ under the chosen decision $D^{\ast}=d$ and the weight for the same outcome $Y(D)=y_{d}$ under the alternative decision $D^{\ast}=1-d$ (up to an additive constant that only depends on covariates).
Although every additive counterfactual loss has a unique corresponding standard loss (up to an additive constant) with the same treatment recommendations, the reverse does not hold. That is, any given standard loss has infinitely many additive counterfactual losses with the same treatment recommendations. Each of such counterfactual losses assigns different values to principal strata. This asymmetry explains why, even in the binary case, additive counterfactual loss has a straightforward interpretation based on principal strata.
We now revisit our binary decision examples given in Section 3.3.
Example 1 (Counterfactual Classification Loss)
The counterfactual loss in Equation (6) can be expressed as a standard loss, $\ell^{\textsc{Std}}(d;y_{d})=\ell_{y_{d}}-\tilde{\ell}_{y_{d}}+c_{d}$, which corresponds to the difference between the realized and counterfactual loss associated with outcome $y_{d}$ with the additional cost term. Thus, the regret is absorbed into the standard loss. If $\tilde{\ell}_{0}<\tilde{\ell}_{1}$ (see Section 2.1.3), the loss difference between correct and incorrect decisions is amplified, since $(\ell_{0}-\tilde{\ell}_{0})-(\ell_{1}-\tilde{\ell}_{1})>\ell_{0}-\ell_{1}$.
Example 2 (Asymmetric Counterfactual Loss)
When $\ell_{0}^{\text{R}}-\ell_{1}^{\text{R}}=\ell_{0}^{\text{H}}-\ell_{1}^{\text{H}}=\Delta$ holds, the loss in Equation (4) can be rewritten as a standard loss with $\ell^{\textsc{Std}}(d;y_{d})=\ell_{y_{d}}+c_{d}-\{\Delta-(\ell_{0}-\ell_{1})\}y_{d}\propto\ell^{\text{H}}_{y_{d}}+c_{d}\propto\ell^{\text{R}}_{y_{d}}+c_{d}$. Thus, the loss reduces to the standard loss in Equation (3).
Next, we show that when the decision is non-binary, no standard loss corresponds directly to the additive counterfactual loss, as long as the counterfactual weight is not additively separable in decision and counterfactual outcome.
Proposition 2 (Additive Counterfactual Risk with Non-binary Decision)
Assume that the decision is non-binary, i.e., $K=|\mathcal{D}|\geq 3$ and Assumptions 13 hold. Consider any additive counterfactual loss $\ell^{\textsc{Add}}(d;\bm{y},\bm{x})$ given in Definition 3. Then, there exists no standard loss $\ell^{\textsc{Std}}(d;y_{d},\bm{x})$ such that the risk difference $R(D^{\ast};\ell^{\textsc{Add}})-R(D^{\ast};\ell^{\textsc{Std}})$ does not depend on $D^{\ast}$ if and only if at least one weight function $\omega_{k}(d,y,\bm{x})$ is not off-diagonal additively separable in $d\neq k$ and $y\in\mathcal{Y}$, i.e., there do not exist functions $f_{k}$ and $g_{k}$ such that
$\omega_{k}(d,y,\bm{x})=f_{k}(d,\bm{x})+g_{k}(y,\bm{x}),$
for all $d\in\mathcal{D}\setminus\{k\},y\in\mathcal{Y}$. If all the weight functions are off-diagonal additively separable, then an equivalent standard loss is given by
$\ell^{\textsc{Std}}(d;,y_{d},\bm{x})=\omega_{d}(d,y_{d},\bm{x})-g_{d}(y_{d},\bm{x})+\sum_{k\in\mathcal{D},k\neq d}f_{k}(d,\bm{x}),$
which is unique up to a constant depending only on covariates.
The proof is given in Appendix E. Proposition 2 shows that in the non-binary setting, standard losses are inadequate when the interactions between decision and counterfactual outcomes matter. In the binary decision case, one weight function is always additively separable. Indeed, since $\mathcal{D}\setminus\{k\}=\{1-k\}$ is a singleton for $\mathcal{D}=\{0,1\}$, we can set $f_{k}(1-k,\bm{x})=0$ and $g_{k}(y,\bm{x})=\omega_{k}(1-k,y,\bm{x})$ to recover Proposition 1.
Example 3 (Trichotomous decision)
The additive counterfactual loss given in Equation (8) is not off-diagonal additively separable. Indeed, the weight function $\omega_{k}(d,y_{k})=\mathds{1}{\left\{d=k\right\}}(\ell_{y_{d}}+c_{d})+\mathds{1}{\left\{d\neq k\right\}}r_{dk}y_{k}\mathds{1}{\left\{k<d\right\}}$ is not separable in the decision and outcome for $d\neq k$ as long as $r_{dk}\neq 0$. Thus, no standard loss can take harm to the arrestee into account.
However, Propositions 1 and 2 also show that there do exist counterfactual losses depending on all potential outcomes that do not generalize the decision-making compared to losses only depending on the observed outcome. We will illustrate this somehow paradoxical behavior with an example.
Example 4 (Additively Separable Regret)
Consider the additive counterfactual loss $\ell(d;\bm{y})=\ell_{d,y_{d}}+\sum_{k\in\mathcal{D},k\neq d}\left(c_{k,d}+\tilde{\ell}_{k,y_{k}}\right)$ for $d\in\mathcal{D}$ and $\bm{y}\in\mathcal{Y}$, which generalizes Equation (6) to multi-valued decision and outcome settings. The corresponding weight function is given by $\omega_{k}(d,y)=\mathds{1}{\left\{d=k\right\}}\ell_{d,y}+\mathds{1}{\left\{d\neq k\right\}}\left(c_{k,d}+\tilde{\ell}_{k,y}\right)$, which is off-diagonal and additively separable with $f_{k}(d)=c_{k,d}$ and $g_{k}(y)=\tilde{\ell}_{k,y}$. By Proposition 2, an equivalent standard loss is $\ell^{\textsc{Std}}(d,y_{d})=\ell_{d,y_{d}}-\tilde{\ell}_{d,y_{d}}+\sum_{k\in\mathcal{D},k\neq d}c_{k,d}$. Computing the loss difference yields $\ell(d,\bm{y})-\ell^{\textsc{Std}}(d,y_{d})=\sum_{k\in\mathcal{D}}\tilde{\ell}_{k,y}$, which does not depend on the decision. Thus, even losses that assign a different regret value to each unrealized counterfactual outcome can be represented as a standard loss, so long as the regret only depends additively on the decision and counterfactual outcome. By Proposition 2, all counterfactual losses that can be written equivalently as standard losses have this form.
3.6 Case with Binary Outcome and Multi-valued Decision
To further build intuition for the above results, we study the special case of binary outcome and multi-valued decision, i.e., $\mathcal{Y}=\{0,1\}$ and $K\geq 3$. For notational simplicity, we ignore covariates. Without loss of generality, assume that $Y=1$ is a desirable outcome while $Y=0$ is undesirable. Under this setting, it is natural to consider the following ordering of weight functions in the additive counterfactual loss in Definition 3 for any given decision $d$:
$\omega_{d}(d,1)\leq 0\leq\omega_{d}(d,0),\quad\{\omega_{d^{\prime}}(d,0)\}_{d^{\prime}\neq d}\leq 0\leq\{\omega_{d^{\prime}}(d,1)\}_{d^{\prime}\neq d},$ (9)
where $\omega_{d}(d,y_{d})$ for $y_{d}=0,1$ accounts for the accuracy of decision, and $\{\omega_{d^{\prime}}(d,y_{d^{\prime}})\}_{d^{\prime}\neq d}$ for $y_{d^{\prime}}=0,1$ capture its difficulty.
We first discuss the accuracy component. The correct “true positive” decision, i.e., $(D^{\ast},Y(d))=(d,1)$ would decrease the counterfactual loss, which is represented by a negative loss of $\omega_{d}(d,1)$. In contrast, the incorrect “false negative” decision, i.e., $(D^{\ast},Y(d))=(d,0)$ would increase the loss, which corresponds to a positive loss of $\omega_{d}(d,0)$.
While accuracy is an essential part of the standard decision theory, difficulty is unique to the counterfactual loss. Avoidance of an incorrect decision (“true negative”), i.e., $(D^{\ast},Y(d))=(d^{\prime},0)$ results in a decrease of the loss, while failure to make a correct “true positive” decision, i.e., $(D^{\ast},Y(d))=(d^{\prime},1)$ increases the loss. Thus, the counterfactual loss is smaller when fewer alternative decisions would yield a desirable outcome (i.e., making a correct decision is difficult). The loss is greater when many alternatives would result in a desirable outcome (making a correct decision is easier). In this way, the counterfactual loss rewards a consequential decision that changes the outcome.
Lastly, since each weight function $\omega_{k}(d,y)$ depends on both the decision and the outcome, we can decompose it into a baseline decision term and an interaction component:
$\omega_{k}(d,y)=\omega_{k}(d,0)+\mathds{1}{\left\{y=1\right\}}\left(\omega_{k}(d,1)-\omega_{k}(d,0)\right).$
The first term corresponds to the baseline cost of making decision $D^{\ast}=d$ on the potential outcome $Y(k)$, while the second term accounts for the interaction of decision and potential outcome.
We formalize the above discussion as the following corollary of Theorem 1. Specifically, we show that the conditional counterfactual risk can be decomposed into four components — accuracy based on the realized outcome $\Pr(D^{\ast}=d,Y(d)=1\mid\bm{X}=\bm{x})$, difficulty based on counterfactual outcomes $\Pr(D^{\ast}=d,Y(k)=1\mid\bm{X}=\bm{x})$ for $k\neq d$, and baseline decision probability $\Pr(D^{\ast}=d\mid\bm{X}=\bm{x})$, and the intercept function $C(\bm{x})$. Each of these terms has a weight function based on the specified additive loss function.
Corollary 3 (Accuracy and Difficulty for the Binary Outcome Case)
Assume $Y\in\{0,1\}$. Suppose that the counterfactual loss function takes the additive form given in Definition 3. Then, under Assumptions 13, the conditional counterfactual risk can be written as,
$R_{\bm{x}}(D^{\ast};\ell^{\textsc{Add}})$
$=$ $\sum_{d\in\mathcal{D}}\zeta_{d}(d,\bm{x})\Pr(D^{\ast}=d,Y(d)=1\mid\bm{X}=\bm{x})+\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D},k\neq d}\zeta_{k}(d,\bm{x})\Pr(D^{\ast}=d,Y(k)=1\mid\bm{X}=\bm{x})$
$+\sum_{d\in\mathcal{D}}\xi(d,\bm{x})\Pr(D^{\ast}=d\mid\bm{X}=\bm{x})+C(\bm{x}).$
where $\zeta_{k}(d,\bm{x})=\omega_{k}(d,1,\bm{x})-\omega_{k}(d,0,\bm{x})$, $\xi(d,\bm{x})=\sum_{k\in\mathcal{D}}\omega_{k}(d,0,\bm{x})$, and $C(\bm{x})=\sum_{\bm{y}\in\{0,1\}^{K}}\varpi(\bm{y},\bm{x})\Pr(\bm{Y}(\mathcal{D})=\bm{y}\mid\bm{X}=\bm{x}).$
The proof is given in Appendix F. If we choose the weights such that Equation (9) holds, then we have $\zeta_{d}(d,\bm{x})\leq 0\leq\zeta_{k}(d,\bm{x})$ for all $k\neq d$. This implies that the risk decreases with greater accuracy, i.e., when the realized outcome is desirable, and increases with counterfactual regret, i.e., when other decisions also would yield a desirable outcome, reflecting the difficulty of the decision. The baseline decision term is a regularizer that penalizes decisions with higher baseline losses.
We next characterize when the additive counterfactual loss differs from the standard loss in the binary outcome case. This result is a corollary of Proposition 2. It shows that an additive counterfactual loss cannot be represented by a standard loss if and only if the former includes a decision-dependent difficulty term.
Corollary 4 (Decision-Dependent Difficulty for the Binary Outcome Case)
Assume $Y\in\{0,1\}$ and $|\mathcal{D}|\geq 3$. Then, under Assumptions 13, for any additive counterfactual loss $\ell^{\textsc{Add}}(d;\bm{y},\bm{x})$ defined in Definition 3, there exists a standard loss $\ell^{\textsc{Std}}(d;y_{d},\bm{x})$ such that the risk difference $R(D^{\ast};\ell^{\textsc{Add}})-R(D^{\ast};\ell^{\textsc{Std}})$ is independent of $D^{\ast}$ if and only if the slope
$\zeta_{k}(d,\bm{x})=\omega_{k}(d,1,\bm{x})-\omega_{k}(d,0,\bm{x})$
is independent of $d$ whenever $d\neq k$ for all $k\in\mathcal{D}$. In this case, an equivalent standard loss takes the form
$\ell^{\textsc{Std}}(d;y_{d},\bm{x})=y_{d}(\zeta_{d}(d,\bm{x})-\tilde{\zeta}_{d}(d,\bm{x}))+\xi(d,\bm{x}),$
where $\tilde{\zeta}_{d}(d,\bm{x}):=\omega_{d}(j,1,\bm{x})-\omega_{d}(j,0,\bm{x})$ for any $j\neq d$ and $\xi(d,\bm{x})=\sum_{k\in\mathcal{D}}\omega_{k}(d,0,\bm{x})$. This loss is unique up to a constant that depends only on covariates.
Thus, a necessary and sufficient condition for an additive counterfactual loss to differ from a standard loss is to have the regret weight function $\zeta_{k}(d,\bm{x})$ in Corollary 3 depend on the decision $d$.
3.7 A Symbolic Algorithm for Identification Verification
Theorem 2 establishes additivity as a necessary and sufficient condition for identifying counterfactual risk differences. Building on this result, we propose a symbolic algorithm that takes a counterfactual loss function as an input and tests whether the corresponding counterfactual risk is identifiable. The algorithm also yields the weight and intercept functions from Definition 3, leading to a decomposition of counterfactual risk into observable marginals. The proposed algorithm further clarifies the constraints that additivity imposes on principal strata.
We begin by noting the fact that the definition of additive counterfactual loss given in Definition 3 induces a system of linear equations. Let $\bm{\ell}=\{\ell^{\textsc{Add}}(d;y_{0},\ldots,y_{K-1})\}_{d\in\mathcal{D},\bm{y}\in\mathcal{Y}^{K}}\in\mathbb{R}^{KM^{K}}$ be the vector of all loss values, $\bm{w}^{\textsc{Add}}=\{\omega_{k}(d,y),\varpi(\bm{y})\}_{d\in\mathcal{D},k\in\mathcal{D},y\in\mathcal{Y},\bm{y}\in\mathcal{Y}^{K}}\in\mathbb{R}^{K^{2}M+M^{K}}$ denote the corresponding vector of weight and intercept functions, and $\bm{A}^{\textsc{Add}}$ represent the $\left|\bm{\ell}\right|\times\left|\bm{w}\right|$ matrix of zeros and ones determined by the additive structure. Then, any additive loss can be written as
$\bm{A}^{\textsc{Add}}\bm{w}^{\textsc{Add}}=\bm{\ell}.$ (10)
[…]
Corollary 5 (Identification Verification of Counterfactual Risk)
Suppose that Assumptions 13 hold. Consider a counterfactual loss defined in Definition 1. The corresponding counterfactual risk is identifiable up to a constant and admits the representation given in Corollary 1 if and only if the linear system given in Equation (10) has a solution.
The linear system in Equation (10) can be checked for a solution and solved using a standard linear solver whose time complexity is polynomial in $M$ but exponential in $K$. Whenever a solution exists, the linear system directly yields the weight and intercept functions, and thus a representation of the counterfactual risk in terms of observable marginals. The solution, however, need not be unique.
For exact identification (as opposed to identification up to a constant), we restrict the linear system in Equation (10) to
$\bm{A}^{\textsc{Ext}}\bm{w}^{\textsc{Ext}}=\bm{\ell},$ (11)
where $\bm{w}^{\textsc{Ext}}=\{\omega_{k}(d,y)\}_{d\in\mathcal{D},k\in\mathcal{D},y\in\mathcal{Y}}\in\mathbb{R}^{K^{2}M}$ contains only the weight functions and $\bm{A}^{\textsc{Ext}}$ is adjusted accordingly. The next corollary shows a necessary and sufficient condition for exact identification.
Corollary 6 (Exact-Identification Verification of the Counterfactual Risk)
Suppose that Assumptions 13 hold. Consider a counterfactual loss defined in Definition 1. The corresponding counterfactual risk is exactly identifiable and admits the representation given in Corollary 2 if and only if the linear system given in Equation (11) has a solution.
Lastly, by further restricting the linear system given in Equation (10) to
$\bm{A}^{\textsc{Std}}\bm{w}^{\textsc{Std}}=\bm{\ell},$ (12)
where $\bm{w}^{\textsc{Std}}=\{\omega_{d}(d,y)\}_{d\in\mathcal{D},\,y\in\mathcal{Y}}\in\mathbb{R}^{KM}$ contains only the diagonal weight functions and $\bm{A}^{\textsc{Std}}$ is defined accordingly, we obtain a necessary and sufficient condition under which a counterfactual loss can be equivalently written as a standard loss.
Li and Pearl [2024b] propose a numerical algorithm for verification of exact identification given experimental data. Corollary 5 extends their result to risk differences, and proposes a symbolic verification algorithm that does not require data and operates directly on a loss function itself. Our algorithm also yields an explicit decomposition of counterfactual risk into observable marginals.
3.8 Restrictions on Principal Strata
Corollaries 5 and 6 characterize the relationship between a counterfactual loss and principal strata required for identification. Let $\text{im}(\bm{A}^{\textsc{Add}})$ denote the image of $\bm{A}^{\textsc{Add}}$, i.e., the space of additive counterfactual losses. If $\bm{\ell}\in\text{im}(\bm{A}^{\textsc{Add}})$, then $R(D^{\ast};\ell)$ is identifiable and there exists a solution $\bm{w}$ to Equation (10); otherwise, the counterfactual risk is not identifiable. By standard linear algebra, $\text{im}(\bm{A}^{\textsc{Add}})=\ker\left((\bm{A}^{\textsc{Add}})^{\top}\right)^{\perp}$ [Lax, 2007], i.e., the orthogonal compliment of the kernel of $(\bm{A}^{\textsc{Add}})^{\top}$. Let $\bm{N}^{\textsc{Add}}$ be a matrix whose columns form a basis for $\ker\left((\bm{A}^{\textsc{Add}})^{\top}\right)$. Then $\bm{\ell}\in\text{im}(\bm{A}^{\textsc{Add}})$ if and only if $\left(\bm{N}^{\textsc{Add}}\right)^{\top}\bm{\ell}=\bm{0}$. This condition specifies the strata-level restrictions on the loss needed for identification. Similar reasoning applies to $\bm{A}^{\textsc{Ext}}$ and $\bm{A}^{\textsc{Std}}$. For simplicity, we illustrate this idea in the binary decision and outcome case (see Appendix H for details), though the result can be generalized to the settings with non-binary decision and outcome.
Principal Strata Decision
$D^{\ast}=0$ $D^{\ast}=1$
Never Survivor
$(Y(0),Y(1))=(0,0)$
$\ell(0;0,0)$ $\ell(1;0,0)$
Harmed
$(Y(0),Y(1))=(1,0)$
$\ell(0;1,0)$ $\ell(1;1,0)$
Responsive
$(Y(0),Y(1))=(0,1)$
$\ell(0;0,1)$ $\ell(1;0,1)$
Always Survivor
$(Y(0),Y(1))=(1,1)$
$\ell(0;1,1)$ $\ell(1;1,1)$
Table 2: Counterfactual Loss for Principal Strata in the Binary Case $\ell(D^{\ast};Y(0),Y(1))$.
Table 2 shows all eight possible loss values (two possible decisions across four principal strata), i.e., $\ell(d;y_{0},y_{1})$ for $d,y_{0},y_{1}\in\{0,1\}$, based on the example given in Section 2.1. Restricting to the class of standard losses given in Equation (12), the equality constraints implied by $\left(\bm{N}^{\textsc{Std}}\right)^{\top}\bm{\ell}=\bm{0}$ are given by,
$\ell(0;0,0)=\ell(0;0,1),\quad\ell(0;1,0)=\ell(0;1,1),\quad\ell(1;0,0)=\ell(1;1,0),\quad\ell(1;0,1)=\ell(1;1,1).$ (13)
These represent the familiar restrictions from standard statistical decision theory, where the loss depends only on the chosen decision and its realized outcome. Indeed, when $D^{\ast}=0$, never survivors and responders, for whom we have $Y(0)=0$, share the same loss. Similarly, always survivors and harmed units, for whom we have $Y(0)=1$, have the same loss. When $D^{\ast}=1$, never survivors and harmed units yield the same loss, as do always survivors and responders. Together, this imposes four equality constraints. In Table 2, these correspond to equality between the light orange and purple cells within each column, and separately between the dark orange and purple cells. In this case, only four of the twelve possible weight functions are non-zero:
$\omega_{0}(0,0)=\ell(0;0,y_{1}),\quad\omega_{0}(0,1)=\ell(0;1,y_{1}),\quad\omega_{1}(1,0)=\ell(1;y_{0},0),\quad\omega_{1}(1,1)=\ell(1;y_{0},1),$
for $y_{0},y_{1}\in\{0,1\}$.
We can further relax these conditions by considering the exact identification case in Equation (11), where $\left(\bm{N}^{\textsc{Ext}}\right)^{\top}\bm{\ell}=\bm{0}$ yields,
$\ell(d;0,0)+\ell(d;1,1)=\ell(d;0,1)+\ell(d;1,0),$ (14)
where for each decision $d$, the sum of losses for never and always survivors equals that for the harmed and responsive units. In Table 2, this corresponds to the orange cells summing to the purple cells within each column. As mentioned earlier, the restriction is identical to the gain equality of Li and Pearl [2019]. Given these two equality restrictions, if we let $\omega_{0}(1,1)=a$ and $\omega_{1}(0,1)=b$, we can express the remaining six weights as follows,
$\omega_{0}(0,0)=\ell(0;0,1)-b,\quad\omega_{0}(0,1)=\ell(0;1,1)-b,$ $\quad\omega_{1}(1,0)=\ell(1;1,0)-a,\quad\omega_{1}(1,1)=\ell(1;1,1)-a,$
$\omega_{0}(1,0)=\ell(1;0,1)-\ell(1;1,1)+a,$ $\quad\omega_{1}(0,0)=\ell(0;1,0)-\ell(0;1,1)+b.$
We can further relax the equality constraints to achieve identification only up to a constant by considering Equation (10). In this case, $\left(\bm{N}^{\textsc{Add}}\right)^{\top}\bm{\ell}=\bm{0}$ imposes a single restriction,
$\ell(1;0,0)-\ell(0;0,0)]+[\ell(1;1,1)-\ell(0;1,1)]=[\ell(1;1,0)-\ell(0;1,0)]+[\ell(1;0,1)-\ell(0;0,1)],$ (15)
which represents the sum of loss differences between the two decisions for never-survivors and always-survivors is equal to the corresponding sum for the harmed and responsive units. In Table 2, this restriction corresponds to the sum of orange cells equaling the sum of purple cells across both columns. Solving the system under this restriction yields the weights,
$\omega_{0}(0,0)=\ell(0;0,1)-b-c,$ $\quad\omega_{0}(0,1)=\ell(0;1,1)-b-e,$
$\omega_{1}(1,0)=\ell(1;1,0)-a-d,$ $\quad\omega_{1}(1,1)=\ell(1;1,1)-a-e,$
$\omega_{0}(1,0)=\ell(1;0,1)-\ell(1;1,1)+a-c+e,$ $\quad\omega_{1}(0,0)=\ell(0;1,0)-\ell(0;1,1)+b-d+e,$
$\varpi_{0,1}(0,0)=\ell(1;0,0)-\ell(1;1,0)$- $\ell(1;0,1)+\ell(1;1,1)+c+d-e,$
where $\omega_{0}(1,1)=a,\omega_{1}(0,1)=b,\varpi_{0,1}(0,1)=c,\varpi_{0,1}(1,0)=d,\varpi_{0,1}(1,1)=e$ are free parameters.
Restrictions on loss functions Degrees of Identification
freedom
light (dark) orange = light (dark) purple
within each column (Equation (13))
4 Exact
sum of orange == sum of purple
within each column (Equation (14))
6 Exact*
sum of orange = sum of purple
across two columns (Equation (15))
7 Constant*
No restriction 8 Unidentifiable
Table 3: Summary of identifiability restrictions under which the risk is either exactly identifiable (“Exact” in green), identifiable up to a constant (“Constant” in yellow), or unidentifiable (in red). Strong ignorability is assumed. An asterisk * signifies a necessary and sufficient condition. The second column shows the number of free parameters in the loss specification.
The third column of Table 3 summarizes the identification results discussed in this subsection. The table illustrates how the identification conditions translate into explicit equality constraints on the loss values across principal strata.
4 Concluding Remarks
In this paper, we generalize the classical statistical decision theory for treatment choice by introducing the counterfactual loss that depends on all potential outcomes. While the standard loss evaluates decisions based solely on realized outcomes, the counterfactual loss additionally takes into account what would have happened under alternative decisions. This consideration enables us to assess not only the accuracy of a decision but also its difficulty, that is, whether the chosen decision makes a difference or other decisions would lead to a similar outcome. The counterfactual loss can account for ethical and other considerations that are not possible to express with the standard loss.
While the counterfactual loss is much more expressive than the standard loss, a central challenge is the difficulty of identification. Since only one potential outcome is observed per unit, counterfactual risks are generally unidentifiable. We show that under the standard assumption of strong ignorability, a counterfactual loss leads to an identifiable risk if and only if it is additive in the potential outcomes. Moreover, we establish that such additive counterfactual losses do not have an equivalent standard loss representation so long as decisions are non-binary. Our results should allow practitioners to consider a wider class of loss functions when evaluating decision rules and learning optimal policies. Promising directions for future research include extending the framework to continuous decisions and outcomes and relaxing the strong ignorability assumption in the identification analysis [see Ben-Michael et al., 2025].
Supplementary Appendix
Appendix A Proof of Theorem 1
By the definition of the counterfactual conditional risk (Definition 2), we have,
$R_{\bm{x}}(D^{\ast};\ell^{\textsc{Add}})$
[…]
$\hskip 72.26999pt\times\Pr(D^{\ast}=d,Y(0)=y_{1},\ldots,Y(K-1)=y_{K-1}\mid\bm{X}=\bm{x})$
[…]
$\quad+\sum_{d\in\mathcal{D}}\sum_{\bm{y}\in\mathcal{Y}^{K}}\varpi(\bm{y},\bm{x})\Pr(D^{\ast}=d,Y(0)=y_{0},\ldots,Y(K-1)=y_{K-1}\mid\bm{X}=\bm{x}),$ (A.1)
where the first equality follows from Definition 3. We now consider each of the two summation terms in Equation (A.1). For notational convenience, we write $\bm{y}_{-k}=(y_{0},\ldots,y_{k-1},y_{k+1},\ldots,y_{K-1})$ for any $k\in\mathcal{D}$. We can rewrite the first term as,
$\sum_{d\in\mathcal{D}}\sum_{\bm{y}\in\mathcal{Y}^{K}}\sum_{k\in\mathcal{D}}\omega_{k}(d,y_{k},\bm{x})\Pr(D^{\ast}=d,Y(0)=y_{0},\ldots,Y(K-1)=y_{K-1}\mid\bm{X}=\bm{x})$
[…]
where the last equality follows from Assumptions 2 and 3.
We next consider the second summation in Equation (A.1).
$\sum_{d\in\mathcal{D}}\sum_{\bm{y}\in\mathcal{Y}^{K}}\varpi(\bm{y},\bm{x})\Pr(D^{\ast}=d,Y(0)=y_{0},\ldots,Y(K-1)=y_{K-1}\mid\bm{X}=\bm{x})$
[…]
This completes the proof. ∎
Appendix B Proof of Theorem 2
Theorem 1 immediately establishes the sufficiency by Definition 4, as the counterfactual risk is a function only of the observable marginals (up to an unknown additive constant). We now prove the necessity. That is, we will show that if the counterfactual risk is identifiable up to an unknown additive constant, then the corresponding loss function must take the additive form given in Definition 3.
By Lemma 2, the counterfactual risk is identifiable if and only if the conditional counterfactual risk is identifiable. It therefore suffices to show that if the conditional counterfactual risk is identifiable up to an unknown additive constant depending only on covariates, then the corresponding loss must be additive.
We begin by fixing $\bm{X}=\bm{x}$. Let $\bm{p}(\bm{x}):=(p_{1}(\bm{x}),\ldots,p_{N}(\bm{x}))^{\top}\in\Delta_{N-1}$ denote the (conditional) probability measure over $\{D^{\ast},Y(0),\ldots,Y(K-1)\}$ given $\bm{X}=\bm{x}$ where $N=K\times M^{K}$ and $\Delta_{N-1}$ is the probability simplex defined as,
$\Delta_{N-1}=\left\{\bm{p}\in\mathbb{R}^{N}\ \mathrel{\Big|}\ \sum_{i=1}^{N}p_{i}=1\text{ and }p_{i}\geq 0\text{ for }i=1,\dots,N\right\}.$
Thus, each element of $\bm{p}(\bm{x})$ is indexed by $i=i(d,\bm{y})$ and is equal to $p_{i}=\Pr(D^{\ast}=d,Y(0)=y_{0},\ldots,Y(K-1)=y_{K-1}\mid\bm{X}=\bm{x})$. Then, by Definition 2, we can write the conditional counterfactual risk as the following linear function of $\bm{p}(\bm{x})$,
$R_{\bm{x}}(D^{\ast};\ell)$ $=\bm{\ell}^{\top}\bm{p}(\bm{x}),$
where $\bm{\ell}=(\ell_{1},\ldots,\ell_{N})^{\top}$ with $\ell_{i}=\ell(d,\bm{y},\bm{x})$ for $i=i(d,\bm{y})$.
By Definition 4 and the subsequent discussion in Section 3.4 an identifiable risk can be written entirely as a function of the observable marginals $\Pr(D^{\ast}=d,Y(k)=y_{k}\mid\bm{X}=\bm{x})$. Since we allow the risk to be identifiable up to a constant, it can also depend on the joint distribution of potential outcomes that do not involve the decision $D^{\ast}$, i.e., $\Pr(\bm{Y}(\mathcal{D})=\bm{y}\mid\bm{X}=\bm{x})$.
Let $Q:\Delta_{N-1}\to[0,1]^{L}$ be the observational mapping from Definition 4 with output $Q(\bm{p}(\bm{x}))=\bm{q}(\bm{x})$ the $L$-dimensional stacked vector of observed distribution $\{\Pr(D^{\ast}=d,Y(k)=y_{k}\mid\bm{X}=\bm{x})\}_{\{k\in\mathcal{D},d\in\mathcal{D},y_{k}\in\mathcal{Y}\}}$ and unobserved one $\{\Pr(\bm{Y}(\mathcal{D})=\bm{y}\mid\bm{X}=\bm{x})\}_{\bm{y}\in\mathcal{Y}^{K}}$ with $L=K^{2}\times M+M^{K}$. By Lemma 3(b) we know that $Q$ is surjective to the space of marginals and
$Q(\bm{p}(\bm{x}))=\bm{C}\bm{p}(\bm{x})=\bm{q}(\bm{p}(\bm{x})),$
for every $\bm{p}(\bm{x})\in\Delta_{N-1}$ and corresponding marginal $\bm{q}(\bm{p}(\bm{x}))$, where $\bm{C}$ is a $L\times N$ matrix of zeros and ones. Its kernel satisfies $\text{ker}(\bm{C})=\{\bm{v}\in\mathbb{R}^{N}:\bm{C}\bm{v}=\mathbf{0}_{L},\mathbf{1}_{N}^{\top}\bm{v}=0\}$ where $\mathbf{0}_{L}$ and $\mathbf{1}_{N}$ are an $L$-dimensional vector of zeros and an $N$-dimensional vector of ones, respectively.
Because the conditional risk $R_{\bm{x}}(D^{\ast};\ell)$ is identifiable up to an additive constant, Definition 4 combined with Lemma 1 guarantees the existence of a map $f$ such that
$R_{\bm{x}}(D^{\ast};\ell)=f(\bm{q}(\bm{p}(\bm{x}))),$ (B.1)
for all $\bm{p}(\bm{x})\in\Delta_{N-1}$.
Equation (B.1) differs from our discussion in Section 3.4, in which identifiability is defined with respect to the conditional probability measures over $\{D,D^{\ast},Y(0),\ldots,Y(K-1)\}$ given $\bm{X}=\bm{x}$ that satisfy Assumptions 13. In contrast, Equation (B.1) considers all conditional probability measures over $\{D^{\ast},Y(0),\ldots,Y(K-1)\}$ given $\bm{X}=\bm{x}$. This difference does not affect our conclusion for two reasons. First, the latter distribution fully characterizes the conditional counterfactual risk $R_{\bm{x}}(D^{\ast};\ell)$. Second, Assumptions 13 do not place any restriction on the distribution of $\{D^{\ast},Y(0),\ldots,Y(K-1)\}$ given $\bm{X}=\bm{x}$. Thus, identifiability must hold uniformly over every possible observable collection of the distribution $\{D^{\ast},Y(0),\ldots,Y(K-1)\}$ given $\bm{X}=\bm{x}$, without any restriction. Thus, Equation (B.1) holds for every $\bm{p}(\bm{x})\in\Delta_{N-1}$.
Finally, Lemma 4 implies that there exists a vector $\bm{w}\in\mathbb{R}^{L}$ such that
$R_{\bm{x}}(D;\ell)$ $=\bm{w}^{\top}\bm{q}(\bm{p}(\bm{x}))$
$=\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D}}\sum_{y\in\mathcal{Y}}w_{k}(d,y,\bm{x})\Pr(D^{\ast}(z)=d,Y(k)=y\mid\bm{X}=\bm{x})$
$\quad+\sum_{\bm{y}\in\mathcal{Y}^{K}}\varpi(\bm{y},\bm{x})\Pr(\bm{Y}(\mathcal{D})=\bm{y}_{\mathcal{S}}\mid\bm{X}=\bm{x})$
where $w_{k}(d,y,\bm{x})$ and $\varpi(\bm{y},\bm{x})$ correspond to different entries of $\bm{w}$. We must show that the additive counterfactual loss given in Definition 3 is the only representation that leads to this conditional counterfactual risk. Assume that there exists another loss function representation $\tilde{\bm{\ell}}$ such that
$R_{\bm{x}}(D;\ell)=\bm{\ell}^{\top}\bm{p}(\bm{x})=\tilde{\bm{\ell}}^{\top}\bm{p}(\bm{x}).$
[…]
Appendix C Proof of Corollary 2
The result follows immediately from the proof of Theorem 2 by restricting $\bm{q}(\bm{p}(\bm{x}))$ to the probabilities $\{\Pr(D^{\ast}=d,Y(k)=y_{k}\mid\bm{X}=\bm{x})\}_{d\in\mathcal{D},k\in\mathcal{D},y_{k}\in\mathcal{Y}}$, while omitting the other terms involving multiple potential outcomes. Then, instead of Lemma 3(b), we apply Lemma 3(a).∎
Appendix D Proof of Proposition 1
We first show the existence of such a standard loss. When $\mathcal{D}=\{0,1\}$, Theorem 1 implies,
$R_{\bm{x}}(D^{\ast};\ell^{\textsc{Add}})$
[…]
$\left.\qquad+\omega_{0}(1,y,\bm{x})\Pr(D^{*}=1,Y(0)=y\mid\bm{X}=\bm{x})+\omega_{1}(0,y,\bm{x})\Pr(D^{*}=0,Y(1)=y\mid\bm{X}=\bm{x})\right\}+C(\bm{x})$
where $C(\bm{x})=\sum_{\bm{y}\in\mathcal{Y}^{K}}\varpi(\bm{y},\bm{x})\Pr(\bm{Y}(\mathcal{D})=\bm{y}\mid\bm{X}=\bm{x})$. Observe,
$\Pr(D^{*}=d,Y(1-d)=y\mid\bm{X}=\bm{x})$
$=$ $\Pr(Y(1-d)=y\mid\bm{X}=\bm{x})-\Pr(D^{*}=1-d,Y(1-d)=y\mid\bm{X}=\bm{x}),$
for $d=0,1$. Substituting this into the above equation yields,
$R_{\bm{x}}(D^{\ast};\ell^{\textsc{Add}})$
$=$ $\sum_{y\in\mathcal{Y}}\omega_{0}(0,y,\bm{x})\left\{\Pr(D^{*}=0,Y(0)=y\mid\bm{X}=\bm{x})+\omega_{1}(1,y,\bm{x})\Pr(D^{*}=1,Y(1)=y\mid\bm{X}=\bm{x})\right$.
$\quad+\omega_{0}(1,y,\bm{x})\left(\Pr(Y(0)=y\mid\bm{X}=\bm{x})-\Pr(D^{*}=0,Y(0)=y\mid\bm{X}=\bm{x})\right)$
$\left.\quad+\omega_{1}(0,y,\bm{x})\left(\Pr(Y(1)=y\mid\bm{X}=\bm{x})-\Pr(D^{*}=1,Y(1)=y\mid\bm{X}=\bm{x})\right)\right\}+C(\bm{x})$
$=$ $\sum_{y\in\mathcal{Y}}\left(\omega_{0}(0,y,\bm{x})-\omega_{0}(1,y,\bm{x})\right)\Pr(D^{*}=0,Y(0)=y\mid\bm{X}=\bm{x})$
$\quad+\left(\omega_{1}(1,y,\bm{x})-\omega_{1}(0,y,\bm{x})\right)\Pr(D^{*}=1,Y(1)=y\mid\bm{X}=\bm{x})+C^{\prime}(\bm{x})$
$=$ $\sum_{d\in\mathcal{D}}\sum_{y\in\mathcal{Y}}\ell^{\textsc{Std}}(d,y,\bm{x})\Pr(D^{*}=d,Y(d)=y\mid\bm{X}=\bm{x})+C^{\prime}(\bm{x})$
$=$ $R_{\bm{x}}(D^{\ast};\ell^{\textsc{Std}})+C^{\prime}(\bm{x}),$
where $\ell^{\textsc{Std}}(d,y,\bm{x})=\omega_{d}(d,y,\bm{x})-\omega_{d}(1-d,y,\bm{x})$ and
$C^{\prime}(\bm{x})=$ $\sum_{y\in\mathcal{Y}}\omega_{0}(1,y,\bm{x})\Pr(Y(0)=y\mid\bm{X}=\bm{x})+\omega_{1}(0,y,\bm{x})\Pr(Y(1)=y\mid\bm{X}=\bm{x})$
$\quad+\sum_{\bm{y}\in\mathcal{Y}^{K}}\varpi(\bm{y},\bm{x})\Pr(\bm{Y}(\mathcal{D})=\bm{y}\mid\bm{X}=\bm{x}).$
Thus,
$R_{\bm{x}}(D^{\ast};\ell^{\textsc{Add}})-R_{\bm{x}}(D^{\ast};\ell^{\textsc{Std}})$ $=C^{\prime}(\bm{x}),$
which does not depend on $D^{\ast}$. Taking the expectation on both sides yields the existence of the desired standard loss. This completes the proof of the first part of the proposition.
Next, we prove the uniqueness. Assume there exists another standard loss $\tilde{\ell}^{\textsc{Std}}$ such that $R(D^{\ast},\ell^{\textsc{Add}})-R(D^{\ast},\tilde{\ell}^{\textsc{Std}})=\tilde{c}$ where $\tilde{c}$ is a constant. Then, this assumption implies that the conditional risk difference is given by
$R_{\bm{x}}(D^{\ast};\ell^{\textsc{Std}})-R_{\bm{x}}(D^{\ast};\tilde{\ell}^{\textsc{Std}})$ $=\widetilde{C}(\bm{x}),$ (D.1)
where $\widetilde{C}(\bm{x})$ is a function that satisfies $\tilde{c}=\mathbb{E}[\widetilde{C}(\bm{X})]$ (otherwise, for any $D^{\ast}$ independent of $\bm{X}$, the resulting risk difference would depend on $D^{\ast}$, contradicting our assumption).
By choosing $D^{\ast}\equiv d$, we obtain
$R_{\bm{x}}(d;\ell^{\textsc{Std}})-R_{\bm{x}}(d;\tilde{\ell}^{\textsc{Std}})$ $=\sum_{y\in\mathcal{Y}}\left(\ell^{\textsc{Std}}(d,y,\bm{x})-\tilde{\ell}^{\textsc{Std}}(d,y,\bm{x})\right)\Pr(Y(d)=y\mid\bm{X}=\bm{x})$
$=\mathbb{E}\left[\ell^{\textsc{Std}}(d,Y(d),\bm{X})-\tilde{\ell}^{\textsc{Std}}(d,Y(d),\bm{X})\ \bigr|\ \bm{X}=\bm{x}\right].$
Since this holds for $d=0,1$ and as the risk difference is invariant to the decision, we obtain the following equality,
$\mathbb{E}\left[\ell^{\textsc{Std}}(0,Y(0),\bm{X})-\tilde{\ell}^{\textsc{Std}}(0,Y(0),\bm{X})\ \bigr|\ \bm{X}=\bm{x}\right]=\mathbb{E}\left[\ell^{\textsc{Std}}(1,Y(1),\bm{X})-\tilde{\ell}^{\textsc{Std}}(1,Y(1),\bm{X})\ \bigr|\ \bm{X}=\bm{x}\right].$
for all $\bm{x}$. The left hand side depends on the distribution of $Y(0)$ given $\bm{X}=\bm{x}$, while the right hand side depends on the distribution of $Y(1)$ given $\bm{X}=\bm{x}$. Since there is no restriction on these distributions, the above equality must hold for every choice of these two conditional distributions. This is only possible if the following equality holds,
$\ell^{\textsc{Std}}(d,Y(d),\bm{x})-\tilde{\ell}^{\textsc{Std}}(d,Y(d),\bm{x})=\overline{C}(d,\bm{x}).$
By Equation (D.1), we must have $\overline{C}(d,\bm{x})\equiv\widetilde{C}(\bm{x})$ for $d=0,1$. That is, the difference in loss does not depend on the decision. Hence, we have shown $\ell^{\textsc{Std}}(d,y,\bm{x})=\tilde{\ell}^{\textsc{Std}}(d,y,\bm{x})+\widetilde{C}(\bm{x})$, implying the uniqueness of $\ell^{\textsc{Std}}(d,Y(d),\bm{x})$ up to an additive constant that only depends on covariates.
Finally, the second part of the proposition concerning surjectivity of the mapping follows immediately because any standard loss is a special case of an additive counterfactual loss. For any standard loss $\ell^{\textsc{Std}}(d,y_{d},\bm{x})$ and $\lambda\in\mathbb{R}$, define the weights
$\omega_{d}(d,y,\bm{x})=(1+\lambda)\ell^{\textsc{Std}}(d,y,\bm{x}),\ \quad\ \omega_{d}(1-d,y,\bm{x})=\lambda\ell^{\textsc{Std}}(d,y,\bm{x}),$
for $d=0,1$ and $y\in\mathcal{Y}$. Then, by the first part of this proposition, this corresponds to an additive counterfactual loss with the same treatment recommendations. Since $\lambda$ is arbitrary, there exist infinitely many additive counterfactual losses that map to the same standard loss. ∎
Appendix E Proof of Proposition 2
Let $\ell^{\textsc{Add}}(d;\bm{y},\bm{x})$ denote an additive counterfactual loss defined in Definition 3. Without loss of generality, we set $\varpi\equiv 0$ because this term does not depend on decision.
We first show if no weights are off-diagonal additively separable, then there exists no equivalent standard loss. We prove this claim through proof by contradiction. Suppose that there exists a standard loss $\ell^{\textsc{Std}}(d;y_{d},\bm{x})$ such that the risk difference $R(D^{\ast};\ell^{\textsc{Add}})-R(D^{\ast};\ell^{\textsc{Std}})$ does not depend on $D^{\ast}$. By applying Lemma 6 to the risk difference, this implies that for fixed $\bm{X}=\bm{x}$, there exists $C(\bm{x})$ such that the following conditional risk difference does not depend on constant decisions,
$\Delta_{\bm{x}}(\pi_{j})=R_{\bm{x}}(\pi_{j};\ell^{\textsc{Add}})-R_{\bm{x}}(\pi_{j};\ell^{\textsc{Std}})=C(\bm{x}),$ (E.1)
where $\pi_{j}=\pi_{j}(\bm{x})=j$ for $j\in\mathcal{D}$ is a constant decision rule. Let
$\Delta_{\bm{x}}(\pi_{j})$ $=\sum_{k\in\mathcal{D}}\sum_{y_{k}\in\mathcal{Y}}\omega_{k}(j,y_{k},\bm{x})\Pr(Y(k)=y_{k}\mid\bm{X}=\bm{x})-\sum_{y_{j}\in\mathcal{Y}}\widetilde{\omega}(j,y_{j},\bm{x})\Pr(Y(j)=y_{j}\mid\bm{X}=\bm{x})$
$=\sum_{k\in\mathcal{D}}\mathbb{E}[\omega_{k}(j,Y(k),\bm{X})\mid\bm{X}=\bm{x}]-\mathbb{E}[\widetilde{\omega}(j,Y(j),\bm{X})\mid\bm{X}=\bm{x}],$
where $\widetilde{\omega}(j,y_{j},\bm{x})=\ell^{\textsc{Std}}(j,y_{j},\bm{x})$ is the weight function for the standard loss. Note that $\Delta_{\bm{x}}(\pi_{j})=C(\bm{x})$ by Equation (E.1) for every $j\in\mathcal{D}$. Let $j,j^{\prime}\in\mathcal{D}$ with $j\neq j^{\prime}$. Then, $\Delta_{\bm{x}}(\pi_{j})-\Delta_{\bm{x}}(\pi_{j^{\prime}})=C(\bm{x})-C(\bm{x})=0$ or equivalently,
$\sum_{k\in\mathcal{D}}\mathbb{E}[\omega_{k}(j,Y(k),\bm{X})-\omega_{k}(j^{\prime},Y(k),\bm{X})\mid\bm{X}=\bm{x}]=\mathbb{E}[\widetilde{\omega}(j,Y(j),\bm{X})-\widetilde{\omega}(j^{\prime},Y(j^{\prime}),\bm{X})\mid\bm{X}=\bm{x}].$ (E.2)
Since the potential outcomes $\{Y(0),\ldots,Y(K-1)\}$ can have an arbitrary conditional distribution given $\bm{X}=\bm{x}$, Equation (E.2) must hold for every choice of this distribution. Since the right-hand side of this equation depends on the conditional distribution of $\{Y(j),Y(j^{\prime})\}$ given $\bm{X}=\bm{x}$, so must its left-hand side. In particular, for every $k\notin\{j,j^{\prime}\}$, $\omega_{k}(j,y_{k},\bm{x})-\omega_{k}(j^{\prime},y_{k},\bm{x})$ cannot depend on $y_{k}$, for all $y_{k}\in\mathcal{Y}$ and $\bm{x}\in\mathcal{X}$. Note that such $k$ exists so long as $K\geq 3$. By Lemma 5, this statement is equivalent to another statement that $\omega_{k}(d,y_{k},\bm{x})$ is off-diagonal additively separable. However, this contradicts our assumption, which proves the claim.
We next show that if no equivalent standard loss exists, then at least one weight function is not additively separable. Again, we show this through proof by contradiction. Suppose all weights of $\ell^{\textsc{Add}}$ are additively-separable, i.e., for any $k\in\mathcal{D}$, there exist $f_{k}$ and $g_{k}$ such that
$\omega_{k}(d,y,\bm{x})=f_{k}(d,\bm{x})+g_{k}(y,\bm{x}),$
for all $d\in\mathcal{D}\setminus\{k\},y\in\mathcal{Y}$. Under Assumptions 13, we can apply Theorem 1 and write the conditional counterfactual risk (recall $\varpi\equiv 0$) as:
$R_{\bm{x}}(D^{\ast};\ell^{\textsc{Add}})$
$=$ $\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D}}\sum_{y\in\mathcal{Y}}\omega_{k}(d,y,\bm{x})\Pr(D^{*}=d,Y(k)=y\mid\bm{X}=\bm{x})$
$=$ $\sum_{d\in\mathcal{D}}\sum_{y\in\mathcal{Y}}\omega_{d}(d,y,\bm{x})\Pr(D^{*}=d,Y(d)=y\mid\bm{X}=\bm{x})$
$+\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D},k\neq d}\sum_{y\in\mathcal{Y}}\omega_{k}(d,y,\bm{x})\Pr(D^{*}=d,Y(k)=y\mid\bm{X}=\bm{x})$
$=$ $\sum_{d\in\mathcal{D}}\sum_{y\in\mathcal{Y}}\omega_{d}(d,y,\bm{x})\Pr(D^{*}=d,Y(d)=y\mid\bm{X}=\bm{x})$
$+\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D},k\neq d}\sum_{y\in\mathcal{Y}}(f_{k}(d,\bm{x})+g_{k}(y,\bm{x}))\Pr(D^{*}=d,Y(k)=y\mid\bm{X}=\bm{x})$
$=$ $\sum_{d\in\mathcal{D}}\sum_{y\in\mathcal{Y}}\omega_{d}(d,y,\bm{x})\Pr(D^{*}=d,Y(d)=y\mid\bm{X}=\bm{x})$
$+\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D},k\neq d}\sum_{y\in\mathcal{Y}}f_{k}(d,\bm{x})\Pr(D^{*}=d,Y(k)=y\mid\bm{X}=\bm{x})$
$+\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D},k\neq d}\sum_{y\in\mathcal{Y}}g_{k}(y,\bm{x})\Pr(D^{*}=d,Y(k)=y\mid\bm{X}=\bm{x})$
$=$ $\sum_{d\in\mathcal{D}}\sum_{y\in\mathcal{Y}}\omega_{d}(d,y,\bm{x})\Pr(D^{*}=d,Y(d)=y\mid\bm{X}=\bm{x})$
$+\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D},k\neq d}f_{k}(d,\bm{x})\sum_{y\in\mathcal{Y}}\Pr(D^{*}=d,Y(k)=y\mid\bm{X}=\bm{x})$
$+\sum_{k\in\mathcal{D}}\sum_{y\in\mathcal{Y}}g_{k}(y,\bm{x})\sum_{d\in\mathcal{D},d\neq k}\Pr(D^{*}=d,Y(k)=y\mid\bm{X}=\bm{x})$
$=$ $\sum_{d\in\mathcal{D}}\sum_{y\in\mathcal{Y}}\omega_{d}(d,y,\bm{x})\Pr(D^{*}=d,Y(d)=y\mid\bm{X}=\bm{x})+\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D},k\neq d}f_{k}(d,\bm{x})\Pr(D^{*}=d\mid\bm{X}=\bm{x})$
$+\sum_{k\in\mathcal{D}}\sum_{y\in\mathcal{Y}}g_{k}(y,\bm{x})(\Pr(Y(k)=y\mid\bm{X}=\bm{x})-\Pr(D^{*}=k,Y(k)=y\mid\bm{X}=\bm{x}))$
$=$ $\sum_{d\in\mathcal{D}}\sum_{y\in\mathcal{Y}}(\omega_{d}(d,y,\bm{x})-g_{d}(y,\bm{x}))\Pr(D^{*}=d,Y(d)=y\mid\bm{X}=\bm{x})$
$+\sum_{d\in\mathcal{D}}\left(\sum_{k\in\mathcal{D},k\neq d}f_{k}(d,\bm{x})\right)\Pr(D^{*}=d\mid\bm{X}=\bm{x})+\sum_{k\in\mathcal{D}}g_{k}(y,\bm{x})\Pr(Y(k)=y\mid\bm{X}=\bm{x})$
Now, define $\ell^{\textsc{Std}}(d;y_{d},\bm{x})=(\omega_{d}(d,y,\bm{x})-g_{d}(y,\bm{x}))+\sum_{k\in\mathcal{D},k\neq d}f_{k}(d,\bm{x})$ as in Proposition 2. Then, by the above calculations,
$R_{\bm{x}}(D^{\ast};\ell^{\textsc{Add}})-R_{\bm{x}}(D^{\ast};\ell^{\textsc{Std}})=\sum_{k\in\mathcal{D}}g_{k}(y,\bm{x})\Pr(Y(k)=y\mid\bm{X}=\bm{x}),$
which does not depend on $D^{\ast}$. This contradicts the existence of no equivalent standard loss.
Finally, we show that $\ell^{\textsc{Std}}$ is unique up to a constant that only depends on covariates. Assume there exists another standard loss $\tilde{\ell}^{\textsc{Std}}$ such that $R(D^{\ast},\ell^{\textsc{Add}})-R(D^{\ast},\tilde{\ell}^{\textsc{Std}})=\tilde{c}$ where $\tilde{c}$ is a constant. As in the proof of Proposition 1, this implies that the conditional risk difference is given by
$R_{\bm{x}}(D^{\ast};\ell^{\textsc{Std}})-R_{\bm{x}}(D^{\ast};\tilde{\ell}^{\textsc{Std}})$ $=\widetilde{C}(\bm{x}),$ (E.3)
where $\widetilde{C}(\bm{x})$ is a function that satisfies $\tilde{c}=\mathbb{E}[\widetilde{C}(\bm{X})]$. Write the loss difference as $\Delta(d,y,\bm{x}):=\ell^{\textsc{Std}}(d,y,\bm{x})-\tilde{\ell}^{\textsc{Std}}(d,y,\bm{x})$, then by definition,
$R_{\bm{x}}(D^{\ast};\ell^{\textsc{Std}})-R_{\bm{x}}(D^{\ast};\tilde{\ell}^{\textsc{Std}})$ $=\sum_{d\in\mathcal{D}}\sum_{y\in\mathcal{Y}}\Delta(d,y,\bm{x})\Pr(D^{*}=d,Y(d)=y\mid\bm{X}=\bm{x})$
By choosing $D^{\ast}=d$, we obtain
$R_{\bm{x}}(d;\ell^{\textsc{Std}})-R_{\bm{x}}(d;\tilde{\ell}^{\textsc{Std}})$ $=\sum_{y\in\mathcal{Y}}\Delta(d,y,\bm{x})\Pr(Y(d)=y\mid\bm{X}=\bm{x}),$
and thus
$\mathbb{E}\left[\Delta(d,Y(d),\bm{x})\mid\bm{X}=\bm{x}\right]=\widetilde{C}(\bm{x})\qquad\forall d\in\mathcal{D}$
by Equation (E.3). The left hand side depends on the conditional distribution of $Y(d)$ given $\bm{X}=\bm{x}$ while the right hand side does not. Since there is no restriction on this conditional distribution, the equality must hold for every choice of this distribution. Furthermore, the equality holds for every $d\in\mathcal{D}$, but the right hand site is independent of $d$. These two observations can only be true if $\Delta(d,y,\bm{x})$ does not depend on $d$ and $y$. Thus, we have,
$\ell^{\textsc{Std}}(d,y,\bm{x})-\tilde{\ell}^{\textsc{Std}}(d,y,\bm{x})=\widetilde{C}(\bm{x}),$
which proves the uniqueness of equivalent standard loss. ∎.
Appendix F Proof of Corollary 3
By Theorem 2, the conditional counterfactual risk can be written as,
$R_{\bm{x}}(D)$ $=\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D}}\sum_{y=0}^{1}\omega_{k}(d,y,\bm{x})\Pr(D=d,Y(k)=y\mid\bm{X}=\bm{x})+C(\bm{x}),$ (F.1)
where $\sum_{\bm{y}\in\{0,1\}^{K}}\varpi(\bm{y},\bm{x})\Pr(\bm{Y}(\mathcal{S})=\bm{y}_{\mathcal{S}}\mid\bm{X}=\bm{x})$. First note,
$\Pr(D^{*}=d,Y(k)=0\mid\bm{X}=\bm{x})=\Pr(D^{*}=d\mid\bm{X}=\bm{x})-\Pr(D^{*}=d,Y(k)=1\mid\bm{X}=\bm{x}).$
Substituting this expression into the first term of Equation (F.1) yields,
$\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D}}\sum_{y=0}^{1}\omega_{k}(d,y,\bm{x})\Pr(D^{*}=d,Y(k)=y\mid\bm{X}=\bm{x})$
[…]
$\quad+\omega_{k}(d,1,\bm{x})\Pr(D^{*}=d,Y(k)=1\mid\bm{X}=\bm{x})$
[…]
$+\sum_{d=1}^{K}\zeta_{d}(d,\bm{x})\Pr(D^{*}=d,Y(d)=1\mid\bm{X}=\bm{x}),$
where $\zeta_{k}(d,\bm{x})=\omega_{k}(d,1,\bm{x})-\omega_{k}(d,0,\bm{x})$, $\xi(d,\bm{x})=\sum_{k\in\mathcal{D}}\omega_{k}(d,0,\bm{x})$. ∎
Appendix G Lemmas and their Proofs
G.1 Lemma 1
Lemma 1
Let $\mathcal{P}$ denote the set of all joint probability distributions under consideration, and let $\mathcal{Q}$ be the set of all observable distributions. Let $Q:\mathcal{P}\to\mathcal{Q}$ denote a surjective mapping that associates each joint distribution with an observable distribution. Let $\theta(P)$ denote the (causal) estimand of interest under distribution $P\in\mathcal{P}$. For a fixed observable distribution $Q_{0}\in\mathcal{Q}$ define $\mathcal{S}_{0}=\{P\in\mathcal{P}:Q(P)=Q_{0}\}.$ Then, the following statements are equivalent:
  1. (a)
    For every $Q_{0}\in\mathcal{Q}$ there exists a constant $\theta_{0}$ such that $\theta(P)=\theta_{0}$ for all $P\in\mathcal{S}_{0}$.
  2. (b)
    There exists a function $f$ such that $\theta(P)=f(Q(P))$ for every $P\in\mathcal{P}$.
Proof.
We first prove that (b) implies (a). Fix any $Q_{0}\in\mathcal{Q}$ and let $P\in\mathcal{S}_{0}$ such that $Q(P)=Q_{0}$. Note that $\mathcal{S}_{0}$ is non-empty by $Q$ surjective. Then, by (b), there exists $f$ such that
$\theta(P)=f(Q(P))=f(Q_{0})=\theta_{0},$
for every $P\in\mathcal{S}_{0}$. Because $Q_{0}$ was arbitrary, the claim follows.
We next prove that (a) implies (b). Let $Q_{0}\in\mathcal{Q}$. Surjectivity of $Q$ ensures $\mathcal{S}_{0}$ is non-empty. By (a), there is a constant $\theta_{0}$ such that $\theta(P)=\theta_{0}$ for all $P\in\mathcal{S}_{0}$. Define $f:\mathcal{Q}\to\mathbb{R}$ by setting $f(Q_{0}):=\theta_{0}$ for every such $Q_{0}\in\mathcal{Q}$. Now, let $P\in\mathcal{P}$ be an arbitrary probability measure and let $Q_{0}=Q(P)$. Then, there exists $\theta_{0}=f(Q_{0})$ such that $\theta(P)=\theta_{0}=f(Q_{0})=f(Q(P))$. This proves the claim. ∎
G.2 Lemma 2
Lemma 2
Given a counterfactual loss $\ell$ defined in Definition 1, the counterfactual risk $R(D^{\ast},\ell)$ defined in Definition 2 is identifiable with respect to $\mathbb{P}(D^{\ast},D,Y,\bm{X})$ if and only if the conditional counterfactual risk $R_{\bm{X}}(D^{\ast},\ell)$ (also defined in Definition 2) is identifiable with respect to $\mathbb{P}(D^{\ast},D,Y\mid\bm{X})$.
Proof.
Let $Q:\mathcal{P}\to\mathcal{Q}$ be the mapping from $P\in\mathcal{P}$ to the distribution $Q(P)=P(D^{\ast},D,Y,\bm{X})$. Further, let $Q_{\bm{X}}:\mathcal{P}\to\mathcal{Q}_{\bm{X}}$ be the mapping from $P\in\mathcal{P}$ to the conditional distribution $Q_{\bm{X}}(P)=P(D^{\ast},D,Y\mid\bm{X})$. We use $P_{\bm{X}}$ to denote the marginal probability distribution of $\bm{X}$. Note that for any $P\in\mathcal{P}$ we can write
$P(D^{\ast},D,Y(\mathcal{D}),\bm{X})$ $=P(D^{\ast},D,Y,Y(\mathcal{D}),\bm{X})$
$=P(Y(\mathcal{D})\mid D^{\ast},D,Y,\bm{X})Q(P)$
$=P(Y(\mathcal{D})\mid D^{\ast},D,Y,\bm{X})Q_{\bm{X}}(P)P_{\bm{X}}(\bm{X}),$
where the first equality follows from Assumption 2. To ease notation, we write $R=R(D^{\ast},\ell)$ and $R_{\bm{x}}=R_{\bm{x}}(D^{\ast},\ell)$.
We first assume that $R$ is identifiable. We use proof by contradiction. Suppose that $R_{\bm{x}}$ is not identifiable. Then, by Definition 4, there exist a conditional observable distribution $Q_{0}(\bm{X})\subseteq\mathcal{Q}_{\bm{X}}$ and distributions $P_{1},P_{2}\in\mathcal{P}$ such that $Q_{\bm{X}}(P_{1})=Q_{\bm{X}}(P_{2})=Q_{0}(\bm{X})$ but $\Delta(\bm{x})=R_{\bm{x}}(P_{1})-R_{\bm{x}}(P_{2})\neq 0$ on some measurable set $\widetilde{\mathcal{X}}\subseteq\mathcal{X}$. Let $\widetilde{P}_{\bm{X}}$ be a probability measure such that
$\int_{\bm{x}\in\mathcal{X}}\Delta(\bm{x})d\widetilde{P}_{\bm{X}}(\bm{x})\neq 0.$
We can always construct such a probability measure $P_{\bm{X}}(\bm{x})$, for example, by concentrating the probability mass on a set $\widetilde{\mathcal{X}}$ where $\Delta(\bm{x})>0$ for all $\bm{x}\in\widetilde{X}$. Next, define another probability measure $\widetilde{P}_{i}$ for $i=1,2$:
$\widetilde{P}_{i}(D^{\ast},D,Y(\mathcal{D}),\bm{X})=P_{i}(Y(\mathcal{D})\mid D^{\ast},D,Y,\bm{X})Q_{0}(\bm{X})\widetilde{P}_{\bm{X}}(\bm{X}).$
where each $\widetilde{P}_{i}\in\mathcal{P}$ satisfies Assumptions 13, as these assumptions do not restrict the marginal of $\bm{X}$. Since the marginal distribution of $\bm{X}$ does not affect the conditional risk, we have $R_{\bm{x}}(\widetilde{P}_{i})=R_{\bm{x}}(P_{i})$ for $i=1,2$ and
$R_{\bm{x}}(\widetilde{P}_{1})-R_{\bm{x}}(\widetilde{P}_{2})=R_{\bm{x}}(P_{1})-R_{\bm{x}}(P_{2})=\Delta(\bm{x}).$
Further, by construction, we have $Q(\widetilde{P}_{1})=Q_{0}(\bm{X})\widetilde{P}_{\bm{X}}(\bm{X})=Q(\widetilde{P}_{2})\in\mathcal{Q}$. Since $R$ is identifiable, Definition 4 implies that $R(\widetilde{P}_{1})=R(\widetilde{P}_{2})$. However,
$R(\widetilde{P}_{1})-R(\widetilde{P}_{2})=$ $\int_{\bm{x}\in\mathcal{X}}R_{\bm{x}}(\widetilde{P}_{1})-R_{\bm{x}}(\widetilde{P}_{2})d\widetilde{P}_{\bm{X}}(\bm{x})$
$=$ $\int_{\bm{x}\in\mathcal{X}}R_{\bm{x}}(P_{1})-R_{\bm{x}}(P_{2})d\widetilde{P}_{\bm{X}}(\bm{x})$
$=$ $\int_{\bm{x}\in\mathcal{X}}\Delta(\bm{x})d\widetilde{P}_{\bm{X}}(\bm{x})$
$\neq$ $0,$
which leads to a contradiction. Thus, $R_{\bm{x}}$ must be identifiable.
Assume next that $R_{\bm{X}}$ is identifiable. Then, by Lemma 1, there exists a function $g$ such that $R_{\bm{X}}(P)=g(Q_{\bm{X}}(P))$. Define a function $f:\mathcal{Q}\to\mathbb{R}$ such that
$f(Q(P))=\int_{x\in\mathcal{X}}g(Q_{\bm{X}}(P))dP_{\bm{X}}(\bm{x}).$
Note that we can define $f$ in such a way as both $Q_{\bm{X}}(P)$ and $P_{\bm{X}}$ are obtainable from $Q$. Hence $R(P)=f(Q(P))$ and so $R$ is identifiable by Lemma 1. ∎
G.3 Lemma 3
Lemma 3
Let $D,Y_{0},\ldots,Y_{K-1}$ be discrete random variables such that $D\in\mathcal{D}=\{0,1,\ldots,K-1\}$ and $Y_{0},\ldots,Y_{K-1}\in\mathcal{Y}=\{0,1,\ldots,M-1\}$. Let $\Delta_{N-1}=\{\bm{p}\in\mathbb{R}^{N}\mid\sum_{i=1}^{N}p_{i}=1\text{ and }p_{i}\geq 0\text{ for }i=1,\ldots,N\}$ denote the probability simplex in $\mathbb{R}^{N}$ where $N=K\times M^{K}$. Then, for every probability distribution $P(D,Y_{0},\ldots,Y_{K-1})$, there exists $\bm{p}\in\Delta_{N-1}$ where $p_{i}=\Pr(D=d,Y_{0}=y_{0},\ldots,Y_{K-1}=y_{K-1})$ for each $i=i(d,\bm{y})\in\{1,\ldots,N\}$. Consider the following two cases that yield the same conclusion.
  1. (a)
    Let $\bm{q}=\bm{q}(\bm{p})\in[0,1]^{L}$ be a vector of probabilities $\{\Pr(D=d,Y_{k}=y_{k})\}_{d\in\mathcal{D},k\in\mathcal{D},y_{k}\in\mathcal{Y}}$ where $q_{j}=\Pr(D=d,Y_{k}=y_{k})$ for $j=j(d,k,y_{k})\in\{1,\ldots,L\}$ and $L=K^{2}\times M$.
  2. (b)
    Let $\bm{q}=\bm{q}(\bm{p})\in[0,1]^{L}$ be the extended vector of probabilities that also includes the joint distribution of $Y_{0},\ldots,Y_{K-1}$, i.e., probabilities $\{\Pr(D=d,Y_{k}=y_{k})\}_{d\in\mathcal{D},\,k\in\mathcal{D},\,y_{k}\in\mathcal{Y}}$ and $\{\Pr(Y_{0}=y_{0},\ldots,Y_{K-1}=y_{K-1})\}_{\bm{y}\in\mathcal{Y}^{K}}$, with $L=K^{2}\times M+M^{K}$.
Then, in each case, there exists an $L\times N$ matrix $\bm{C}$ of zeros and ones such that $\bm{q}=\bm{C}\bm{p}$ for all $\bm{p}\in\Delta_{N-1}$ and
$\text{ker}(\bm{C})=\left\{\bm{v}\in\mathbb{R}^{N}:\sum_{\bm{y}_{-k}\in\mathcal{Y}^{K-1}}v_{d,\bm{y}}=0\ \text{for all }d\in\mathcal{D},k\in\mathcal{D},y_{k}\in\mathcal{Y},\ \mathbf{1}_{N}^{\top}\bm{v}=0\ \text{and}\ \bm{C}\bm{v}=\mathbf{0}_{L}\right\},$
where $\bm{y}_{-k}=(y_{0},\ldots,y_{k-1},y_{k+1},\ldots,y_{K-1})$ and $\mathbf{1}_{N}$ represents an $N$-dimensional vector of ones. Furthermore, $\bm{C}$ is a surjective map from the space of joint distribution $P(D,Y_{0},\ldots,Y_{K-1})$ onto the space of the respective marginal distribution.
We now prove each case in turn.
Case (a).
For every joint probability distribution $P(D,Y_{0},\ldots,Y_{K-1})$, we can express the marginal distribution of $P(D,Y_{k})$ as,
$\Pr(D=d,Y_{k}=y_{k})=\sum_{d^{\prime}\in\mathcal{D}}\sum_{\bm{y}^{\prime}\in\mathcal{Y}^{K}}\mathds{1}{\left\{d^{\prime}=d,y^{\prime}_{k}=y_{k}\right\}}\Pr(D=d^{\prime},Y_{0}=y^{\prime}_{0},\ldots,Y_{K-1}=y^{\prime}_{K-1}).$ (G.1)
Define a matrix $\bm{C}\in\{0,1\}^{L\times N}$ whose $(j,i)$ entry is given by,
$C_{\,j(d,k,y_{k}),\,i(d^{\prime},\bm{y}^{\prime})}=\mathds{1}{\left\{d^{\prime}=d,y^{\prime}_{k}=y_{k}\right\}}.$
Then, Equation (G.1) can be written as,
$q_{j(d,k,y_{k})}=\sum_{d^{\prime}\in\mathcal{D}}\sum_{\bm{y}^{\prime}\in\mathcal{Y}^{K}}C_{\,j(d,k,y_{k}),\,i(d^{\prime},\bm{y}^{\prime})}p_{i(d^{\prime},\bm{y}^{\prime})},$
implying $\bm{q}=\bm{C}\bm{p}$. Since every marginal distribution has a corresponding joint probability distribution, $\bm{C}$ is a surjective map from the space of joint probability distributions onto the space of marginal distributions.
We proceed to examine the kernel of the matrix $\bm{C}$. Let $\bm{v}=(v_{d,\bm{y}})\in\mathbb{R}^{N}$. The definition of $\bm{C}$ implies that for all $j=j(d,k,y_{k})\in\{1,\ldots,L\}$, we have,
$(\bm{C}\bm{v})_{j(d,k,y_{k})}=\sum_{d^{\prime}\in\mathcal{D}}\sum_{\bm{y}^{\prime}\in\mathcal{Y}^{K}}\mathds{1}{\left\{d^{\prime}=d,y^{\prime}_{k}=y_{k}\right\}}\bm{v}_{(d^{\prime},\bm{y})}=\sum_{\bm{y}_{-k}\in\mathcal{Y}^{K-1}}v_{d,\bm{y}_{-k}}.$
Therefore, $\bm{C}\bm{v}=\mathbf{0}_{L}$ if and only if $\sum_{\bm{y}_{-k}\in\mathcal{Y}^{K-1}}v_{d,\bm{y}}=0$ for all $d\in\mathcal{D},\;k\in\mathcal{D},\;y_{k}\in\mathcal{Y}$. This proves the first part of the kernel property. The second part follows immediately by observing,
$\mathbf{1}_{N}^{\top}\bm{v}=\sum_{d\in\mathcal{D}}\sum_{\bm{y}\in\mathcal{Y}^{K}}v_{d,\bm{y}}=\sum_{d\in\mathcal{D}}\sum_{k\in\mathcal{D}}\sum_{y_{k}\in\mathcal{Y}}\sum_{\bm{y}_{-k}\in\mathcal{Y}^{K-1}}v_{d,\bm{y}}=0.$
Case (b).
The existence and surjectivity of $\bm{C}$ follow from the same arguments as in the above proof of Case (a). Thus, we only need to verify the kernel property of $\bm{C}$. We start by partitioning
$\bm{C}=\begin{bmatrix}\bm{C}_{1}\\ \bm{C}_{2}\\ \end{bmatrix}$,
where $\bm{C}_{1}$ is an $K^{2}M\times N$ matrix of zeros and ones, containing all the rows corresponding to the marginal distributions $\{\Pr(D=d,Y_{k}=y_{k})\}_{d\in\mathcal{D},k\in\mathcal{D},y_{k}\in\mathcal{Y}}$, which are included in $\bm{C}$ by assumption. Let $\bm{v}\in\text{Ker}(\bm{C})$, then,
$\bm{C}\bm{v}=\begin{bmatrix}\bm{C}_{1}\bm{v}\\ \bm{C}_{2}\bm{v}\\ \end{bmatrix}=\begin{bmatrix}\bm{0}\\ \bm{0}\\ \end{bmatrix}$,
implying $\bm{C}_{1}\bm{v}=0$. Thus, $\bm{v}\in\text{ker}(\bm{C}_{1})$ and consequently, $\text{ker}(\bm{C})\subseteq\text{ker}(\bm{C}_{1})$. Since $\bm{C}_{1}$ corresponds to the matrix described in Case (a), this completes the proof. ∎
G.4 Lemma 4
Lemma 4
Let $\Delta_{N-1}=\{\bm{p}\in\mathbb{R}^{N}\mid\sum_{i=1}^{N}p_{i}=1\text{ and }p_{i}\geq 0\text{ for }i=1,\ldots,N\}$ denote the probability simplex in $\mathbb{R}^{N}$. Let $\bm{C}\in\mathbb{R}^{L\times N}$ be a matrix such that $\text{ker}(\bm{C})=\left\{\bm{v}\in\mathbb{R}^{N}:\mathbf{1}_{N}^{\top}\bm{v}=0\ \text{and}\ \bm{C}\bm{v}=\mathbf{0}_{L}\right\}$. Set $\bm{C}(\Delta_{n-1})=\{\bm{q}\in\mathbb{R}^{L}\mid\exists\bm{p}\in\Delta_{n-1}:\bm{q}=\bm{C}\bm{p}\}$ to the image of $\bm{C}$ on $\Delta_{n-1}$. Let $f_{\bm{\ell}}:\bm{C}(\Delta_{n-1})\to\mathbb{R}$ be a function for which there exists $\bm{\ell}\in\mathbb{R}^{N}$ such that for all $\bm{q}(\bm{p})\in\bm{C}(\Delta_{n-1})$, we have $f_{\bm{\ell}}(\bm{q})=f_{\bm{\ell}}(\bm{C}\bm{p})=\bm{\ell}^{\top}\bm{p}$.
[…]
Proof.
Observe that if $\bm{\ell}\in\text{im}(\bm{C}^{\top})$, then there exists $\bm{w}\in\mathbb{R}^{L}$ such that $\bm{\ell}=\bm{C}^{\top}\bm{w}$. Then, for any $\bm{q}=\bm{C}\bm{p}$ we have
$f_{\bm{\ell}}(\bm{q})=\bm{\ell}^{\top}\bm{p}=\bigl(\bm{C}^{\top}\bm{w}\bigr)^{\!\top}\bm{p}=\bm{w}^{\top}\bm{C}\bm{p}=\bm{w}^{\top}\bm{q}.$
Hence, $f_{\bm{\ell}}(\bm{q})=\bm{w}^{\top}\bm{q}$.
Assume now that $\bm{\ell}\notin\text{im}(\bm{C}^{\top})$. We will prove that this yields a contradiction. By standard linear algebra [Lax, 2007, Theorem 5, Chapter 3], $\text{im}(\bm{C}^{\top})=\text{ker}(\bm{C})^{\perp}$. This implies that $\exists\bm{v}\in\text{ker}(\bm{C})\setminus\{\bm{0}\}:\bm{\ell}^{\top}\bm{v}\neq 0$. Let $\bm{p}_{0}\in\mathring{\Delta}_{n-1}$ be in the interior of the simplex with strictly positive entries and set $\bm{q}_{0}=\bm{C}\bm{p}_{0}$. Define
$\bm{p}=\bm{p}_{0}+\alpha\bm{v}\quad\text{with}\ 0<\alpha<\frac{\min_{i\in[N]}{p_{0}}_{i}}{\max_{i\in[N]}\left|v_{i}\right|}.$
Note that such an $\alpha$ exists. Indeed, by assumption, $0<{p_{0}}_{i}$ for all entries and $\bm{v}$ has at least one non-zero entry as $\bm{v}\neq\bm{0}$. Observe,
$\min_{i\in[n]}p_{i}\geq\min_{i\in[n]}{p_{0}}_{i}-\alpha\max_{i\in[n]}\left|v_{i}\right|>0,$
and,
$\sum_{i=1}^{n}p_{i}=\sum_{i=1}^{n}{p_{0}}_{i}+\alpha\sum_{i=1}^{n}v_{i}=\sum_{i=1}^{n}{p_{0}}_{i}=1,$
where the second equality follows from $\sum_{i=1}^{n}v_{i}=0$ by $\bm{v}\in\text{ker}(\bm{C})$. Thus, $\bm{p}\in\Delta_{N-1}$.
Because $\bm{v}\in\text{ker}(\bm{C})$,
$\bm{C}\bm{p}=\bm{C}\bigl(\bm{p}_{0}+\alpha\bm{v}\bigr)=\bm{C}\bm{p}_{0}=\bm{q}_{0}.$
This implies that $f_{\bm{\ell}}(\bm{q}_{0})=f_{\bm{\ell}}(\bm{C}\bm{p})=f_{\bm{\ell}}(\bm{C}\bm{p}_{0})$ and,
$0=f_{\bm{\ell}}(\bm{C}\bm{p})-f_{\bm{\ell}}(\bm{C}\bm{p}_{0})=\bm{\ell}^{\top}\bm{p}-\bm{\ell}^{\top}\bm{p}_{0}=\bm{\ell}^{\top}\left(\bm{p}-\bm{p}_{0}\right)=\alpha\bm{\ell}^{\top}\bm{v}\neq 0,$
as $\bm{\ell}^{\top}\bm{v}\neq 0$ by assumption. This yields a contradiction. ∎
G.5 Lemma 5
Lemma 5
Let $\left|\mathcal{D}\right|\geq 3$ and fix $k\in\mathcal{D},\bm{x}\in\mathcal{X}$. Consider the weight function $\omega_{k}:\mathcal{D}\times\mathcal{Y}\times\mathcal{X}\to\mathbb{R}$. Then, the following statements are equivalent:
  1. (a)
    For all $d,d^{\prime}\in\mathcal{D}\setminus\{k\}$ the difference $\omega_{k}(d,y,\bm{x})-\omega_{k}(d^{\prime},y,\bm{x})$ does not depend on $y\in\mathcal{Y}$
  2. (b)
    There exist functions $f_{k}$ and $g_{k}$ such that $\omega_{k}(d,y,\bm{x})=f_{k}(d,\bm{x})+g_{k}(y,\bm{x})$ for all $d\in\mathcal{D}\setminus\{k\}$
Proof.
We first prove that (a) implies (b). Fix $d_{0}\in\mathcal{D}\setminus\{k\}$ and define $f_{k}(d,\bm{x}):=\omega_{k}(d,y,\bm{x})-\omega_{k}(d_{0},y,\bm{x})$ for any fixed $y\in\mathcal{Y}$. By assumption, $f_{k}(d,\bm{x})$ does not depend on $y$ for $d\neq k$. Next, define $g_{k}(y,\bm{x}):=\omega_{k}(d_{0},y,\bm{x})$ such that we obtain
$\omega_{k}(d,y,\bm{x})=f_{k}(d,\bm{x})+\omega_{k}(d_{0},y,\bm{x})=f_{k}(d,\bm{x})+g_{k}(y,\bm{x}),$
for $d\neq k$, which proves the claim. We next prove that (b) implies (a). If $\omega_{k}(d,y,\bm{x})=f_{k}(d,\bm{x})+g_{k}(y,\bm{x})$ for all $d\in\mathcal{D}\setminus\{k\}$, then for any $d,d^{\prime}\in\mathcal{D}\setminus\{k\}$ we have
$\omega_{k}(d,y,\bm{x})-\omega_{k}(d^{\prime},y,\bm{x})=f_{k}(d,\bm{x})+g_{k}(y,\bm{x})-(f_{k}(d^{\prime},\bm{x})+g_{k}(y,\bm{x}))=f_{k}(d,\bm{x})-f_{k}(d^{\prime},\bm{x}),$
which is independent of $y$. This proves the claim. ∎
G.6 Lemma 6
Lemma 6
If $R(D^{\ast};\ell)$ is constant across $D^{\ast}$, then there exists $C(\bm{x})$ such that $P_{\bm{X}}$-almost surely for every $\bm{x}\in\mathcal{X}$,
$R_{\bm{x}}(d;\ell)=C(\bm{x}),$
for every $d\in\mathcal{D}$.
Proof.
Fix $d\neq d^{\prime}$ and let $A\subset\mathcal{X}$ measurable. Define the two policies
$\pi_{A}(\bm{x})$ $=d\mathds{1}{\left\{\bm{x}\in A\right\}}+d^{\prime}\mathds{1}{\left\{\bm{x}\in A^{c}\right\}},$
$\tilde{\pi}_{A}(\bm{x})$ $=d^{\prime}\mathds{1}{\left\{\bm{x}\in A\right\}}+d\mathds{1}{\left\{\bm{x}\in A^{c}\right\}}.$
By assumption we have
$0=R(\pi_{A};\ell)-R(\tilde{\pi}_{A};\ell)=\int_{\bm{x}\in A}[R_{\bm{x}}(d;\ell)-R_{\bm{x}}(d^{\prime};\ell)]dP_{\bm{X}}(\bm{x})-\int_{\bm{x}\in A^{c}}[R_{\bm{x}}(d;\ell)-R_{\bm{x}}(d^{\prime};\ell)]dP_{\bm{X}}(\bm{x}).$
Rearranging yields
$\int_{\bm{x}\in A}[R_{\bm{x}}(d;\ell)-R_{\bm{x}}(d^{\prime};\ell)]dP_{\bm{X}}(\bm{x})=\int_{\bm{x}\in A^{c}}[R_{\bm{x}}(d;\ell)-R_{\bm{x}}(d^{\prime};\ell)]dP_{\bm{X}}(\bm{x}).$
Adding the left hand side to both sides, we obtain
$2\int_{\bm{x}\in A}[R_{\bm{x}}(d;\ell)-R_{\bm{x}}(d^{\prime};\ell)]dP_{\bm{X}}(\bm{x})=R(d;\ell)-R(d^{\prime};\ell)=0,$
by assumption. This holds for every measurable $A\subseteq\mathcal{X}$. Thus, by standard arguments, we have $R_{\bm{x}}(d;\ell)=R_{\bm{x}}(d^{\prime};\ell)$ almost surely. Since the choice of $d,d^{\prime}$ is arbitrary, we can define $C(\bm{x}):=R_{\bm{x}}(d;\ell)$, which proves the claim.
Appendix H Derivation of Weights
We illustrate the results from Section 3.8 by working out the linear systems in Equations (10), (11), and (12) for the binary case $D,Y\in\{0,1\}$. In this setting, the loss can take eight distinct values $\ell(d;y_{0},y_{1})$, corresponding to two decisions across four principal strata.
Recall that we seek to solve the linear system
$\bm{A}^{b}\bm{w}^{b}=\bm{\ell},$
where $b\in\{\textsc{Std},\textsc{Ext},\textsc{Add}\}$ indexes the standard, exactly identifiable, and additive cases, respectively, as described in Section 3.8. For $b=\textsc{Add}$, the system is unrestricted:
$A^{\textsc{Add}}=\left[\begin{array}[]{cccccccccccc}1&0&0&0&0&0&1&0&1&0&0&0\\ 0&0&1&0&1&0&0&0&1&0&0&0\\ 0&1&0&0&0&0&1&0&0&0&1&0\\ 0&0&1&0&0&1&0&0&0&0&1&0\\ 1&0&0&0&0&0&0&1&0&1&0&0\\ 0&0&0&1&1&0&0&0&0&1&0&0\\ 0&1&0&0&0&0&0&1&0&0&0&1\\ 0&0&0&1&0&1&0&0&0&0&0&1\end{array}\right],\qquad\bm{w}^{\textsc{Add}}=\begin{bmatrix}\omega_{0}(0,0)\\ \omega_{0}(0,1)\\ \omega_{1}(1,0)\\ \omega_{1}(1,1)\\ \omega_{0}(1,0)\\ \omega_{0}(1,1)\\ \omega_{1}(0,0)\\ \omega_{1}(0,1)\\ \varpi_{0,1}(0,0)\\ \varpi_{0,1}(0,1)\\ \varpi_{0,1}(1,0)\\ \varpi_{0,1}(1,1)\end{bmatrix},\qquad\bm{\ell}=\begin{bmatrix}\ell(0;0,0)\\[5.0pt] \ell(1;0,0)\\[5.0pt] \ell(0;1,0)\\[5.0pt] \ell(1;1,0)\\[5.0pt] \ell(0;0,1)\\[5.0pt] \ell(1;0,1)\\[5.0pt] \ell(0;1,1)\\[5.0pt] \ell(1;1,1)\end{bmatrix}$.
The other cases correspond to restrictions of this system:
$\bm{A}^{\textsc{Ext}}=\bm{A}^{\textsc{Add}}_{1:6},\quad\bm{w}^{\textsc{Ext}}=\bm{w}^{\textsc{Add}}_{1:6},\qquad\bm{A}^{\textsc{Std}}=\bm{A}^{\textsc{Add}}_{1:4},\quad\bm{w}^{\textsc{Std}}=\bm{w}^{\textsc{Add}}_{1:4},$
[…]
Furthermore, since $\bm{A}^{b}$ is given, its left null space can be represented by a matrix $\bm{N}^{b}$ whose columns form a basis for $\ker\!\left((\bm{A}^{b})^{\top}\right)$. By standard linear algebra, $\bm{\ell}\in\text{im}(\bm{A}^{b})$ (i.e., the system admits a solution) if and only if
$(\bm{N}^{b})^{\top}\bm{\ell}=\bm{0}.$
This directly yields the equality constraints on $\ell(d;y_{0},y_{1})$ needed for identification in terms of principal strata.
In the standard case, i.e., $b=\textsc{Std}$, the null space can be represented as
$\bm{N}^{\textsc{Std}}=\left[\begin{array}[]{cccc}0&-1&0&0\\ -1&0&0&0\\ 0&0&-1&0\\ 1&0&0&0\\ 0&1&0&0\\ 0&0&0&-1\\ 0&0&1&0\\ 0&0&0&1\end{array}\right],$
with $\operatorname{rank}(\bm{N}^{\textsc{Std}})=4$. This recovers the four constraints of Equation (13), leaving four degrees of freedom to choose the loss $\ell$, as indicated in Table 3.
In the exactly identifiable case, i.e., $b=\textsc{Ext}$, the null space takes the form
$\bm{N}^{\textsc{Ext}}=\left[\begin{array}[]{cc}1&0\\ 0&1\\ -1&0\\ 0&-1\\ -1&0\\ 0&-1\\ 1&0\\ 0&1\end{array}\right],$
with $\operatorname{rank}(\bm{N}^{\textsc{Ext}})=2$, yielding the two constraints of Equation (14) and leaving six degrees of freedom for the loss.
Finally, in the general additive case with $b=\textsc{Add}$, this reduces to
$\bm{N}^{\textsc{Add}}=\left[\begin{array}[]{c}-1\\ 1\\ 1\\ -1\\ 1\\ -1\\ -1\\ 1\end{array}\right],$
with $\operatorname{rank}(\bm{N}^{\textsc{Add}})=1$, giving the single restriction of Equation (15) and leaving seven degrees of freedom.
The above derivation illustrates how the identification conditions translate into explicit equality constraints on the loss values $\ell(d;y_{0},y_{1})$ in terms of principal strata.